### v2

# Knowledge Augmentation in Federation: Rethinking What Collaborative Learning Can Bring Back to Decentralized Data

Abstract
Data, as an observable form of knowledge, has become one of the most important factors of production for the development of Artificial Intelligence (AI). Meanwhile, increasing legislation and regulations on private and proprietary information results in scattered data sources also known as the “data islands”. Although some collaborative learning paradigms such as Federated Learning (FL) can enable privacy-preserving training over decentralized data, they have inherent deficiencies in fairness, costs and reproducibility because of being learning-centric, which greatly limits the way how participants cooperate with each other. In light of this, we present a knowledge-centric paradigm termed Knowledge Augmentation in Federation (KAF), with focus on how to enhance local knowledge through collaborative effort. We provide the suggested system architecture, formulate the prototypical optimization objective, and review emerging studies that employ methodologies suitable for KAF. On our roadmap, with a three-way categorization we describe the methods for knowledge expansion, knowledge filtering, and label and feature space correction in the federation. Further, we highlight several challenges and open questions that deserve more attention from the community. With our investigation, we intend to offer new insights for what collaborative learning can bring back to decentralized data.
Index Terms:
Decentralized data, knowledge extraction, knowledge augmentation, federated learning, data representation, data engineering.
I Introduction
As Artificial Intelligence (AI) exhibits its potential in transforming human society, we are entering a new era where the amount of data and the value generated from data both increase in an unprecedented speed. Statistics show that the total volume of data or information created, captured, copied and consumed globally has experienced a tenfold growth in the past decade and is expected to reach 394 zetabytes in 2028 [1]. This is in large part attributed to the development of data acquisition and sensing techniques as well as the pervasive deployment of Machine Learning (ML) models. For example, research predicts that a single autonomous vehicle can generate as much as 10 gigabytes data per second in the future [2]. The ever-increasing availability of data makes it possible for us to build and train more advanced models such as the Broad Learning System (BLS) [3, 4] and Large Language Models (LLMs) [5, 6].
Figure 1: The conceptual overview of Knowledge Augmentation in Federation (KAF) as a knowledge-centric paradigm for local domain enhancement.
While AI applications are more data-hungry than ever, privacy and proprietary issues arise. The accessibility of private data is in many cases limited by legislation and regulations such as the General Data Protection Regulation (GDPR) [^1] https://gdpr-info.eu/ and the California Consumer Privacy Act (CCPA) [^2] https://www.oag.ca.gov/privacy/ccpa. This to a great extent leads to a paradigm shift from centralized learning to collaborative learning from decentralized data. To address the problem of “data islands”, McMahan et al. [7] from Google proposed Federated learning (FL) [8], a privacy-preserving paradigm based on the concept of Federated Optimization designed for collaborative learning over autonomic user devices. The general workflow of FL involves global model distribution, local training and model upload, and global model aggregation. Essentially, the key factors that distinguish Federated learning (or Federated Optimization) concern the massive system scale, restricted communication and most importantly, the non-Independent and Identically Distributed (non-IID) data assumption [9]. Although extensive research has been conducted to enhance the efficacy and efficiency of FL in diverse scenarios, existing methods hardly go beyond model-based knowledge sharing where models [^3] Models and model parameters are used interchangeably if no ambiguity is caused., model updates, or gradients are exchanged between different parties as the carrier of knowledge. This helps in protecting the raw data but also results in several deficiencies:
  • Intractable fairness: No matter the participants cooperate towards a consensus model for common welfare or personalized models for individual interest, practical fairness is hard to realize. This is because the payoff and cost differ greatly between the parties given pathological data distribution, but learning-centric paradigms such as FL generally struggle to balance the two.
  • Prohibitive costs: Models are “heavy”, and so are the model updates and the gradients. They are generally in the same tensor structure and can take up to gigabytes of network traffic for each transmission. Besides, the cost for local training cannot be overlooked considering the problem of slow convergence over non-IID data [10].
  • Poor reproducibility: models are the carriers of knowledge but are hard to reproduce in massive collaboration due to the randomness in participation. Any participant has to request massive cooperation in order to reproduce a model from scratch.
Is it possible to find an alternative to model-centric collaboration and tackle the above-mentioned issues? In this paper we offer our answer by presenting a knowledge-centric collaboration paradigm termed Knowledge Augmentation in Federation (KAF) or Federated Knowledge Augmentation. The key idea behind KAF is to redefine the way that participants collaborate in a federated system by shifting the focus from learning models to enhancing knowledge across the domains, as illustrated in Fig. 1. Technically, the key difference that KAF brings to the field mainly concerns i) the introduction of knowledge-centric objectives for collaborative optimization in the federated system, and ii) the adoption of more flexible carriers of knowledge without compromising privacy or proprietary rights. More details will be elaborated in Sec. III.
On the one hand, typical distributed learning paradigms seek to build a strong model by aggregating models, gradients or other forms of knowledge from separate domains. Even in heterogeneity-aware approaches such as personalized Federated Learning (pFL) [11], the optimization is still mostly limited to task-specific models which are hardly generalizable. By contrast, Knowledge Augmentation in Federation emphasizes the enhancement of local knowledge associated with each individual participant. On the other hand, The knowledge brought in can take whatever form, e.g, models, gradients, predictions, synthetic data, etc., as long as it serves to enrich or refine the knowledge implied in local data (Fig. 2). This calls for new protocols that enable cost-effective exchange of knowledge. We will elaborate on these points in Sec. III. Compared to existing paradigms, KAF has potential advantages summarized as follows:
  1. 1.
    Improving global knowledge coherence across non-IID data holders,
  2. 2.
    Enhancing local knowledge representations in quantity, quality and utility for better local training or independent learning, and
  3. 3.
    Knowledge-centric collaboration for augmenting local domains with learning-friendly data rather than task-specific models.
Figure 2: The conceptual relationship between data, models and knowledge.
Moreover, we believe knowledge augmentation in federation shows a new roadmap for autonomic intelligence, where each participant can learn independently from knowledge gained collaboratively. In fact, there exist a number of studies that adopted inspiring methodologies in relevance to KAF, but a comprehensive review is so far missing in the literature. As far as we know, this review is the first to assemble related studies in the field under the concept of KAF.
I-A Contributions and Related Surveys
Extensive research has been carried out in the fields of deep learning (DL), representation learning, and data engineering. Meanwhile, demands for data privacy have given rise to broad interest in collaborative learning paradigms such as federated learning, swarm learning and fog learning. In the following we briefly review the representative surveys under these topics and then highlight the difference made in our work.
Deep learning and representation learning. There exist a number of comprehensive surveys on DL-related techniques and applications. Dong et al. [12] briefly reviewed the development of Deep Neural Networks (DNNs) as well as other fundamental elements for DL such as activation functions, model architectures and programming libraries. Similarly, Sarker [13] categorized typical DL approaches in a taxonomy based on the availability of labels. Realistic tasks for supervised learning, unsupervised learning, hybrid learning and other learning paradigms are systematically reviewed. From a different angle, Alzubaidi et al. [14] paid attention to the impact of data scarcity on DL model training and how to deal with it. Besides, how to learn strong representations from the data is believed to be the key to DL algorithm design [15]. Considering the connection between data representations with domain-specific knowledge, surveys have been conducted with focuses on various domains such as network representation learning [16], multi-modal representation learning [17], and graph representation learning [18].
Data engineering and management. The development of Big Data techniques not only paves the way for complex learning systems, but also calls for more learning-centric data engineering. Generally speaking, data engineering entails the whole lifecycle of data management from acquisition and pre-processing to utilization, each with significant impact on the outcome of learning [19]. Among the problems in the field, data imbalance has attracted the most attention, and a diversity of pre-processing and learning-based methods were investigated in the literature [20, 21]. In addition, surveys such as [22, 23, 24] shed light on the bias and fairness in data. However, collaborative engineering of knowledge beyond data is yet to be covered in existing surveys.
Federated learning and its variants. Kairouz et al. [9] presented one of most comprehensive surveys on FL, covering the efficiency, effectiveness, privacy, robustness and fairness issues. They also highlighted the challenges related to ill-distributed data for both cross-device and cross-silo FL scenarios. Some surveys focus on particular aspects of an FL system such as privacy preservation [25], decentralization [26], and heterogeneous FL [27]. Application-specific studies are also reviewed in the literature, e.g., FL for Internet of Things (IoT) [28], smart cities [29], and healthcare [30, 31, 32]. Ji et al. [33] termed the variants of FL paradigms as “Federated X Learning”, which represents the fusion of FL with other emerging ML paradigms. Nonetheless, these surveys, without any exception, are learning-centric and did not investigate what FL can bring back to the data.
Table I highlights the aspects that distinguish this review paper from the rest in the literature.
TABLE I: The comparison between our work and the surveys in related fields regarding the research themes.
Knowledge-centric collaboration learning data & feature engineering
DL & representation learning (e.g., [12, 13, 14, 15, 16, 17, 18])
Data engineering & management (e.g., [19, 20, 21, 22, 23, 24])
Federated learning (e.g., [9, 25, 26, 27, 28, 29, 30, 31, 32, 33])
Ours
Our contributions. In this paper we advocate a knowledge-centric vision and focus on what learning can bring back to data in a federated system. Main contributions of this work are summarized as follows:
  • A revisit to data processing and learning across “data islands” from the perspective of knowledge.
  • New understanding of collaborative learning as a bridge to knowledge augmentation in federation.
  • Insights for what KAF can bring up for the evolution of AI and the new regimes of collaborative intelligence.
II Background and Preliminaries
In this section we provide a brief introduction to the basics of federated learning and data engineering, which helps to better understand the ideas and topics covered in our work.
II-A Federated Learning
Federated learning is a distributed machine learning paradigm for privacy-preserving collaboration across a multitude of devices. The basic system model of FL includes user devices (often termed clients) that possess private data and a central server that coordinates the learning process [34]. The standard process of FL [7] is organized in rounds each comprised of the following steps: i) Client selection: a fraction of clients are selected for the upcoming round of training; ii) Model distribution: the server synchronizes with the selected clients by sending them the global model (consensus model); iii) Local training: each selected client trains the model received on its local data using gradient-based methods, which produces a local model; iv) Model upload: selected clients upload their local models to the server; and v) Model aggregation: the server aggregates the local models into a new global model. These steps cycle for a number of rounds until some criterion is met.
A unique challenge that distinguishes federated learning from other learning paradigms is the multi-party ownership of data. This significantly complicates the way we operate across multiple disjoint sets of data, i.e., “data islands”. On the one hand, in the FL setting we need to deal with a second layer of data distribution, that is, a distribution of similar but different domains represented by all the client-owned local datasets in the system. On the other hand, we still face the notorious problems, such as data imbalance, sample corruption and label noise, probably within each of the local datasets. More importantly, the orchestrator (e.g., a central server in most cast) is typically agnostic of local data distribution due to privacy constraints.
II-B Data Engineering
Data engineering in a general sense refers to the practice in the management of data spanning its whole lifecycle [23]. Data engineering mainly deals with the problems in data collection and integration, data storage, data processing, and data analytics [^4] Systematic aspects such as data pipeline automation are also elements of the theme but do not fall within the scope of this paper.. The boom of ML and AI applications has brought increasing attention to learning-centric data engineering methods particularly including data augmentation, data balancing and feature engineering. In the following we briefly explain these concepts.
Data augmentation and balancing. The goal of data augmentation is to facilitate learning or analytics by artificially enhancing a dataset in size and sample diversity. Widely adopted in image, text and rational data pre-processing, data augmentation usually involves the generation of synthetic samples and the transformation of raw data. On the one hand, data synthesis introduces artificial samples to the dataset by exploiting interpolation, extrapolation, statistical models, generative models, or any combination of these techniques [35]. Note that data synthesis is also a common strategy for data balancing or re-sampling where the distribution of samples is amended by over-sampling the minority class. This is the very rationale behind the classic methods such as Synthetic Minority Oversampling Technique (SMOTE) [36], ADAptive SYNthetic (ADASYN) [37], and Mixup [38]. On the other hand, data transformation helps to diversify the features of data by applying certain types of transforms. The transformation of data are usually domain-specific, e.g., rotation, flipping and rescaling for image data, and random deletion, insertion and rephrasing for text data.
Feature Engineering (FE). Raw data are more often than not represented by attributes that can be noisy, highly correlated and even corrupted. The goal of feature engineering is to create a new set of features that better represent the knowledge implied and filter out the noise and misinformation. Typical recipes for feature engineering includes feature selection and feature transformation and extraction. Specifically, feature selection reduces the dimensionality of the feature space by filtering for the most important features, which is usually performed based on statistical evidence or measures from information theory [39]. As for feature transformation and extraction, it not only applies numerical operation(s) like rescaling to individual features, but also involves re-encoding the data with new combinations of features [40]. Considering the target value (label) as a special feature, methods such as label correction and generation can also be categorized as the implementation of FE. In practice, domain knowledge can be leveraged in the FE process. The confluence of feature engineering and representation learning unleashes the power of AI over massive data.
III Conceptual Overview
In this section we provide a holistic view of the proposed KAF paradigm. We start with introducing the (suggested) system architecture and move on to the problem setting. This is followed by the outline of key approaches that will be elaborated in Secs. IV, V and VI.
III-A System Architecture
As shown in Fig. 1, the system model of KAF is similar to that of (horizontal) federated learning. To be specific, a central server is deployed as the coordinator at the center of the hub-and-spoke topology. A multitude of participants $\mathcal{K}=\{u_{1},u_{2},\ldots\}$ connect to the server to send and receive shared knowledge. Each participant $u_{k}$ has the exclusive access to its local data $D_{k}$ that is not allowed to be shared for privacy or proprietary reasons.
A key difference is that each participant in KAF is not necessarily capable of training deep ML models. This is because the optimization objective for KAF is not learning-centric but knowledge-centric, which will be elaborated in the following section.
III-B Problem Statement
The fundamental problem behind privacy-sensitive collaboration is the “data islands” that causes non-IID data across the participants (i.e., clients in FL). From the view of knowledge, this problem can be interpreted as a global domain segmented into multiple overlapped sub-domains [^5] Also simply referred to as (local) domains if no ambiguity is caused. corresponding to the local data partitions on the clients.
The proposed paradigm, namely knowledge augmentation in federation, seeks to address the problem by exchanging and enhancing the knowledge in private data independent on task-specific model learning. In other words, the primary goal of KAF is to maximize the effective capacity of knowledge for each participant [^6] The terms participant and client are used interchangeably. $u_{k}\in\mathcal{K}$ that holds a private dataset $D_{k}$, where $\mathcal{K}$ represents the entire population in the system. In addition, we also expect each local domain to align with the global domain through collaboration but without deviating too much from its original data distribution $\mathcal{D}_{k}$. This translates to a prototypical objective function expressed in (1),
$\max_{\pi_{k}}F_{k}=\underbrace{\nu\big{(}\Theta_{\pi_{k}}(D_{k};\mathcal{K})\big{)}}_{enhancement}-\\ \alpha_{k}\Big{(}\beta_{k}\cdot\underbrace{\sum_{u_{i}\in\mathcal{K}/u_{k}}d\big{(}\Theta_{\pi_{k}}(D_{k};\mathcal{K}),\Theta_{\pi_{k}}(D_{i};\mathcal{K})\big{)}}_{alignment\;loss}+\\ (1-\beta_{k})\cdot\underbrace{d\big{(}\Theta_{\pi_{k}}(D_{k};\mathcal{K}),D_{k}\big{)}}_{regularization}\Big{)},$ (1)
where the function $\nu(\cdot)$ evaluates the capacity of knowledge given a dataset, $\Theta_{\pi}(\cdot;\mathcal{K})$ denotes the operations following a strategy (or protocol) $\pi$ on local data collaboratively performed across the federation $\mathcal{K}$, and $d(\cdot,\cdot)$ measures the distributional distance (e.g., Kullback-Leibler (KL) divergence) between two datasets. The two client-specific hyper-parameters $\alpha_{k}$ and $\beta_{k}$ weight the three terms in the objective.
Collectively, one can derive a system-wide objective from (1) if all the participants are cooperative and stick to their strategies $\{\pi_{1},\pi_{2},\ldots\}$:
$\max_{\{\pi_{1},\pi_{2},\ldots\}}\sum_{k=1}^{|\mathcal{K}|}F_{k}$ (2)
Besides, objective (1) indicates that each participant in KAF is self-motivated and contributes towards its own benefit. Particularly, the augmented knowledge represented in local data makes it possible for each $u_{k}$ to carry out independent model training or any other downstream tasks, making the cooperation knowledge-centric rather than learning-centric. On this point, KAF also differs from the objective of pFL [41] where each client aims to minimize a task-specific loss w.r.t. a specific model.
III-C Outline of Approaches
Through in-depth investigation, we figure out that there has been early effort in related areas especially in the conjunction of FL and data engineering. Even though the motivation behind these studies do not explicitly align with the goal of KAF, we found their methodologies somewhat suitable for serving as the building blocks for KAF. By categorizing these studies, we present a suggested roadmap for KAF comprised of three approaches: federated knowledge expanding, federated knowledge filtering, and federated label and feature space correction (Fig. 3.
Figure 3: The structure of this article.
IV Methods for Knowledge Expansion
As discussed previously, a federated system entails a set of knowledge domains that collectively comprise the global domain. Intuitively, it is easy to understand that greater overlaps between the local domains can facilitate faster and better learning. For example, it is commonly agreed that the global model’s convergence can be improved by reducing the “non-IIDness” of local data. In fact, this can be achieved by outspreading the knowledge contained in each local domain.
In this section, we investigate popular approaches for collaborative knowledge expansion under the FL setting, where the implications of expansion are two-fold. First, this represents a roadmap for extending the boundaries of local knowledge domains through introducing exogenous information from other clients. Second, the expansion of knowledge also refers to the augmentation of data with the aims of re-balancing local distribution and generating diversified instances. Hence, the key criteria for this research category are:
  • Seeking to extend local domains of knowledge through collaboration;
  • Improving the diversity of data samples within each local domain based on information exchange.
We summarize that there are two major lines of research in this direction. In the following content, we first introduce existing studies for extrapolating the knowledge in local data by information sharing and reconstruction in a federated system. Then we revisit the research on enhancing local data in terms of size, quality and distribution via data augmentation.
IV-A Federated Knowledge Extrapolation
We conceptually describe knowledge extrapolation as the methodology of filling the missing part of knowledge that is not explicitly expressed by the raw data. This process usually involves predicting or approximating implicit samples locally by utilizing some global knowledge contributed and shared by other participants in the system (Fig. 4). In the context of federated learning, knowledge extrapolation can be very useful when all or most of the local datasets are sparse in terms of information density. For example, Sanyal et al. [42] considered a healthcare scenario where numerous Internet of Medical Things (IoMT) devices are simultaneously working towards sensing a particular phenomenon. In this case each IoMT device acts as an independent data collector that simultaneously generate time series data. The challenge here is that these readings need to be collectively utilized for analysis, but it is not encouraged to continuously push them to a central server due to privacy concerns and network constraints. To this end, the authors proposed a Federated Filtering Framework (FFF) [42], where the key component on each local device is a predictor for approximating the true sensing data. The IoMT devices communicate with a fog server that aggregates and utilizes the predictors in order to reconstruct the full data matrix. For communication efficiency, the timing of local model update and aggregation is determined by dynamic thresholds.
Figure 4: A schematic showing the general framework of knowledge extrapolation. The goal is to obtain the missing part of knowledge in the local domain.
Sparsity is one of the major obstacles that can largely undermine the efficacy of machine learning. The problem has been extensively explored in the context of recommender systems and the design of recommendation algorithms. More specifically, data are represented as a user-item interaction matrix that is huge in dimensionality but low in information density, which is caused by the natural pattern of user behavior and the scale of e-commerce platforms. Although there are many solutions to it in the literature, most of the existing studies make the assumption that the whole data matrix is available for analysis. However, a centralized collection of user behavior data may not be permitted in many situations such as cross-border e-commerce.
As one of the most popular approaches to recommender systems, Collaborative Filtering (CF) an be viewed as a data model that is able to extrapolate incomplete knowledge (i.e., unrecorded interactions) using the dense representation of factor matrices. Approved by the Netflix Prize, Matrix Factorization (MF) is the classic implementation of CF, where the raw user-item interaction matrix $\boldsymbol{R}\in\mathbb{R}^{m\times n}$ can be approximated as
$\boldsymbol{R}\sim\boldsymbol{X}^{T}\boldsymbol{Y},$ (3)
where $\boldsymbol{X}\in\mathbb{R}^{k\times m}$ and $\boldsymbol{Y}\in\mathbb{R}^{k\times n}$, with $m$ and $n$ being the number of users and items, respectively.
By adapting MF to FL settings, researchers have found new paths to personalized recommendation without a centralized storage of user vectors (which are likely to contain user-private information). Ammad-ud-din et al. [43] first described this new approach as Federated Collaborative Filtering (FCF), where all the clients keep their interaction data private on their own devices. Based on a reconstruction loss and stochastic gradient descent (SGD), clients in FCF update their user-specific (user-factor) matrices locally given the latest item-factor matrix from the server. In the meantime, each client also calculates its corresponding part of the item-specific gradients and feeds back to the server for item-factor matrix update. Following a similar framework, Yang et al. [44] further investigated a multi-behavior scenario, where both item viewing and purchase are recorded but the former is kept private by each user. The authors proposed to binarize both factor matrices and utilized a caching mechanism to reduce communication payload for collaborative parameter update. Considering that a single form of information may not be sufficient for accurate recommendation, the knowledge provided by implicit and explicit feedback can be integrated using a probabilistic model. In this case, a specific decomposing strategy is needed for model updating without exposing users’ private information [45].
Neural Collaborative Filtering (NCF) [46] replaces the matrix factorization process of CF by using deep neural networks (DNNs) to learn the user embedding vectors $\boldsymbol{X}=\{\boldsymbol{x_{1}},\boldsymbol{x_{2}},\ldots\}$ and item embedding vectors $\boldsymbol{Y}=\{\boldsymbol{y_{1}},\boldsymbol{y_{2}},\ldots\}$. With an embedding dimensionality of $k$, these learned latent vectors are then concatenated and fed into a Multi-Layer Perceptron (MLP) that outputs the predicted score of preference.
NCF differs from MF-based approaches in the sense that knowledge is encapsulated in deep neural networks (DNNs), rather than factor matrices. In such way each user and item can be encoded independently and fed into a prediction model (e.g., another DNN) for training and inference. Inspired by this property, Perifanis and Efraimidis [47] designed a federated NCF framework that extends that basic NeuMF model where each client holds three sets of parameters including their user profile, item profile, and the DNN’s weights. With user profiles kept private, they proposed a secure aggregation algorithm, MF-SecAgg, to perform server-side parameter update, with separate steps for item profile update and DNN update respectively. With a similar motivation, Jiang et al. [48] introduced an adaptive Differential Privacy (DP) mechanism with decaying noise into the communication process of Federated NCF. This protects the exchange of model gradients against inference attacks.
For effective knowledge extrapolation, it is intricate to strike a balance between preserving learned representations locally and leveraging the common knowledge shared globally. To address this, Li et al. [49] designed a gating dual-encoder structure based on Variational Auto-Encoder (VAE) for personalized CF in a federated setting. They introduced a gating network followed by two encoders, one for pertaining local knowledge and the other for learning user representations in the global latent space. By mapping the user-item interaction vectors to separate latent subspaces, the loss of personalized knowledge (due to federated aggregation) is effectively mitigated.
Generally, user-private knowledge needs to be stored locally, no matter as a factor matrix (for MF-based methods) or a neural network’s parameters (for NCF). This facilitates global knowledge fusion but also increases the burden on end devices. One may ask that is it possible to make local knowledge “invisible”? Singhal et al. [50] offered their answer by presenting a novel reconstruction-based framework FedRecon. Inspired by Model-Agnostic Meta Learning (MAML), FedRecon allows stateless participants to rebuild their local model every time they join the federation with the knowledge provided by the global model. This facilitates more flexible collaboration at scale and, more importantly, indicates that personalized knowledge can be extrapolated by leveraging globally shared knowledge.
IV-B Federated Data Augmentation
Data augmentation, as a powerful technique for diversifying the data, can also enable effective expansion of knowledge when deployed in a federated system. The value of federated data augmentation is two-fold. On the one hand, it directly introduces new knowledge—typically represented by synthetic samples—to the set of local data. On the other hand, multiple data owners can collaboratively perform data augmentation under the guidance of a global coordinator, which brings higher consistency of data distribution across the parties in the federation.
The implementations of augmentation generally involve applying transforms on existing samples or synthesizing new samples based on a priori. Note that the synthesis of data and the training of downstream models can be decoupled, hence a simple implementation is to perform data augmentation in the local pre-processing stage without interleaving with the following FL process. Kundu et al. [51] followed this approach and developed a federated monkeypox disease detection model on the basis of offline image augmentation. Specifically, a Cycle GAN is pre-trained and deployed locally to generate additional samples to enhance local training. Similarly, Ullah et al. [52] developed an FL-based smart healthcare system, where they employed oversampling in order to address the class imbalance of chest X-ray imaging data. More specifically, SMOTE is adopted to mitigate the scarcity of minority-class samples such as those with “COVID-19” and “Viral” labels [53]. One potential limit is that unreasonable samples could be generated without using domain-specific knowledge as constraints. Despite being easy-to-implement, offline data augmentation is agnostic of any global knowledge and thus may not serve well for the global model’s convergence.
The potential limit of offline data augmentation can be explained by considering the boundary of knowledge for each domain, wherein offline data augmentation is essentially analogous to linear and non-linear interpolation within the domain. By contrast, through introducing exogenous information for local data augmentation (i.e., collaborative data augmentation), more generalizable knowledge can be obtained to expand the domain. This could lead to higher gains for the downstream tasks.
TABLE II: Summarizing the studies related to federated knowledge expansion in terms of methodological highlights, drawbacks, datasets for experiments, and the form of knowledge for exchange.
Refs. Category Knowledge carrier Highlights Drawbacks Datasets
[42] Fed. knowledge extrapolation predictors & filters $\bullet$ remote data reconstruction via predictors
$\bullet$ dynamic communication frequency based on error monitoring
$\bullet$ uniform filter parameters across all devices MHEALTH
[43, 45] Fed. knowledge extrapolation factor matrices $\bullet$ decoupling private and shareable knowledge via matrix factorization $\bullet$ linear growth of communication cost with the raw matrix’s dimensions MovieLens, FilmTrust, Epinions, and in-house data
[44] Fed. knowledge extrapolation binarized matrices $\bullet$ finer categorization of private knowledge
$\bullet$ binarization and caching
$\bullet$ limited to CF applications JD, Tmall, User Behavior, and MovieLens
[47, 48, 49] Fed. knowledge extrapolation latent vectors and prediction models $\bullet$ better protection of private knowledge
$\bullet$ dilution of knowledge in model aggregation due to sparsity
$\bullet$ extra costs for local encoding and model transmission
$\bullet$ knowledge redundancy in the shared latent vectors
MovieLens, Lastfm, Foursquare NYC, Ta-Feng, Amazon-Instant-Video, and QB-article
[50] Fed. knowledge extrapolation the global model $\bullet$ model-agnostic and stateless collaboration
$\bullet$ generalizable to diverse tasks
$\bullet$ lack of consideration of data heterogeneity MovieLens and Stack Overflow
[51, 52] Federated data augmentation task-specific models $\bullet$ easy to implement
$\bullet$ low risk of privacy leakage
$\bullet$ only local knowledge exploited in augmentation
$\bullet$ hard to generalize
COVID-19 radiography, pox virus, and chest X-ray images
[54] Federated data augmentation attention scores from generative models $\bullet$ CVAE for local augmentation
$\bullet$ light-weight knowledge exchange
$\bullet$ high local training cost Fashion-MNIST and CIFAR-10
[55, 56] Federated data augmentation the average of local samples $\bullet$ controllable anonymity
$\bullet$ low communication costs
$\bullet$ hard to generalize for being loss-dependent FEMNIST, FMNIST, CIFAR-10, CIFAR-100, CINIC-10, Airfoil, NO2, and UTKFace
[57, 58] Federated data augmentation prediction model $\bullet$ Augmented images with cross-institute knowledge $\bullet$ tightly-coupled with the task fastMRI and IXI
[59] Federated data augmentation shared generator $\bullet$ decoupled knowledge with decoupled models $\bullet$ imbalanced amount of knowledge for the generator and discriminator fastMRI, IXI, BRATS, and in-house data
Although the collaboration of different clients naturally provides a rich set of knowledge sources for data augmentation, how to share the knowledge without compromising user privacy remains the key challenge. Essentially, a client needs to perceive the distribution of data outside its domain without raw data exchange. On this point, sharing a generator (i.e., a generative model that produces new samples) has proved to be an useful mechanism. Wen et al. [54] proposed a generative models-based federated data augmentation strategy (FedDA), where a Conditional AutoEncoder (CVAE) is deployed and collaboratively trained on each client to tackle non-IID data distribution. Instead of sharing the whole generator, they extracted the decoder’s hidden-layer features as the shared knowledge and introduced a knowledge distillation loss to guide the training. This effectively reduces communication overheads. For privacy preservation, noise samples are mixed into each local dataset. Intuitively, one can adapt traditional data augmentation, such as the popular MixUp [38] strategy, to the federated setting. However, the difficulty here is that raw samples need to be shared if we want to mix up local and global knowledge. Over the generated samples $(\boldsymbol{x}^{\prime},y^{\prime})$ this “global MixUp” strategy theoretically corresponds to minimizing the following loss:
$\ell_{rawmix}(h_{\theta}(\boldsymbol{x}^{\prime}),y^{\prime})=\ell\Big{(}h_{\theta}\big{(}(1-\lambda)\boldsymbol{x}_{i}+\lambda\boldsymbol{x}_{j}\big{)},(1-\lambda)y_{i}+\lambda y_{j}\Big{)}$ (4)
where $\ell$ and $h_{\theta}$ is the task-specific loss criterion and the model to be trained.
To achieve privacy preservation, Yoon et al. [55] first presented Mean Augmented Federated Learning (MAFL), a simple framework that enables clients to exchange the average of local samples (called “mashed” data) along with model updates. Further, through approximating the global MixUp loss (Eq. 4) with Taylor expansion, they derived that the approximation can be done using local samples and the averaged samples from other clients (Eq. (5)):
$\ell_{FedMix}=\mu\ell\big{(}h_{\theta}(\mu\boldsymbol{x}_{i}),y_{i}\big{)}+\lambda\ell\big{(}h_{\theta}(\mu\boldsymbol{x}_{i}),\bar{y}_{j}\big{)}+\lambda\frac{\partial\ell}{\partial\boldsymbol{x}}\bar{\boldsymbol{x}}_{j},$ (5)
where $\mu=1-\lambda$, $\bar{\boldsymbol{x}}_{j}=\frac{1}{|\xi|}\sum_{j\in\xi}\boldsymbol{x}_{j}$ and $\bar{y}_{j}=\frac{1}{|\xi|}\sum_{j\in\xi}y_{j}$ represent the mean of features and labels over a private data batch $\xi$ from other clients, respectively. The derivative $\frac{\partial\ell}{\partial\boldsymbol{x}}$ is evaluated at $\boldsymbol{x}=\mu\boldsymbol{x}_{i}$ and $y=y_{i}$.
This motivated them to devise the FedMix [55] algorithm that performs collaborative MixUp augmentation and training based on a decomposable loss function. Zhang et al. [56] further extended this framework for regression tasks using bilateral neighborhood expansion and evaluated it on various public datasets. Empirical results show that their algorithms yielded stronger performance than FedMix [55].
A special case of federated data augmentation is the reconstruction of samples using knowledge from multiple domains, which has proved very useful for accelerated medical imaging such as Magnetic Resonance Imaging (MRI) [57, 58]. To adapt to the disparity of imaging operators between different sites, Elmas et al. [59] deployed an unconditional GAN at each site to generate synthetic high-quality samples, with a private discriminator to fit site-specific knowledge and a globally shared generator for prior knowledge embedding. A tailored reconstruction loss was used to guide the training of local generators.
IV-C Summary
In this section we revisit two representative lines of research that share the common idea of expanding user/client-specific domains of knowledge in a federation. Federated knowledge extrapolation mainly deals with data sparsity and missing information in local domains, whilst federated data augmentation focuses on mitigating non-IID local data distribution. The two approaches methodologically differ in the two aspects.
  • The way new knowledge is introduced: knowledge extrapolation typically relies on representation learning to map multiple sources of information into one or several vector spaces, where sparse features are transformed into dense embeddings. These embeddings and the encoders act as the carriers in the process of knowledge aggregation and distribution. On this basis, predictions can be made to extend the local domain of knowledge, for example, by extrapolating unobserved interactions in the recommender system. By contrast, data augmentation can be performed independently to enhance the representations of knowledge, which is essentially analogous to linear interpolation [^7] In fact, it is exactly the idea behind the MixUp strategy.. Through collaboration, knowledge from outside the domain can be introduced by information fusion at sample level. This directly results in an expanded size of data and typically does not depend on any task-specific models.
  • Motivation and objectives: the extrapolation of knowledge is usually coupled with the optimization of task-specific objectives such as the accuracy of item recommendations. For example, in federated collaborative filtering, knowledge from different users are implicitly fused in the process of the global item vector update and local user vector update, guided by a cost function that quantifies the reconstruction loss. Federated data augmentation, on the other hand, aims to regulate the local data distribution towards the IID setting. In most cases, augmentation is performed without an explicit objective of optimization.
Each of the two approaches has certain limitations, as summarized in Table II. For collaborative knowledge extrapolation, domain-specific knowledge is embedded as factor matrices or vectors that are optimized under the supervision of known observations. This means that inaccurate information could be inferred and disseminated once the training samples are very noisy or manipulated by malicious clients. For collaborative data augmentation, neither interpolation-based or generative model-based schemes can theoretically guarantee the fidelity of synthetic samples. Certain mechanisms are necessary to constrain the process of augmentation against invalid patterns [60, 61]. So far, this problem still needs further investigation in the federated settings.
V Methods for Knowledge Filtering
A federated system is usually characterized by a rich set of data sources which, on the one hand, facilitate a collaborative way of learning but, on the other hand, bring about knowledge redundancy as well as the risk of data poisoning [62, 63]. The issue has become increasingly concerning with the advance of generative models that are able to fabricate adversarial samples. For example, a hyperdimensional computing (HDC) model is employed in [64] to craft poisoned samples that can bypass byzantine-robust defenses. This calls for more effective mechanisms for knowledge filtering.
We in this paper define knowledge filtering as the techniques for excluding misinformation in form of improper data (and its derivatives [^8] Including but not limited to embeddings, models and gradients.) and distinguishing different regions of the domain in terms of value for learning. Knowledge filtering can involve engineering and algorithm design at both levels across and within the domains. This section reviews the emerging studies that seek to address the discrepancy of knowledge regarding its validity and utility for the downstream task.
V-A Client-level Filtering
Poisoned datasets can severely threaten the validity of knowledge, which in turn causes performance degradation of the federated model learned. Being targeted or untargeted, data poisoning attacks are generally assumed to be carried out by one or multiple participants. Therefore, a straightforward approach to defending against data poisoning is to identify and exclude these malicious clients.
Although a variety of schemes have been developed for the generation of adversarial samples that are difficult to detect individually, they still inevitably change the distribution of data, for instance, with flipped labels or fake samples blended in. Through empirical studies on label-flipping attacks, Khuu et al. [65] found that, with a single round of attack, a small subset of adversarial clients can result in dramatic drop in the model’s recall on the targeted class as well as notable decrease of overall accuracy. Inspired by the Shapley Additive Explanation (SHAP), they developed a client filtering algorithm using Support Vector Machines (SVMs) to classify clients based on four features derived from SHAP values. However, SHAP values are computation-intensive for each client as we need to calculate a score matrix of all image-label pairs. By contrast, detecting data poisoning from the perspective of model is low-cost and more easy-to-implement. Model-based filtering approaches are based on the fact that local training on poisoned data will result in model update that are likely to deviate from that of benign clients. For example, the FedZZ [66] framework detects untargeted data poisoning attack based on the cosine similarity between a client’s update vector and a reference zone-aggregated vector. The proposed zone-based detection scheme groups the clients into multiple zones and calibrates the zone set progressively by accuracy. Similarly, Wei et al. [67] investigated the impact of targeted data poisoning attacks on local training by examining the eigenvalues of the covariances of gradient updates. Through theoretical analysis and empirical study, they proved that gradients from adversarial clients and honest clients form clearly separable clusters, which provides forensic evidence for accurate client filtering. Nonetheless, challenges remain as it was also observed that the timing of attacks greatly influences the effectiveness of statistic-based defense strategies [67].
Apart from adversarial clients with malicious intention, low-quality information such as noise and irrelevant samples may exist in local data. In our previous work [68], we conducted empirical studies inspecting the impact of low-quality information on the process of collaborative learning (Fig. 5). The result demonstrates that poor data conditions significantly slow down the convergence whilst undermining the global model’s performance.
Figure 5: Results of our experiments over 100 clients under four disparate conditions of data that involve IID, non-IID and low-quality on-device data. Experiment details are described in [68], with more observations included.
Instead of dirty labels, in [68] we examined the performance drop caused by low-quality clients that trained on local data affected by various forms of noise. We discovered that the engagement of low-quality knowledge can be as detrimental as pathological data distribution, and more importantly, the existence of both led to the worst results of convergence and final accuracy. To address the problem, a novel client filtering strategy FedProf [68] is presented, where the latent representations from the global model’s hidden layer are extracted and compressed into client-specific profiles. Based on the Kullback–Leibler (KL) divergence between the profiles, low-quality clients are filtered out before each FL round probabilistically.
Noise data filtering at client level can be practically imperative especially for Internet of Things (IoT) applications. This is because real-time sensing is inherently susceptible to environmental disturbance and hardware flaws. For example, real-time data captured by image sensors are sensitive to violent camera shake and occlusion in the Internet of Vehicles (IoV) scenario. Targeting at this problem, Lei et al. [69] presented a filtering framework OES-Fed that works on the vehicle level. The framework first identifies a group of reference vehicles (i.e., vehicles with high-quality data) based on model accuracy, and sets their model parameters as the initial centers for K-means clustering. Vehicles that deviate significantly from the centers in the parameter space are detected as noise data carrier. In addition, Kalman filters and exponential smoothing are employed to incorporate the historical changes of model parameters. This enables dynamic assessment of local knowledge in response to Continual Learning (CL) scenarios [70] where new data keep rolling in.
V-B Sample-level Filtering
Studies on sample-level filtering endeavor to improve the quality of local knowledge by differentiating each sample within the local datasets. In contrast to the coarse-grained approaches that either keep or filter out entire local data domains, sample-level filtering corresponds to a fine-grained roadmap for enhancing the gain of local training and in turn facilitating global model convergence. However, we face unique challenges as it is typically not allowed in FL to access the raw attributes of each individual instance of data. This problem is especially intractable in the healthcare scenarios where the Electronic Health Records (EHRs) of patients are highly sensitive and probably heterogeneous across different organizations in the federation. By heterogeneous it means that samples from different clients scatter in inconsistent feature spaces that correspond to different permutations of features. To address this, Thakur et al. [71] proposed a knowledge abstraction and filtering framework, where the core component is the client-specific filtering modules that can be cascaded. Each filtering module starts with a native encoder that map local data to a unified latent space and ends with a gating operator that uses a globally shared knowledge abstraction vector to filter local samples’ representations. The novelty of this framework boils down to the knowledge abstraction vector which is trained collaboratively to embed consensus knowledge and guides the filtering of low knowledge. Nonetheless, the mechanism cannot directly identify and filter out problematic samples. Motivated by the knowledge distillation paradigm [72, 73], Lu et al. [74] proposed to employ conditional generators on each client to generate proxy samples for ensemble distillation. To address the loss of fidelity in synthetic data, the authors connected a filter to the head of each local model whose outputs are then modulated based on the reference labels. This helps in refining local knowledge without the need for a public data pool.
As a matter of fact, knowledge filtering can be performed simultaneously at client level and sample level. Adopting such a hybrid approach, Xu et al. [75] cast light on cloud-edge collaboration where “dirty” data on untrustworthy devices may significantly compromise the learning process. A novel framework called Safe is presented, which at the client level resorts to Alternating Direction Method of Multipliers (ADMM) for detecting suspicious clients that behave like data poisoning attackers. At sample level, data samples on these (potentially) malicious are then filtered based on K-means clustering and Euclidean distance. A shortcoming here is that the Euclidean distance in the raw feature space may not serve as a good criterion in the sample-level filtering in the presence of high dimensionality and sparsity. Moreover, Safe was only evaluated on the learning task of Support Vector machines (SVMs), which means that whether the framework can generalize to deep learning tasks remains questionable.
TABLE III: Summarizing the studies related to federated knowledge filtering in terms of methodological highlights, drawbacks, datasets for experiments, and the form of knowledge for exchange.
Refs. Category Knowledge carrier Highlights Drawbacks Datasets
[65] Client-level filtering derived features from SHAP $\bullet$ model-independency and statistical interpretability $\bullet$ high computational cost of SHAP MNIST and Fashion-MNIST
[66, 67] Client-level filtering gradients or update vectors $\bullet$ scalable and easy-to-implement
$\bullet$ effective against diverse types of data pollution
$\bullet$ relatively high cost of communication
$\bullet$ vulnerability to collusion
Fashion-MNIST, EMNIST, CIFAR10, and LFW
[68] Client-level filtering statistically compressed representations $\bullet$ very low cost of communication
$\bullet$ stronger privacy protection
$\bullet$ requirement for dense layers in the model structure GasTurbine, EMNIST, and CIFAR-10
[69] Client-level filtering task-specific models $\bullet$ scalable and effective against diverse types of data pollution
$\bullet$ historical behavior included for smoothing
$\bullet$ relatively high cost of communication MNIST, CIFAR-10, and Vehicle Classification dataset
[71] Sample-level filtering knowledge abstraction vector and the prediction model $\bullet$ abstraction of global knowledge for local filtering
$\bullet$ localized filtering module adaptable to heterogeneous domains
$\bullet$ relatively high cost of communication
$\bullet$ vulnerability to erroneous and adversarial samples
CURIAL, eICU-CRD, and MIMIC-III
[…]
$\bullet$ global synthesis regulated by local knowledge ensemble
$\bullet$ high cost of communication
$\bullet$ vulnerability to adversaries
Fashion-MNIST, SVHN, CIFAR-10, and Digits-5
[75] Client+sample filtering model parameters $\bullet$ knowledge filtering at both levels using a multi-stage design
$\bullet$ edge-cloud hierarchical system for stronger privacy
$\bullet$ distance measure ill-suited to high dimensionality and sparsity
$\bullet$ applicable to DSVM only
Iris and WineQuality
[76] Fed. data re-weighting task-specific models $\bullet$ IWDS for achieving fast convergence without loss of accuracy $\bullet$ knowledge outside local domains unexploited CIFAR-10 and Fashion-MNIST
[77] Fed. data re-sampling task-specific models $\bullet$ easy to implement with any classic re-sampling methods $\bullet$ knowledge outside local domains unexploited MIMIC-III
[78] Fed. data re-weighting Gini coefficients and models $\bullet$ learnable strategy for client-specific re-weighting control
$\bullet$ local Gini coefficients reflecting local distributional knowledge
$\bullet$ huge state space for the DRL agent
$\bullet$ high communication costs
MNIST, FMNIST, and CIFAR-10
V-C Federated Data Re-balancing
The notorious problem of non-IID data across the local domains in the federation not only hampers the convergence of the global model, but also leads to the drop of accuracy when the final model is evaluated on a presumably balanced test set. From the angle of knowledge, these phenomena can be explained from two aspects. First, theoretical evidence [79, 80] has revealed the convergence speed of the global model, on a certain ML task, is very much determined by the divergence of parameter updates (or accumulated gradients) from the clients. This essentially comes down to discrepancy in local knowledge as reflected by the conditional distribution of features and labels, which results in diverging local gradients towards different local optima. Second, non-IID data across the system usually implies that clients could have conflicting knowledge. Once learned and aggregated into a global model, client-specific knowledge gets diluted or neutralized, which consequently undermines the model’s ability to generalize.
Data re-balancing has proved its effectiveness in class-imbalanced learning where re-sampling is the most common technique adopted in the literature [81, 82]. By scaling up (down) the sampling probabilities of tail (dominant) classes, the local model is learned on a reshaped distribution of data. This can also be interpreted as a means of knowledge filtering where class-specific knowledge is kept by a different ratio. Considering that local data can be extremely imbalanced in non-IID settings, it is natural to explore collaborative knowledge filtering by means of data re-balancing. Tang et al. [76] investigated the impact of the celebrated class-balanced loss (CBLoss) [83] when applied to the local training of FL. Two interesting observations were made. First, data re-sampling narrowed the gap between the probability distributions of label sampling between different clients, which consequently accelerates the global model’s convergence. Second, though reaching the plateau at a faster rate, re-sampling resulted in an inferior model that yielded lower accuracy on local data. These two findings motivated them to devise a Imbalanced Weight Decay Sampling (IWDS) strategy that gradually smooths an imbalanced sampling towards uniform sampling. More specifically, given $n_{k,c}$ samples of class $c$ on client $u_{k}$, the sampling weight for a class-$c$ instance is:
$w^{(t)}_{k,c}=\frac{1-\beta_{t}}{1-\beta_{t}^{n_{k,c}}},$ (6)
with $\beta_{t}<1$ being the sampling parameter that decays from $\beta_{m}$ to $\beta_{0}$:
$\beta_{t}=\beta_{m}+(\beta_{0}-\beta_{m})\cdot\gamma^{t}$ (7)
where $\gamma$ is the decay rate parameter and $t$ is the index of the federated round. Using this dynamic re-sampling results in a round-varied ratio of probability that a class-$c_{i}$ instance gets sampled to that of a class-$c_{j}$ instance:
$r^{(t)}_{k,i,j}=\frac{w^{(t)}_{k,i}}{w^{(t)}_{k,j}}=\frac{1-\beta_{t}^{n_{k,j}}}{1-\beta_{t}^{n_{k,i}}}.$ (8)
This strategy leverages imbalanced re-sampling to speed up the convergence at the early stage of FL and gradually approaches uniform sampling for higher final accuracy [76], i.e., $r^{(t)}_{k,i,j}$ gets closer to 1 with $\beta_{t}$ decreasing. Although distributional information is utilized locally to enhance local training, knowledge outside the local domain is left unexploited.
Ouyang et al. [78] formulated a different weighting rule for each sample as a power function of class frequency. After normalization the sampling probability $P_{k,c}$ for a class-$c$ instance is:
$P_{k,c}=\frac{n_{k,c}^{\beta_{k}}}{\sum_{i=1}^{C}n_{k,i}^{\beta_{k}}},$ (9)
where $C$ is total number of classes and the client-specific parameter $\beta_{k}$ controls the sampling probability distribution on client $u_{k}$. To dynamically determine $\{\beta_{k}|k=1,2,\ldots\}$ for the training cohort, a global coordinating agent based on Deep Reinforcement Learning (DRL) was developed. More specifically, the agent collects local and global model parameters as input state, makes decisions on $\{\beta_{k}|k=1,2,\ldots\}$, and receives rewards defined based on class-specific accuracy and Gini coefficients of local distribution.
Most of the classic re-sampling techniques for imbalanced learning can be easily applied to the local training step in FL. This corresponds to a naïve implementation of federated data re-balancing where clients act independently based on local distribution. Chen [77] conducted empirical studies to investigate the impacts of local data re-balancing on the efficacy of FL. Over-sampling, under-sampling and the combination of these two were evaluated on a healthcare-related logistic regression task over ten clients. They also filtered out the outliers in the pre-processing stage by z-score. Two interesting finds were reported: i) Under-sampling caused a slight performance decrease in the final model’s accuracy and precision, and ii) the hybrid method (combining over-sampling and under-sampling) showed the best performance across the board in all metrics. These observations imply a strong potential of data re-balancing in the federation even without any knowledge exchange between the domains. However, whether the phenomenon still holds for more complex tasks remains under-explored.
V-D Summary
In this section we demystify the representative approaches to knowledge filtering within a federation. With rising concerns of data-centric attacks, accurately identifying adversarial clients that train on poisoned local data can be a top priority with regard to system security and robustness. We argue that fine-grained filtering at sample level can be employed jointly with client selection to provide more reliable evidence for adversary detection. As of data re-balancing, existing approaches stand on an implicit assumption that uniform data distribution is ideal for local training. However, this may not always hold. On the one hand, the local learner could under-fit on the “special knowledge” in local data when more attention is paid to the tail classes [76]. On the other hand, there is no sufficient theoretical evidence that supports imbalanced sampling, provided that the test data distribution is unknown. Although extensive effort has been made in addressing class-imbalanced learning, collaborative knowledge filtering for regression tasks, e.g. drug response modeling, is hardly explored. Moreover, existing studies pay little attention to the subtle line between data imbalance and non-IID data, the understanding of which can help us better reshape the knowledge locally and globally. In Table III we summarize the strength and limits of existing studies in relevance to knowledge filtering in the federation.
VI Label and Feature Space Correction
In this section we review the advance of research on label and feature space correction, by which we refer to the branch of studies that employ collaborative mechanisms to make correction or transformation on data labels or features. Provided that the quality of local data can be dubious especially in a cross-device federated system, it is important to rectify the samples in each local domain. In contrast to knowledge filtering that aims to exclude misinformation, amendments of samples usually involves data engineering techniques and feature engineering methods. Hence, in the following content we survey the state-of-the-art under two different categories: federated data correction and federated feature engineering.
VI-A Federated Data Correction
As discussed in Section V, data noise commonly exists in a federation. Among various types and forms of noise, mislabeled samples are believed to have salient impact on supervised learning. Many researchers endeavored to address the problem of noisy labels in the context of FL, that is, federated noisy label learning (FNLL) [84, 85, 86, 87]. Considering the tie between a client and its local domain, it is intuitive to first identify the “noisy clients” [88, 89]. Xu et al. [88] proposed FedCorr, a multi-stage label denoising framework that streamlines a pre-processing stage and a fine-tuning stage prior to the traditional FL process, all executed in a federated fashion. In the federated pre-processing stage, two distinct mechanisms are employed for detecting noisy clients and correcting noisy labels. On the one hand, each client computes an estimate of the Local Intrinsic Dimensionality (LID) based on its local model’s predictions, which are collected by the central server for noisy client detection. As an useful measure that connects dimensionality and discriminability [90, 91], the LID for a vector $\boldsymbol{x}\sim\mathcal{D}$ is defined as:
${LID}_{x}=\lim_{r\rightarrow 0}\lim_{\epsilon\rightarrow 0}\frac{\log F_{d_{x}}((1+\epsilon)r)-\log F_{d_{x}}(r)}{\log(1+\epsilon)},$ (10)
where $d_{x}$ is the distance between $\boldsymbol{x}$ and a randomly drawn vector $\boldsymbol{x}^{\prime}$ from $\mathcal{D}$ and $F_{d_{x}}(\cdot)$ represents the cumulative distribution function (CDF) of $d_{x}$. As models trained on noisy data tend to report larger LID scores, the server then uses this rule to spot these clients for label correction.
FedCorr [88] adopts a confidence-based rule to find and correct mislabeled data on noisy clients. Specifically, the subset of samples (controlled by a proportion $\pi$) that yield the highest losses is first selected, following which the predictions of the global model $f_{G}$ are used for relabeling. Formally, the final subset of samples $D^{\prime}_{k}$ (for a noisy client $u_{k}$) to be relabeled is defined in (11):
$D^{\prime}_{k}=\big{\{}(\boldsymbol{x},y)|(\boldsymbol{x},y)\in\underset{\hat{D}\subseteq D_{k}\atop|\hat{D}|=\pi|D_{k}|}{\arg\max}\sum_{(\boldsymbol{x},y)\in\hat{D}}\ell(f_{G}(\boldsymbol{x}),y)\\ \text{ and }\max f_{G}(\boldsymbol{x})\geq\theta\big{\}},$ (11)
where $\ell$ is the loss function and $D_{k}$ is the local data from client $u_{k}$. The labels of $D^{\prime}_{k}$ are then corrected to the predictions by the global model $f_{G}$.
There are two major defects in FedCorr [88]: i) incompatibility to general loss functions other than cross entropy, and ii) susceptibility to dishonest clients who tamper with their LID scores. Targeting at the second issue, Wang et al. [89] built upon FedCorr and redesigned its federated pre-processing stage to achieve robust label correction. The key idea behind is to craft an efficient Zero-Knowledge Proof (ZKP) [92] protocol by which the server (as the verifier) can verify the computation integrity of the LID scores reported by the clients (as the provers) in a batched manner. The proofs are generated both before and after label correction. Only calculation in the form of arithmetic circuits is verifiable by the protocol, which limits its potential application to other data correction schemes. Similarly, the FedELC framework by Jiang et al. [84] detects clients that carry noisy data and corrects them in two separate stages. The first stage is almost the same as in [88, 85], whilst the second stage creates learnable soft labels for data correction optimized towards the minimization of a linear combination of three different entropy losses. FedELC was evaluated on a rich set of noise data settings including the CIFAR-N [93] dataset from Amazon Mechanical Turk which features a high noise rate of human-annotated labels.
Noisy label detection and correction can also be executed directly at sample level. Based on label uncertainty estimation [94], Zeng et al. [95] started from utilizing a well-trained global model’s predictions to estimate the probability that samples of a class $i\in\mathcal{C}$ are correctly labeled in a local dataset $D_{k}$:
$P_{k,i}=$ $\sum_{j\in\mathcal{C}}p(y=i|y^{*}=j)p(y^{*}=j|y=i)$
$\approx$ $\sum_{j\in\mathcal{C}}p(y=i|y^{*}=j)\mathbb{E}_{\boldsymbol{x}\in D_{k,y=i}}p(y^{*}=j;\boldsymbol{x},\theta_{G})$
$=$ $\mathbb{E}_{\boldsymbol{x}\in D_{k,y=i}}\sum_{j\in\mathcal{C}}p(y=i|y^{*}=j)p(y^{*}=j;\boldsymbol{x},\theta_{G})$
$=$ $\mathbb{E}_{\boldsymbol{x}\in D_{k,y=i}}\sum_{j\in\mathcal{C}}p(y=i|y^{*}=j;\boldsymbol{x},\theta_{G})p(y^{*}=j;\boldsymbol{x},\theta_{G})$
$=$ $\mathbb{E}_{\boldsymbol{x}\in D_{k,y=i}}\sum_{j\in\mathcal{C}}p(y=i,y^{*}=j;\boldsymbol{x},\theta_{G})$
$=$ $\mathbb{E}_{\boldsymbol{x}\in D_{k,y=i}}p(y=i;\boldsymbol{x},\theta_{G})$
$=$ $\frac{1}{|D_{k,y=i}|}\sum_{v=1}^{|D_{k,y=i}|}p(y_{v}=i;\boldsymbol{x}_{v},\theta_{G}),$ (12)
where $D_{k,y=i}$ is the data subset on client $u_{k}$ labeled as class $i$, $\mathcal{C}\subset\mathbb{W}$ is the label space, and $\theta_{G}$ denotes the global model’s parameter. Next, the global estimate of the probability of correctly-labeled class $i$ can be obtained using weighted average consensus [95] over the population $\mathcal{K}$:
$P_{i}=\sum_{k\in\mathcal{K}}\frac{|D_{k,y=i}|}{\sum_{k\in\mathcal{K}}|D_{k,y=i}|}P_{k,i},$ (13)
which resembles the model aggregation rule in the canonical FedAvg [96].
A practical concern of federated data correction is that the participants can be self-interested, which means that they may not stick to the protocol if their gain of participation cannot cover the cost of data correction. In light of this issue, Yan et al. [97] first modeled the payoff of clients as a function of the model’s accuracy, payment from the task publisher, and the data correction cost. It is worth noting that the authors employed an iterative denoising strategy that corrects a fraction of $\alpha$ noise samples per round. To reduce the fraction of noisy data from $\epsilon$ to $\epsilon^{\prime}$ given a denoising cost per sample $c$ and data size $|D_{k}|$, the total cost of data correction for client $u_{k}$ is:
${Cost}_{k}=|D_{k}|c(\log_{\alpha}\epsilon-\log_{\alpha}\epsilon^{\prime}).$ (14)
Yan et al. [97] considered two different situations of federated denoising. In the first case, clients are cooperative and work towards to the optimal social welfare, i.e., the maximization of the sum of client payoffs. In the second case, every client is assumed self-interested and only motivated to maximize its own payoff, which corresponds to a non-cooperative game.
VI-B Federated Feature Engineering
Prior to the boom of deep learning, feature engineering had proved very useful in alleviating the “curse of dimensionality” and allowing the model to learn more generalizable knowledge from the data. In the context of KAF, by federated feature engineering we refer to a new thread of research with particular focus on collaborative knowledge augmentation by means of aligning, selecting and transforming the feature space (Fig. 6). This process can either be independent of the downstream task or somehow coupled with a specific federated learning protocol. In both cases, feature engineering across multiple domains in the federation poses new challenges in several aspects.
  • Disordered or misaligned attributes: in cross-institutional cooperation it is common to deal with different views of data where the local datasets are subject to different permutations of attributes [71].
  • Heterogeneous feature space: from a single domain to a multitude of domains with partial overlap in the feature space, the heterogeneity brings great difficulty to the design of autonomic feature engineering.
  • Varied significance of features: non-IID settings usually mean bias and special knowledge within each domain, which will greatly influence the importance of features when measured statistically.
These challenges have catalyzed rising interest in federated heterogeneous learning and also motivated cutting-edge studies on new feature engineering paradigms. In this section we discuss the advance in federated feature engineering, with particular focuses on collaborative feature selection and feature transformation.
Figure 6: A schematic illustration of collaborative feature engineering under the KAF paradigm.
Fed-FiS [98] represents an early exploration of collaborative feature selection in the federation based on information theory. The study follows a training-free, task-agnostic approach to assessing the importance of features. Locally, a client $u_{k}$ independently evaluates the feature-feature mutual information (FFMI) and the feature-class mutual information (FCMI) over the local feature and label set $\{f^{(k)}_{1},f^{(k)}_{2},\ldots\}\cup C^{(k)}$ by Eqs. (15) and (16):
${FFMI}^{k}_{i,j}=\sum_{x\in f^{(k)}_{i}}\sum_{y\in f^{(k)}_{j}}p(x,y)\log\frac{p(x,y)}{p(x)p(y)},$ (15)
${FFCI}^{k}_{i}=\sum_{x\in f^{(k)}_{i}}\sum_{c\in C^{(k)}}p(x,c)\log\frac{p(x,c)}{p(x)p(c)}.$ (16)
Based on these mutual information-based measures, each client performs clustering to select a subset of features which are then uploaded to the central server. In the second stage, the server uses FFMI and FCMI again to score each feature, which in turn helps the clients identify the best local feature subset. Banerjee et al. [99] further extended this framework using an improved ranking scheme based on Pareto optimization where each feature is treated as a solution characterized by FFMI and FCMI. The features are then ranked based on the relation of domination and reported back to the clients.
Fu et al. [100] studied the feature selection problem in the Vertical Federated Learning (VFL) setting where the set of entities are associated with consistent identifies but represented by different features. The proposed framework, FEAST, involves three types of clients for collaborative feature selection: a selected party, an active party, and candidate parties. In the pre-processing stage, discretization is first applied to each feature, combined with stratified sampling to reduce computation and communication costs. This is followed by iterative operations and interactions between the parties. Specifically, a party is selected to report its statistical variables which are generated by grouping local features and re-encoding the instances. The candidate parties receive the statistical variables and score each feature based on Conditional Mutual Information (CMI). The feature scores are then collected by an active party to perform global ranking to produce the best feature subset $S_{k}$ (s.t., $|S_{k}|=m$ given $m$ as the number of local features to select) for each client $u_{k}$, which collectively represents the solution to the following optimization:
$\min_{S_{k}\subseteq\mathcal{X}_{k}}H(Y|\{S_{1},S_{2},\ldots\}).$ (17)
where $H(Y|\{S_{1},S_{2},\ldots\})$ is conditional entropy of the label $Y$ given the union of locally selected feature subsets $\{S_{1},S_{2},\ldots\}$. This framework offers useful insights for communication-efficient exchange of knowledge for collaborative feature selection, but the protocol does not consider dishonest parties (clients) in the system. As a matter of fact, feature engineering techniques can be utilized to defend against data poisoning attacks strengthened by feature manipulation. This is also termed feature poisoning (FP) in the literature [101]. Experimental evidence has demonstrated that, despite being highly disguised, FP attacks can be effectively countered by the Random Deep Feature Selection (RDFS) [102] strategy.
Beside feature selection, feature transformation represents another typical approach to data engineering. By mapping samples to a new space of features derived from raw data, feature transformation not only helps to reduce information redundancy but also refines the representation of knowledge that is difficult to learn. However, it is not trivial for data holders in a federated system to perform feature transformation collaboratively. A practical problem that stands out is the computation cost incurred by performance evaluation provided numerous combinations of transformed features. Moreover, privacy issues arise since transformation can involve multiple features coming from different domains, which necessitates Secure Multi-Party Computation (MPC). To address these issues, Fang et al. [103] designed FLFE, a federated feature engineering framework for efficient, privacy-preserving multi-party collaboration. The core techniques employed include a sketching scheme for quantizing local features, a DNN-based classifier for judging the value of features, and a mask-based feature exchange protocol for transformation. This brings advantages in communication efficiency and also preserves data privacy without homomorphic encryption that limits the diversity of transformations.
TABLE IV: Summarizing the studies related to federated label and feature space correction in terms of methodological highlights, drawbacks, datasets for experiments, and the form of knowledge for exchange.
Refs. Category Knowledge carrier Highlights Drawbacks Datasets
[88, 89] Fed. label correction local intrinsic dimensionality (LID) and models $\bullet$ global pre-processing plus “clean” clients-only fine-tuning
$\bullet$ local knowledge encapsulated in LID
$\bullet$ incompatibility to general loss functions CIFAR-10, CIFAR-100, Clothing1M, and simulated data
[84] Fed. label correction warm-up models and class-wise loss vectors $\bullet$ distance-aware aggregation to mitigate noisy knowledge
$\bullet$ learnable soft labels for local correction
$\bullet$ sensitive to the setting of multiple hyper-parameters CIFAR-10/100, CIFAR-10/100-N, Clothing1M
[95] Fed. label correction class-wise confidence thresholds and models $\bullet$ consensus-based method to ensemble decentralized knowledge $\bullet$ reliance on a pre-trained model for label uncertainty estimation MNIST, USC-HAD, CIFAR-10, and Clothing1M
[97] Fed. label correction denoising rate decisions and models $\bullet$ payoff modeling for game-based interaction
$\bullet$ iterative denoising with a tunable fraction
$\bullet$ reliance on accurate estimate of noise rate of private data MNIST and CIFAR-10
[98, 99] Fed. feature selection MI-based scores and feature subsets $\bullet$ low communication costs
$\bullet$ model-agnostic and task-independent
$\bullet$ lack of personalization
$\bullet$ applicable to single-label tasks only
SL-KDD99 and anonymized credit card transactions (ACC)
[100] Fed. feature selection statistical variables and feature scores $\bullet$ MI-based and model-independent
$\bullet$ efficient exchange of statistical knowledge
$\bullet$ vulnerability to dishonest parties
$\bullet$ limited to VFL on tabular data
MIMIC III, PhysioNet, Census-Income, and Nomao
[104] Fed. feature selection feature subsets $\bullet$ low communication cost
$\bullet$ unsupervised and model-agnostic feature selection
$\bullet$ training required for each local OCSVM
$\bullet$ lack of personalization
Bot-IoT, ACC, KDD99, and DEFT
[105] Fed. feature transform. expression of derived features $\bullet$ low communication cost and reduced search space $\bullet$ very high cost of feature subset evaluation OpenML586 and Airfoil
[106] Fed. feature selection probability vectors associated with features $\bullet$ probabilistic feature subset selection and Bayesian aggregation
$\bullet$ significantly reduced network traffic
$\bullet$ convergence guarantee for IID data only MAV and WEarable Stress and Affect Detection (WESAD)
[103] Fed. feature transform. sketches of features $\bullet$ Quantile Data Sketch and mask-based feature exchange
$\bullet$ model-free evaluation
$\bullet$ pre-training required for the DNN-based feature judge 166 datasets from OpenML
A major obstacle for adapting Automated Feature Engineering (AutoFE) methods (e.g., DIFER [107], OpenFE [108] and Cognito [109]) to the federated system is the prohibitive cost of communication. To this end, Overman and Klabjan [105] have made early effort and devised a federated framework for collaborative AutoFE. In the horizontal FL setting, clients independently carry out AutoFE and report the expression of derived features to the server. The server then randomly applies masks to the union of these features to generate feature subsets. An iterative halving strategy is used to find the best mask (e.g., optimal feature subset) after running multiple rounds of standard FL for performance evaluation. For vertical FL setting, feature values across the entire local dataset need to be shared after homomorphic encryption so as to fuse features from different domains. For hybrid settings, importance of common features across multiple domains can be assessed progressively with the training process.
Federated feature engineering can facilitate the refinery of knowledge in a broad range of scenarios. For example, Zhang et al. [104] proposed a federated feature selection framework for Internet of Things (IoT) applications where each device (as a client) has very limited computational resources. In this framework, local features $\{f^{(k)}_{1},f^{(k)}_{2},\ldots\}$ on each client $u_{k}$ are first filtered using a threshold defined based on Symmetric Uncertainty (SU) and Feature Average Relevance (FAR):
${SU}(f^{(k)}_{i},f^{(k)}_{j})=2\Big{[}\frac{{MI}(f^{(k)}_{i},f^{(k)}_{j})}{H(f^{(k)}_{i})+H(f^{(k)}_{j})}\Big{]},$ (18)
${FAR}(f^{(k)}_{i})=\frac{1}{d_{k}}\sum_{j=1}^{d_{k}}{SU}(f^{(k)}_{i},f^{(k)}_{j}),$ (19)
where $d_{k}$ is the feature dimensionality for client $u_{k}$, $H(\cdot)$ is the entropy measure and $MI(\cdot,\cdot)$ denotes the mutual information between two features. Next, One-Class Support Vector Machine (OCSVM) and hierarchical clustering are employed on each client to identify the subset of most significant features. The server then calculates the intersection of these subsets and synchronizes with the clients. The solution was evaluated on four classification tasks in relevance to IoT applications. Experimental results demonstrated that competitive and even higher performance could be achieved with the federated feature subset, with improved training efficiency.
Cassará et al. [106] argued that a small subset of features are adequate for collaborative learning in interconnected Autonomous Driving Systems (ADSs). To this end, they developed a federated feature selection algorithm that combines on-vehicle feature filtering and edge-based feature subset aggregation. Locally, each participating vehicle (as a client) optimizes a feature selection probability vector such that the conditional entropy $H(y|S)$ (given feature subset $S$ and labels $y$) is minimized, which corresponds to an Associated Stochastic Problem (ASP). The edge server then aggregates these probability vectors into a consensus one and feeds it back to the vehicles. This interaction keeps going iteratively until the global vector converges. The Kolmogorov-Smirnov test with preset parameters are used for checking convergence. By evaluating this algorithm, the authors claimed that a 99% compression rate for network traffic was achieved since as few as 24 (out of 2166) features were selected for collaborative training on the MAV dataset, with negligible impact on model performance. This finding indicates that federated feature engineering can be as effective as model-based compression methods (e.g., quantization, pruning, etc.) [110] in terms of saving learning costs.
VI-C Summary
In this section we highlight emerging studies that work on the dimensions in which data are represented through collaborative effort from multiple parties. We argue that they symbolize an important approach to knowledge augmentation in federation. On the one hand, federated data correction attempts to amend the misinformation carried by the data to avoid erroneous knowledge, with particular focus on the labels. Reference knowledge, in most of the label-correction methods, is encapsulated in the global model. On the other hand, federated feature engineering focuses on identifying the most important attributes or the most expressive combination of them. In this context, the statistical properties of features are often utilized as the measure of knowledge they can provide.
Although different in methodologies (summarized in Table IV, the two lines of research have some common points in terms of implementation. First, both data correction and feature engineering are typically implemented as part of a multi-stage procedure in the federation. For example, data correction is usually cooperated by the clients in the pre-processing stage before the actual training process such as FL. Similarly, federated feature engineering usually involves repetitive local feature assessment and client-server information exchange. In addition, we discover that both approaches can be sensitive to the discrepancy of data distribution across different domains. A unified posterior (that describes the conditional probability of labels) or an optimal feature subspace may not work well for every individual client in the federated system.
VII Open Challenges and Questions
Through revisiting diverse approaches towards more effective, efficient learning from various perspectives of knowledge, we argue that there is a large space ahead to explore in the domain. The advance in knowledge discovery combined with the promising potential of federated learning is expected to open up new opportunities for knowledge-centric, learning-driven collaboration in the future. In the paper we highlight urgent challenges and open questions that involve but are not limited to the ways knowledge is learned, represented and exchanged as well as how to enhance Large Language Models (LLMs) with KAF. We expect to offer unique insights for research beyond knowledge for learning.
VII-A Decentralized Peer-to-Peer Knowledge Augmentation
Extensive effort has been made in decentralized machine learning since it has natural advantages in scalability, flexibility and byzantine-robustness [111, 112]. However, existing approaches are mostly learning-centric, which means that their objectives are shaped towards the optimization of a specific group of models or a family of relevant tasks. This makes the collaboration costly and hard to reproduce, because the model as the knowledge carrier is very sensitive to the uncertainty in federated training. This uncertainty boils down to the population involved as well as the sequence of knowledge propagation in the network, which also raises the cost of reproducing the model.
Recent studies on decentralized matrix factorization (DMF) present a domain-specific path to expand local knowledge domain via peer-to-peer (P2P) communications [113, 114, 115]. DMF can be regarded as a special case of decentralized knowledge augmentation. With a novel vision, Abdelmoniem et al. [116] described a decentralized knowledge routing framework where a client can describe the knowledge requested and find another well-matched client for collaborative learning. This is done by sharing artificial samples with the proxy devices (called knowledge discoverers) responsible for request routing and handling. Even though the objective employed is still task-specific and learning-oriented (deficiencies discussed in Sec. I still apply), it inspires us to explore new directions for decentralized knowledge exchange and knowledge matching between domains.
Clues for further study: knowledge abstraction and matching, blockchain, zero-knowledge proof, and knowledge routing.
VII-B Light-weight Knowledge Courier and Efficient Exchange
A key factor in knowledge augmentation in federation is the exchange of knowledge between distributed local domains. For example, the item factor matrix is shared in the federation as the carrier of domain-specific knowledge in federated collaborative filtering, whilst the global model takes this role in federated label denoising. Whatever form it takes for knowledge exchange, the reduction of communication cost persists as a top priority in practice.
On the application layer, the roadmap for efficient knowledge exchange mainly encompasses two approaches: i) decreasing the communication payload incurred by the courier of knowledge, and ii) reducing the number of communication rounds if iterative interactions are needed. Although FL is more communication-efficient than traditional distributed learning, it is worth noting that models, model updates and gradients are “heavy”. Provided that it is not necessary to exchange the “full picture” of knowledge in many situations such as federated knowledge filtering and label space correction, how to craft a flexible protocol to avoid information redundancy emerges as new problem to be resolved.
Clues for further study: model-free knowledge representation, information bottleneck, and Bayesian compression.
VII-C Federated LLM Augmentation with Exogenous Knowledge
It is impossible for any model, even as complex as the latest LLMs [5, 6, 117], to learn and store comprehensive knowledge that covers all domains of training data. Moreover, language models usually suffer in reproducing accurate knowledge, a.k.a. the problem of hallucinations. Recent studies have paid attention to the federated Retrieval Augmentation Generation (RAG) [118], which is a promising technique for introducing exogenous knowledge to assist in model reasoning. These attempts target at addressing the “information scattering” effect [119] by searching multiple domains, yet the trade-off between performance gain and retrieval cost is mostly overlooked. The effectiveness and efficiency of federated RAG hinge on two factors, that is, whether a selected exogenous knowledge base well matches the given query, and whether the content retrieved can provide qualified knowledge. On the one hand, it is very expensive in time to reliably determine whether a huge database contains the piece of knowledge desired. On the other hand, the scarcity of user feedback brings difficulties for assessing the usefulness of contents retrieved. To harness the potential of federated retrieval for enhanced LLM services, deeper investigation into the evaluation of exogenous knowledge is needed.
Clues for further study: federated RAG, federated search systems, query-content evaluation, and knowledge base summary.
VII-D KAF for Unlabeled data
We observed that the majority of researches in relevance to knowledge augmentation and federated learning were designed for or evaluated on supervised learning tasks. This is mainly attributed to the prior knowledge implied by the labels that usually serve as the most important criterion for many operations. For example, feature engineering typically relies on feature-label mutual information, and data re-balancing strategies re-weight data points by class label. It is also worth mentioning that most existing empirical studies on federated learning associate non-IIDness to the marginal distribution of labels. However, this largely limits their application to a broader categories of unlabeled data such as text corpora, surveillance footage and browsing records. We argue that the main challenge for KAF over unlabeled data lies in the perplexity of knowledge within each domain. This means that, without the a priori from annotations, the increased uncertainty of data (e.g., measured in entropy) may result in lower confidence in the knowledge learned. In this case, more attention should be paid to the intrinsic structures of data and the impact of non-IID data beyond classification.
Clues for further study: active learning, contrastive learning, soft labels, and knowledge distillation.
VII-E Can KAF Make Independent Training Superior over FL?
We argue that refined representation of knowledge is what collaborative learning can bring back to data. This is the key point that distinguishes KAF from what have been explored in the relevant field such as Federated Learning. As discussed, knowledge augmentation in federation can enhance the local knowledge in each domain by means of expansion, filtering and correction. This creates possibility for each participant to learn independently after the augmentation. But how competitive will it be as compared to Federated Learning? The answer is uncertain and hinges on multiple factors including local data size, the test data distribution, and the utility of knowledge gained via KAF. In terms of performance, KAF can increase the size and diversity of samples necessary for stochastic gradient-based local training, with augmented knowledge from foreign domains for generalizable model. In terms of cost, FL struggles in communication efficiency with large models and huge client population, whilst KAF plus independent training has tempting potential when equipped with efficient knowledge exchange protocols. From both perspectives, the superiority of KAF is still an open-ended question.
Clues for further study: benchmarking, sample-efficient learning, and marginal gains of collaboration.
VII-F Can Knowledge be Borrowed and Unlearned in a Federation?
“The right to be forgotten”, as stated in privacy-protecting regulations, has given rise to the interest in the research of Machine Unlearning [120, 121]. With potential value in achieving security, privacy, usability, and fidelity in various application scenarios [121], the concept of machine unlearning has also been adopted in FL systems to enable a specific client’s knowledge to be removed from what has been learned by the model [122]. However, existing approaches are basically limited to the unlearning of class-oriented knowledge [123]. More importantly, the process of unlearning can cost as much as re-training in terms of time and resources. This calls for further study on approximate federated unlearning to enable efficient and generalizable knowledge erasure. In the context of KAF, collaboratively augmented data can be reckoned as “borrowed” knowledge. New mechanisms are required in order to accurately identify the synthetic data points that characterize the designated part of knowledge to be forgotten.
Clues for further study: federated machine unlearning, model inversion, gated knowledge transfer, noise learning, and influence-aware filtration.
VIII Conclusion
As the value of data assets and privacy issues arise, it is of utmost importance for organizations to make full use of proprietary data whilst seeking cooperation with their counterparts. Although the collaborative paradigm of federated learning has paved the way for the sharing of knowledge across domains with privacy preservation, research in the field hardly explored the possibilities beyond model-centric learning. We argue that knowledge sharing is not just limited to the purposes of model optimization, but can also serve as the key to local domain augmentation, which substantially diversifies the way how participants can benefit from the federation. With this vision, we in this paper present knowledge augmentation in federation, a conceptual paradigm for the expansion, filtration and correction of knowledge across decentralized domains. We revisit emerging research in the field via the lens of knowledge and provide a category of studies according to their methodologies. Through comparison and discussion we point out that the carrier of knowledge and the protocol for knowledge exchange are the most influential factors for KAF to be effective and efficient. Based on our investigation, we further highlight a number of open challenges that are not adequately addressed by existing effort, and also pose two open-ended questions to inspire more theoretical and empirical exploration in the relevant threads of research.
Acknowledgment
This work is supported by the National Natural Science Foundation of China, National Social Science Foundation of China Post-funded project, Basic and Applied Basic Research Project of Guangzhou, and the Fundamental Research Funds for the Central Universities.

### v3

@@
-Knowledge Augmentation in Federation: Rethinking What Collaborative Learning Can Bring Back to Decentralized Data
+Task-Agnostic Federation over Decentralized Data: Research Landscape and Visions
 Abstract
-Data, as an observable form of knowledge, has become one of the most important factors of production for the development of Artificial Intelligence (AI).
-Meanwhile, increasing legislation and regulations on private and proprietary information results in scattered data sources also known as the “data islands”.
-Although some collaborative learning paradigms such as Federated Learning (FL) can enable privacy-preserving training over decentralized data, they have inherent deficiencies in fairness, costs and reproducibility because of being learning-centric, which greatly limits the way how participants cooperate with each other.
-In light of this, we present a knowledge-centric paradigm termed Knowledge Augmentation in Federation (KAF), with focus on how to enhance local knowledge through collaborative effort.
-We provide the suggested system architecture, formulate the prototypical optimization objective, and review emerging studies that employ methodologies suitable for KAF.
-On our roadmap, with a three-way categorization we describe the methods for knowledge expansion, knowledge filtering, and label and feature space correction in the federation.
+Increasing legislation and regulations on private and proprietary information results in scattered data sources also known as the “data islands”.
+Although Federated Learning-based paradigms can enable privacy-preserving collaboration over decentralized data, they have inherent deficiencies in fairness, costs and reproducibility because of being learning-centric, which greatly limits the way how participants cooperate with each other.
+In light of this, we investigate the possibilities to shift from resource-intensive learning to task-agnostic collaboration especially when the participants have no interest in a common goal.
+We term this new scenario as Task-Agnostic Federation (TAF), and investigate several branches of research that serve as the technical building blocks.
+These techniques directly or indirectly embrace data-centric approaches that can operate independently of any learning task.
+In this article, we first describe the system architecture and problem setting for TAF.
+Then, we present a three-way roadmap and categorize recent studies in three directions: collaborative data expansion, collaborative data refinement, and collective data harmonization in the federation.
 Further, we highlight several challenges and open questions that deserve more attention from the community.
-With our investigation, we intend to offer new insights for what collaborative learning can bring back to decentralized data.
+With our investigation, we intend to offer new insights about how autonomic parties with varied motivation can cooperate over decentralized data beyond learning.
 Index Terms:
-Decentralized data, knowledge extraction, knowledge augmentation, federated learning, data representation, data engineering.
+Generic knowledge, Decentralized data, knowledge exchange, data engineering, federated systems.
 I Introduction
 As Artificial Intelligence (AI) exhibits its potential in transforming human society, we are entering a new era where the amount of data and the value generated from data both increase in an unprecedented speed.
-Statistics show that the total volume of data or information created, captured, copied and consumed globally has experienced a tenfold growth in the past decade and is expected to reach 394 zetabytes in 2028 [1].
-This is in large part attributed to the development of data acquisition and sensing techniques as well as the pervasive deployment of Machine Learning (ML) models.
-For example, research predicts that a single autonomous vehicle can generate as much as 10 gigabytes data per second in the future [2].
-The ever-increasing availability of data makes it possible for us to build and train more advanced models such as the Broad Learning System (BLS) [3, 4] and Large Language Models (LLMs) [5, 6].
-Figure 1: The conceptual overview of Knowledge Augmentation in Federation (KAF) as a knowledge-centric paradigm for local domain enhancement.
-While AI applications are more data-hungry than ever, privacy and proprietary issues arise.
-The accessibility of private data is in many cases limited by legislation and regulations such as the General Data Protection Regulation (GDPR) [^1] https://gdpr-info.eu/ and the California Consumer Privacy Act (CCPA) [^2] https://www.oag.ca.gov/privacy/ccpa.
-This to a great extent leads to a paradigm shift from centralized learning to collaborative learning from decentralized data.
-To address the problem of “data islands”, McMahan et al.
-[7] from Google proposed Federated learning (FL) [8], a privacy-preserving paradigm based on the concept of Federated Optimization designed for collaborative learning over autonomic user devices.
-The general workflow of FL involves global model distribution, local training and model upload, and global model aggregation.
-Essentially, the key factors that distinguish Federated learning (or Federated Optimization) concern the massive system scale, restricted communication and most importantly, the non-Independent and Identically Distributed (non-IID) data assumption [9].
-Although extensive research has been conducted to enhance the efficacy and efficiency of FL in diverse scenarios, existing methods hardly go beyond model-based knowledge sharing where models [^3] Models and model parameters are used interchangeably if no ambiguity is caused., model updates, or gradients are exchanged between different parties as the carrier of knowledge.
-This helps in protecting the raw data but also results in several deficiencies:
+Statistics show that the total volume of data or information created, captured, copied and consumed globally has experienced a tenfold growth in the past decade and is expected to reach 394 zetabytes in 2028 [statista2028].
+A research predicted that a single autonomous vehicle can generate as much as 10 gigabytes data per second in the future [yang2020vehicle].
+This signifies two trends in the development of Machine Learning (ML) and data mining techniques.
+On the one hand, vast data lays the foundation for large-scale learning systems [guo2024bls, guo2025deepseek, zhong2024evaluation].
+On the other hand, knowledge sharing across multiple data sources matters more than ever [zhao2024domain, guan2025effective].
+In reality, the accessibility of private data is in many cases limited by legislation and regulations such as the General Data Protection Regulation (GDPR) [^1] https://gdpr-info.eu/ and the California Consumer Privacy Act (CCPA) [^2] https://www.oag.ca.gov/privacy/ccpa.
+This to a great extent leads to a paradigm shift from centralized learning to collaborative training over decentralized data.
+The most popular solution to “data islands” is Federated learning (FL) [mcmahan2016arxiv, mcmahan2017communication], which was developed on the basis of Federated Optimization for collaborative model training [kairouz2021fl].
+Although extensive researches have been conducted following the FL paradigm, they hardly go beyond learning-centric, task-specific knowledge sharing.
+This protects the raw data but results in several deficiencies:
   • Intractable fairness: No matter the participants cooperate towards a consensus model for common welfare or personalized models for individual interest, practical fairness is hard to realize.
@@
 They are generally in the same tensor structure and can take up to gigabytes of network traffic for each transmission.
-Besides, the cost for local training cannot be overlooked considering the problem of slow convergence over non-IID data [10].
-  • Poor reproducibility: models are the carriers of knowledge but are hard to reproduce in massive collaboration due to the randomness in participation.
+Besides, the cost for local training cannot be overlooked considering the problem of slow convergence over non-IID data [amiri2022conv].
+  • Poor reproducibility: models can carry knowledge but are hard to reproduce in massive collaboration due to the randomness in participation.
 Any participant has to request massive cooperation in order to reproduce a model from scratch.
-Is it possible to find an alternative to model-centric collaboration and tackle the above-mentioned issues?
-In this paper we offer our answer by presenting a knowledge-centric collaboration paradigm termed Knowledge Augmentation in Federation (KAF) or Federated Knowledge Augmentation.
-The key idea behind KAF is to redefine the way that participants collaborate in a federated system by shifting the focus from learning models to enhancing knowledge across the domains, as illustrated in Fig.
+Figure 1: The conceptual overview of Task-Agnostic Federation (TAF), which is a data-centric scenario that liberates the participants from any common goal and employs flexible protocols for generic knowledge exchange.
+Is it possible to find an alternative to task-oriented collaboration and tackle the above-mentioned issues?
+In this paper we offer our answer by describing a new scenario termed Task-Agnostic Federation (TAF), under which we gather emerging studies that enable new ways of collaboration through the exchange of generic knowledge in data.
+TAF liberates the participants by allowing for diverse motivation of collaboration in a federated system.
+This is done by shifting the focus from learning models to enhancing data across the domains, as illustrated in Fig.
 1.
-Technically, the key difference that KAF brings to the field mainly concerns i) the introduction of knowledge-centric objectives for collaborative optimization in the federated system, and ii) the adoption of more flexible carriers of knowledge without compromising privacy or proprietary rights.
-More details will be elaborated in Sec.
+This generalizes the benefits of collaboration because the enhanced data can take various forms and serve for any downstream purposes beyond learning.
+The concept of TAF will be elaborated in Sec.
 III.
-On the one hand, typical distributed learning paradigms seek to build a strong model by aggregating models, gradients or other forms of knowledge from separate domains.
-Even in heterogeneity-aware approaches such as personalized Federated Learning (pFL) [11], the optimization is still mostly limited to task-specific models which are hardly generalizable.
-By contrast, Knowledge Augmentation in Federation emphasizes the enhancement of local knowledge associated with each individual participant.
-On the other hand, The knowledge brought in can take whatever form, e.g, models, gradients, predictions, synthetic data, etc., as long as it serves to enrich or refine the knowledge implied in local data (Fig.
+Traditional distributed learning paradigms seek to build a strong model by aggregating models, gradients or other forms of knowledge from separate domains.
+Even in heterogeneity-aware approaches such as personalized Federated Learning (pFL) and Federated Multi-Task Learning (FMTL) [rahman2024tvt, liu2025pfed, skovajsova2025fmtl], the optimization is still mostly limited to task-specific models which are hardly generalizable.
+By contrast, Task-Agnostic Federation emphasizes the enhancement of private data with generic knowledge across the system.
+Particularly, The knowledge brought in can take whatever form, e.g, statistical information, synthetic data, generators, class prototypes, predictions, etc., as long as it serves to enrich or refine the knowledge implied in local data (Fig.
 2).
-This calls for new protocols that enable cost-effective exchange of knowledge.
-We will elaborate on these points in Sec.
-III.
-Compared to existing paradigms, KAF has potential advantages summarized as follows:
-  1. 1.
-    Improving global knowledge coherence across non-IID data holders,
-  2. 2.
-    Enhancing local knowledge representations in quantity, quality and utility for better local training or independent learning, and
-  3. 3.
-    Knowledge-centric collaboration for augmenting local domains with learning-friendly data rather than task-specific models.
-Figure 2: The conceptual relationship between data, models and knowledge.
-Moreover, we believe knowledge augmentation in federation shows a new roadmap for autonomic intelligence, where each participant can learn independently from knowledge gained collaboratively.
-In fact, there exist a number of studies that adopted inspiring methodologies in relevance to KAF, but a comprehensive review is so far missing in the literature.
-As far as we know, this review is the first to assemble related studies in the field under the concept of KAF.
-I-A Contributions and Related Surveys
-Extensive research has been carried out in the fields of deep learning (DL), representation learning, and data engineering.
-Meanwhile, demands for data privacy have given rise to broad interest in collaborative learning paradigms such as federated learning, swarm learning and fog learning.
-In the following we briefly review the representative surveys under these topics and then highlight the difference made in our work.
-Deep learning and representation learning.
-There exist a number of comprehensive surveys on DL-related techniques and applications.
-Dong et al.
-[12] briefly reviewed the development of Deep Neural Networks (DNNs) as well as other fundamental elements for DL such as activation functions, model architectures and programming libraries.
-Similarly, Sarker [13] categorized typical DL approaches in a taxonomy based on the availability of labels.
-Realistic tasks for supervised learning, unsupervised learning, hybrid learning and other learning paradigms are systematically reviewed.
-From a different angle, Alzubaidi et al.
-[14] paid attention to the impact of data scarcity on DL model training and how to deal with it.
-Besides, how to learn strong representations from the data is believed to be the key to DL algorithm design [15].
-Considering the connection between data representations with domain-specific knowledge, surveys have been conducted with focuses on various domains such as network representation learning [16], multi-modal representation learning [17], and graph representation learning [18].
-Data engineering and management.
-The development of Big Data techniques not only paves the way for complex learning systems, but also calls for more learning-centric data engineering.
-Generally speaking, data engineering entails the whole lifecycle of data management from acquisition and pre-processing to utilization, each with significant impact on the outcome of learning [19].
-Among the problems in the field, data imbalance has attracted the most attention, and a diversity of pre-processing and learning-based methods were investigated in the literature [20, 21].
-In addition, surveys such as [22, 23, 24] shed light on the bias and fairness in data.
-However, collaborative engineering of knowledge beyond data is yet to be covered in existing surveys.
+This calls for new protocols that enable cost-effective exchange of knowledge under privacy constraints.
+Compared to existing paradigms, TAF has advantages in several aspects.
+First, it directly enhances local data in quantity, quality or utility.
+Second, it brings generic, reusable knowledge to users that can be utilized later for diverse custom purposes.
+Besides, it leads to sustainable gain for every participating party despite how much they are engaged [^3] By contrast, FL relies on persistent client engagement to produce strong models..
+This is practically important for the participants who have very limited resources to invest.
+In fact, an increasing number of studies have adopted inspiring methodologies in relevance to TAF, but a comprehensive review is so far missing in the literature.
+As far as we know, this review is the first to assemble related studies in the field under this concept.
+Our contributions.
+In this paper we advocate a data-centric vision and focus on how multiple parties can cooperate without a common goal in a federated system.
+We believe TAF distinguishes itself as a new roadmap for autonomic intelligence, where each participant seeks and shares knowledge that is generic and reusable.
+The value of this article is summarized as follows:
+  • A revisit to multi-party collaboration across “data islands” from the standpoint of task-irrelevant interest and long-lasting benefits.
+  • New understanding of collaboration as a bridge to augmenting, refining and transforming local data for general purposes rather than learning.
+  • Literature review and insights for what TAF can bring up for the new forms of federated systems where autonomy matters.
+TABLE I: The comparison between our work and the surveys in related fields regarding the research themes.
+Theme Surveys Task-agnostic Collaboration Learning Local benefit
+Data engineering & knowledge management (e.g., [balayn2021biasunfairness, cui2025kadl, xie2024dataquality, gonzales2023synth, caton2024fairness, mehrabi2022biasfairness])
+Federated learning and variants (e.g., [kairouz2021fl, chen2024flprivacy, martinez2023decenfl, sabah2024model, ye2023heterofl, ji2024fxl, skovajsova2025fmtl]) ✓(pFL&FMTL)
+Federated analytics (e.g., [wang2025fa, wang2022fa, elkordy2023federated])
+Ours
+Figure 2: Data is the raw material and the most flexible form of knowledge. Information presented in statistics, patterns and models is derived from data and carries a specific subset of knowledge.
+II Background and Related Surveys
+In this section we briefly review some closely related concepts and draw comparison with related surveys to highlight our unique perspectives.
+Data engineering and knowledge management.
+Data engineering is something usually “behind the screen” but vital to any learning or mining pipeline [balayn2021biasunfairness], whilst the management of knowledge is pervasive in practice spanning from data abstraction to expertise transfer.
+The rise of these topics dates back to decades ago when rule-based symbolism and expert systems prevailed.
+An early work by Studer et al.
+[studer1998knowledge] signified a shift towards modeling view and outlined two diverged approaches to refining knowledge, whilst a recent survey by Cui et al.
+[cui2025kadl] reviews how deep learning can be augmented by domain knowledge.
+A diversity of data-centric studies were also investigated in the literature, with topics covering data balancing [xie2024dataquality, gonzales2023synth, chawla2002smote], feature engineering [chandrashekar2014survey, salau2019extract] as well as data bias and fairness issues [caton2024fairness, balayn2021biasunfairness, mehrabi2022biasfairness].
+However, new scenarios like collaborative data engineering and decentralized knowledge augmentation is yet to be covered in existing surveys.
 Federated learning and its variants.
+Federated learning is a distributed machine learning paradigm for privacy-preserving collaboration across a multitude of devices [bonawitz2019flsys].
+The most unique challenge for FL is the multi-party ownership of data.
 Kairouz et al.
-[9] presented one of most comprehensive surveys on FL, covering the efficiency, effectiveness, privacy, robustness and fairness issues.
-They also highlighted the challenges related to ill-distributed data for both cross-device and cross-silo FL scenarios.
-Some surveys focus on particular aspects of an FL system such as privacy preservation [25], decentralization [26], and heterogeneous FL [27].
-Application-specific studies are also reviewed in the literature, e.g., FL for Internet of Things (IoT) [28], smart cities [29], and healthcare [30, 31, 32].
+[kairouz2021fl] presented one of the most comprehensive surveys on FL, covering the efficiency, effectiveness, privacy, robustness and fairness issues.
+Some surveys focus on particular aspects of an FL system such as privacy preservation [chen2024flprivacy], decentralization [martinez2023decenfl], personalized FL (pFL) [sabah2024model], and heterogeneous FL [ye2023heterofl].
+Application-specific studies are also reviewed in the literature, e.g., FL for Internet of Things (IoT) [nguyen2021fliot], smart cities [pandya2023flcities], and healthcare [taha2023flhealthcare].
 Ji et al.
-[33] termed the variants of FL paradigms as “Federated X Learning”, which represents the fusion of FL with other emerging ML paradigms.
-Nonetheless, these surveys, without any exception, are learning-centric and did not investigate what FL can bring back to the data.
+[ji2024fxl] termed the variants of FL paradigms as “Federated X Learning”, which represents the fusion of FL with other emerging ML paradigms.
+Nonetheless, these surveys, without any exception, overlooked what learning can bring back to the data.
+Federated Analytics.
+Federated analytics (FA) [googleai2020federated], usually developed on information seeking, confidential computing and cryptography, aims to obtain global statistics with minimal private information leakage.
+The motivation and unique challenges for FA are discussed in [wang2022fa], while Wang et al.
+[wang2025fa] provide a comprehensive view of research landscape, with a taxonomy that entails analytical tasks, algorithms as well as privatization.
+Elkordy et al.
+[elkordy2023federated] review relevant techniques to FA from a query-oriented perspective.
+Yet, FA does not bring knowledge back to the data owners, making what we are interested in uncovered in these surveys.
+In contrast to existing surveys, we advocate a task-agnostic vision and investigate how multiple parties in a federated system can benefit from data-centric cooperation.
 Table I highlights the aspects that distinguish this review paper from the rest in the literature.
-TABLE I: The comparison between our work and the surveys in related fields regarding the research themes.
-Knowledge-centric collaboration learning data & feature engineering
-DL & representation learning (e.g., [12, 13, 14, 15, 16, 17, 18])
-Data engineering & management (e.g., [19, 20, 21, 22, 23, 24])
-Federated learning (e.g., [9, 25, 26, 27, 28, 29, 30, 31, 32, 33])
-Ours
-Our contributions.
-In this paper we advocate a knowledge-centric vision and focus on what learning can bring back to data in a federated system.
-Main contributions of this work are summarized as follows:
-  • A revisit to data processing and learning across “data islands” from the perspective of knowledge.
-  • New understanding of collaborative learning as a bridge to knowledge augmentation in federation.
-  • Insights for what KAF can bring up for the evolution of AI and the new regimes of collaborative intelligence.
-II Background and Preliminaries
-In this section we provide a brief introduction to the basics of federated learning and data engineering, which helps to better understand the ideas and topics covered in our work.
-II-A Federated Learning
-Federated learning is a distributed machine learning paradigm for privacy-preserving collaboration across a multitude of devices.
-The basic system model of FL includes user devices (often termed clients) that possess private data and a central server that coordinates the learning process [34].
-The standard process of FL [7] is organized in rounds each comprised of the following steps: i) Client selection: a fraction of clients are selected for the upcoming round of training; ii) Model distribution: the server synchronizes with the selected clients by sending them the global model (consensus model); iii) Local training: each selected client trains the model received on its local data using gradient-based methods, which produces a local model; iv) Model upload: selected clients upload their local models to the server; and v) Model aggregation: the server aggregates the local models into a new global model.
-These steps cycle for a number of rounds until some criterion is met.
-A unique challenge that distinguishes federated learning from other learning paradigms is the multi-party ownership of data.
-This significantly complicates the way we operate across multiple disjoint sets of data, i.e., “data islands”.
-On the one hand, in the FL setting we need to deal with a second layer of data distribution, that is, a distribution of similar but different domains represented by all the client-owned local datasets in the system.
-On the other hand, we still face the notorious problems, such as data imbalance, sample corruption and label noise, probably within each of the local datasets.
-More importantly, the orchestrator (e.g., a central server in most cast) is typically agnostic of local data distribution due to privacy constraints.
-II-B Data Engineering
-Data engineering in a general sense refers to the practice in the management of data spanning its whole lifecycle [23].
-Data engineering mainly deals with the problems in data collection and integration, data storage, data processing, and data analytics [^4] Systematic aspects such as data pipeline automation are also elements of the theme but do not fall within the scope of this paper..
-The boom of ML and AI applications has brought increasing attention to learning-centric data engineering methods particularly including data augmentation, data balancing and feature engineering.
-In the following we briefly explain these concepts.
-Data augmentation and balancing.
-The goal of data augmentation is to facilitate learning or analytics by artificially enhancing a dataset in size and sample diversity.
-Widely adopted in image, text and rational data pre-processing, data augmentation usually involves the generation of synthetic samples and the transformation of raw data.
-On the one hand, data synthesis introduces artificial samples to the dataset by exploiting interpolation, extrapolation, statistical models, generative models, or any combination of these techniques [35].
-Note that data synthesis is also a common strategy for data balancing or re-sampling where the distribution of samples is amended by over-sampling the minority class.
-This is the very rationale behind the classic methods such as Synthetic Minority Oversampling Technique (SMOTE) [36], ADAptive SYNthetic (ADASYN) [37], and Mixup [38].
-On the other hand, data transformation helps to diversify the features of data by applying certain types of transforms.
-The transformation of data are usually domain-specific, e.g., rotation, flipping and rescaling for image data, and random deletion, insertion and rephrasing for text data.
-Feature Engineering (FE).
-Raw data are more often than not represented by attributes that can be noisy, highly correlated and even corrupted.
-The goal of feature engineering is to create a new set of features that better represent the knowledge implied and filter out the noise and misinformation.
-Typical recipes for feature engineering includes feature selection and feature transformation and extraction.
-Specifically, feature selection reduces the dimensionality of the feature space by filtering for the most important features, which is usually performed based on statistical evidence or measures from information theory [39].
-As for feature transformation and extraction, it not only applies numerical operation(s) like rescaling to individual features, but also involves re-encoding the data with new combinations of features [40].
-Considering the target value (label) as a special feature, methods such as label correction and generation can also be categorized as the implementation of FE.
-In practice, domain knowledge can be leveraged in the FE process.
-The confluence of feature engineering and representation learning unleashes the power of AI over massive data.
 III Conceptual Overview
-In this section we provide a holistic view of the proposed KAF paradigm.
-We start with introducing the (suggested) system architecture and move on to the problem setting.
+In this section we provide a holistic view of the proposed scenario of TAF, starting with a general system architecture and moving on to the problem setting.
 This is followed by the outline of key approaches that will be elaborated in Secs.
@@
 As shown in Fig.
-1, the system model of KAF is similar to that of (horizontal) federated learning.
-To be specific, a central server is deployed as the coordinator at the center of the hub-and-spoke topology.
-A multitude of participants $\mathcal{K}=\{u_{1},u_{2},\ldots\}$ connect to the server to send and receive shared knowledge.
-Each participant $u_{k}$ has the exclusive access to its local data $D_{k}$ that is not allowed to be shared for privacy or proprietary reasons.
-A key difference is that each participant in KAF is not necessarily capable of training deep ML models.
-This is because the optimization objective for KAF is not learning-centric but knowledge-centric, which will be elaborated in the following section.
-III-B Problem Statement
-The fundamental problem behind privacy-sensitive collaboration is the “data islands” that causes non-IID data across the participants (i.e., clients in FL).
-From the view of knowledge, this problem can be interpreted as a global domain segmented into multiple overlapped sub-domains [^5] Also simply referred to as (local) domains if no ambiguity is caused.
-corresponding to the local data partitions on the clients.
-The proposed paradigm, namely knowledge augmentation in federation, seeks to address the problem by exchanging and enhancing the knowledge in private data independent on task-specific model learning.
-In other words, the primary goal of KAF is to maximize the effective capacity of knowledge for each participant [^6] The terms participant and client are used interchangeably.
-$u_{k}\in\mathcal{K}$ that holds a private dataset $D_{k}$, where $\mathcal{K}$ represents the entire population in the system.
-In addition, we also expect each local domain to align with the global domain through collaboration but without deviating too much from its original data distribution $\mathcal{D}_{k}$.
-This translates to a prototypical objective function expressed in (1),
-$\max_{\pi_{k}}F_{k}=\underbrace{\nu\big{(}\Theta_{\pi_{k}}(D_{k};\mathcal{K})\big{)}}_{enhancement}-\\ \alpha_{k}\Big{(}\beta_{k}\cdot\underbrace{\sum_{u_{i}\in\mathcal{K}/u_{k}}d\big{(}\Theta_{\pi_{k}}(D_{k};\mathcal{K}),\Theta_{\pi_{k}}(D_{i};\mathcal{K})\big{)}}_{alignment\;loss}+\\ (1-\beta_{k})\cdot\underbrace{d\big{(}\Theta_{\pi_{k}}(D_{k};\mathcal{K}),D_{k}\big{)}}_{regularization}\Big{)},$ (1)
-where the function $\nu(\cdot)$ evaluates the capacity of knowledge given a dataset, $\Theta_{\pi}(\cdot;\mathcal{K})$ denotes the operations following a strategy (or protocol) $\pi$ on local data collaboratively performed across the federation $\mathcal{K}$, and $d(\cdot,\cdot)$ measures the distributional distance (e.g., Kullback-Leibler (KL) divergence) between two datasets.
-The two client-specific hyper-parameters $\alpha_{k}$ and $\beta_{k}$ weight the three terms in the objective.
-Collectively, one can derive a system-wide objective from (1) if all the participants are cooperative and stick to their strategies $\{\pi_{1},\pi_{2},\ldots\}$:
-$\max_{\{\pi_{1},\pi_{2},\ldots\}}\sum_{k=1}^{|\mathcal{K}|}F_{k}$ (2)
-Besides, objective (1) indicates that each participant in KAF is self-motivated and contributes towards its own benefit.
-Particularly, the augmented knowledge represented in local data makes it possible for each $u_{k}$ to carry out independent model training or any other downstream tasks, making the cooperation knowledge-centric rather than learning-centric.
-On this point, KAF also differs from the objective of pFL [41] where each client aims to minimize a task-specific loss w.r.t.
-a specific model.
+1, the system model in TAF is similar to that of (horizontal) federated learning.
+Considering the diversity of data sources, a central server is usually needed for coordinating the scope and protocol of cooperation.
+On this hub-and-spoke topology, a multitude of participants $\mathcal{U}=\{u_{1},u_{2},\ldots\}$ connect to the server to send and receive shareable knowledge.
+Theoretically, TAF can also employ Peer-to-Peer (P2P) decentralized topology if the community is self-organized.
+In both cases, each participant $u_{k}$ has the exclusive access to its local data $D_{k}$ that is not allowed to be shared for privacy or proprietary reasons.
+Meanwhile, a certain knowledge exchange protocol has to be established between the participating parties in the first place.
+Remark.
+A key feature of TAF is that any participant is not necessarily capable of or interested in training.
+This is because generic knowledge does not always come from model training.
+For example, participants can swap data statistics, feature rankings, or synthetic samples for their specific purposes.
+III-B Formal Definition
+The fundamental setting of TAF determines that it is established over “data islands”.
+This means that any participant $u_{k}$ in the federation $\mathcal{U}$ is associated with a private dataset $D_{k}$.
+Each participant $u_{k}$ exchanges messages with the coordinator by a customized protocol $\pi_{k}$, which determines what it needs to send and receive and how.
+The formal definition is given as follows:
+Definition 1 (Task-Agnostic Federation (TAF)).
+Involving a set of autonomic participants $\mathcal{U}$ each with a private dataset $D_{k}$, TAF is a collaborative system or scenario where each participant $u_{k}$ seeks to maximize its data-related gain $g_{k}(D_{k},\mathcal{U},\pi_{k})$, where $g_{k}$ is a participant-specific function reflecting how much $D_{k}$ is enhanced.
+Each participant operates by a custom protocol $\pi_{k}$ without knowing any other participant’s objective.
+Example.
+To make the objective more understandable, here we sketch a mock algorithm.
+First, we use the effective number of samples proposed in [cui2019cbloss] to define $\nu(D)\triangleq(1-\beta^{|D|})/(1-\beta)$, with $\beta$ set as a priori (e.g., $\beta=0.999$).
+Next, each participant performs data clustering, and reports its centroids to the server in exchange for other participants’ centroids following the protocol $\pi_{k}$ [^4] We are not concerned about privacy leakage in this toy example..
+The shared centroids are then used for some local data augmentation $\Theta_{\mathcal{U},\pi_{k}}:D_{k}\rightarrow D^{\prime}_{k}$ (e.g., interpolation), which translates to $u_{k}$’s gain expressed as $g_{k}(D_{k},\mathcal{U},\pi_{k})=\Delta\nu=\nu(D^{\prime}_{k})-\nu(D_{k})$.
+The gain increases as the algorithm iterates.
+Remark.
+The gain implies that each participant in TAF is self-motivated and remains agnostic of any task (including its own and other participants’).
+On this point, TAF differs from pFL and FMTL as they aim to minimize a set of task-specific losses w.r.t.
+a set of specific models.
+Further, TAF brings sustainable benefit to the participants since their local data, once enhanced, can serve any downstream purposes beyond training.
+Figure 3: (Left) Taxonomy of research related to task-agnostic federation, and (right) the couriers of knowledge employed in the studies.
 III-C Outline of Approaches
-Through in-depth investigation, we figure out that there has been early effort in related areas especially in the conjunction of FL and data engineering.
-Even though the motivation behind these studies do not explicitly align with the goal of KAF, we found their methodologies somewhat suitable for serving as the building blocks for KAF.
-By categorizing these studies, we present a suggested roadmap for KAF comprised of three approaches: federated knowledge expanding, federated knowledge filtering, and federated label and feature space correction (Fig.
-3.
-Figure 3: The structure of this article.
-IV Methods for Knowledge Expansion
-As discussed previously, a federated system entails a set of knowledge domains that collectively comprise the global domain.
-Intuitively, it is easy to understand that greater overlaps between the local domains can facilitate faster and better learning.
-For example, it is commonly agreed that the global model’s convergence can be improved by reducing the “non-IIDness” of local data.
-In fact, this can be achieved by outspreading the knowledge contained in each local domain.
-In this section, we investigate popular approaches for collaborative knowledge expansion under the FL setting, where the implications of expansion are two-fold.
-First, this represents a roadmap for extending the boundaries of local knowledge domains through introducing exogenous information from other clients.
-Second, the expansion of knowledge also refers to the augmentation of data with the aims of re-balancing local distribution and generating diversified instances.
-Hence, the key criteria for this research category are:
-  • Seeking to extend local domains of knowledge through collaboration;
-  • Improving the diversity of data samples within each local domain based on information exchange.
+Through an in-depth investigation, we collect a rich set of studies that align or partially align with the characterized scenario of TAF.
+Some of these papers rest in the scope of federated learning but employ data-centric methods suitable for task-agnostic scenarios.
+We adopt a two-layer taxonomy: in the top layer we define three types of data-centric collaboration: collaborative data expansion, collaborative data refinement, and collective harmonization of data, and in the bottom layer we group the papers into multiple branches based on their technical design.
+Fig.
+3 visualizes the taxonomy of studies covered by this survey.
+IV Collaborative Data Expansion
+In this section, we investigate popular approaches for collaborative data expansion under the federated setting.
+By expansion we refer to methods that outspread the knowledge contained in each local domain by diversifying data samples.
 We summarize that there are two major lines of research in this direction.
-In the following content, we first introduce existing studies for extrapolating the knowledge in local data by information sharing and reconstruction in a federated system.
+In the following content, we first introduce existing studies for extrapolating the samples in local data by information sharing and reconstruction.
 Then we revisit the research on enhancing local data in terms of size, quality and distribution via data augmentation.
-IV-A Federated Knowledge Extrapolation
-We conceptually describe knowledge extrapolation as the methodology of filling the missing part of knowledge that is not explicitly expressed by the raw data.
+IV-A Federated Sample Extrapolation
+We conceptually describe sample extrapolation as the methodology of filling the missing part of information that is not explicitly expressed by the raw data.
 This process usually involves predicting or approximating implicit samples locally by utilizing some global knowledge contributed and shared by other participants in the system (Fig.
 4).
-In the context of federated learning, knowledge extrapolation can be very useful when all or most of the local datasets are sparse in terms of information density.
+In a federation, sample extrapolation can be very useful when all or most of the local datasets are sparse in terms of information density.
 For example, Sanyal et al.
-[42] considered a healthcare scenario where numerous Internet of Medical Things (IoMT) devices are simultaneously working towards sensing a particular phenomenon.
-In this case each IoMT device acts as an independent data collector that simultaneously generate time series data.
-The challenge here is that these readings need to be collectively utilized for analysis, but it is not encouraged to continuously push them to a central server due to privacy concerns and network constraints.
-To this end, the authors proposed a Federated Filtering Framework (FFF) [42], where the key component on each local device is a predictor for approximating the true sensing data.
+[Sanyal2019fff] considered a healthcare scenario where numerous Internet of Medical Things (IoMT) devices act simultaneously as independent data collectors.
+The challenge here is that their readings need to be collectively utilized for analysis, but it is not encouraged to continuously push them to a central server.
+To this end, the authors proposed a Federated Filtering Framework (FFF) [Sanyal2019fff], where the key component on each local device is a predictor for approximating the true sensing data.
 The IoMT devices communicate with a fog server that aggregates and utilizes the predictors in order to reconstruct the full data matrix.
-For communication efficiency, the timing of local model update and aggregation is determined by dynamic thresholds.
-Figure 4: A schematic showing the general framework of knowledge extrapolation. The goal is to obtain the missing part of knowledge in the local domain.
-Sparsity is one of the major obstacles that can largely undermine the efficacy of machine learning.
-The problem has been extensively explored in the context of recommender systems and the design of recommendation algorithms.
-More specifically, data are represented as a user-item interaction matrix that is huge in dimensionality but low in information density, which is caused by the natural pattern of user behavior and the scale of e-commerce platforms.
+Figure 4: A schematic showing the general framework of federated sample extrapolation. The goal is to obtain the missing part of knowledge in the local domain.
+Sparsity is one of the major obstacles that can largely undermine the efficacy of machine learning and data mining.
+The problem has been extensively explored in the context of recommender systems where the data are represented as a user-item interaction matrix that is huge in dimensionality but low in information density, which is caused by the natural pattern of user behavior and the scale of e-commerce platforms.
 Although there are many solutions to it in the literature, most of the existing studies make the assumption that the whole data matrix is available for analysis.
 However, a centralized collection of user behavior data may not be permitted in many situations such as cross-border e-commerce.
-As one of the most popular approaches to recommender systems, Collaborative Filtering (CF) an be viewed as a data model that is able to extrapolate incomplete knowledge (i.e., unrecorded interactions) using the dense representation of factor matrices.
+As one of the most popular approaches to recommender systems, Collaborative Filtering (CF) can be viewed as a data model that is able to extrapolate incomplete information (i.e., unrecorded interactions) using the dense representation of factor matrices.
 Approved by the Netflix Prize, Matrix Factorization (MF) is the classic implementation of CF, where the raw user-item interaction matrix $\boldsymbol{R}\in\mathbb{R}^{m\times n}$ can be approximated as
@@
 where $\boldsymbol{X}\in\mathbb{R}^{k\times m}$ and $\boldsymbol{Y}\in\mathbb{R}^{k\times n}$, with $m$ and $n$ being the number of users and items, respectively.
-By adapting MF to FL settings, researchers have found new paths to personalized recommendation without a centralized storage of user vectors (which are likely to contain user-private information).
+It is hard to predict unknown interactions (i.e., new knowledge) without a centralized storage of user vectors (which are likely to contain user-private information).
 Ammad-ud-din et al.
-[43] first described this new approach as Federated Collaborative Filtering (FCF), where all the clients keep their interaction data private on their own devices.
-Based on a reconstruction loss and stochastic gradient descent (SGD), clients in FCF update their user-specific (user-factor) matrices locally given the latest item-factor matrix from the server.
-In the meantime, each client also calculates its corresponding part of the item-specific gradients and feeds back to the server for item-factor matrix update.
+[FCF2019] first described this new model as Federated Collaborative Filtering (FCF), where all the participants keep their interaction data private.
+Based on a reconstruction loss and stochastic gradient descent (SGD), participants in FCF update their user-specific (user-factor) matrices locally given the latest item-factor matrix from the server.
+In the meantime, each participant also calculates its corresponding part of the item-specific gradients and feeds back to the server for item-factor matrix update.
 Following a similar framework, Yang et al.
-[44] further investigated a multi-behavior scenario, where both item viewing and purchase are recorded but the former is kept private by each user.
+[yang2024discrete] further investigated a multi-behavior scenario, where both item viewing and purchase are recorded but the former is kept private by each user.
 The authors proposed to binarize both factor matrices and utilized a caching mechanism to reduce communication payload for collaborative parameter update.
-Considering that a single form of information may not be sufficient for accurate recommendation, the knowledge provided by implicit and explicit feedback can be integrated using a probabilistic model.
-In this case, a specific decomposing strategy is needed for model updating without exposing users’ private information [45].
-Neural Collaborative Filtering (NCF) [46] replaces the matrix factorization process of CF by using deep neural networks (DNNs) to learn the user embedding vectors $\boldsymbol{X}=\{\boldsymbol{x_{1}},\boldsymbol{x_{2}},\ldots\}$ and item embedding vectors $\boldsymbol{Y}=\{\boldsymbol{y_{1}},\boldsymbol{y_{2}},\ldots\}$.
-With an embedding dimensionality of $k$, these learned latent vectors are then concatenated and fed into a Multi-Layer Perceptron (MLP) that outputs the predicted score of preference.
-NCF differs from MF-based approaches in the sense that knowledge is encapsulated in deep neural networks (DNNs), rather than factor matrices.
-In such way each user and item can be encoded independently and fed into a prediction model (e.g., another DNN) for training and inference.
-Inspired by this property, Perifanis and Efraimidis [47] designed a federated NCF framework that extends that basic NeuMF model where each client holds three sets of parameters including their user profile, item profile, and the DNN’s weights.
+Considering that a single form of information may not be sufficient for accurate recommendation, the information provided by implicit and explicit feedback can be integrated using a probabilistic model [yang2022dpmf].
+Neural Collaborative Filtering (NCF) [NCF2017] replaces the matrix factorization process of CF by using deep neural networks (DNNs) to learn the user embedding vectors $\boldsymbol{X}=\{\boldsymbol{x_{1}},\boldsymbol{x_{2}},\ldots\}$ and item embedding vectors $\boldsymbol{Y}=\{\boldsymbol{y_{1}},\boldsymbol{y_{2}},\ldots\}$.
+NCF differs from MF-based approaches in the sense that knowledge is encapsulated in deep neural networks (DNNs) rather than factor matrices.
+Inspired by this property, Perifanis and Efraimidis [perifanis2022federated] designed a federated NCF framework that extends the basic NeuMF model, where each participant holds three sets of parameters including their user profile, item profile, and the DNN’s weights.
 With user profiles kept private, they proposed a secure aggregation algorithm, MF-SecAgg, to perform server-side parameter update, with separate steps for item profile update and DNN update respectively.
 With a similar motivation, Jiang et al.
-[48] introduced an adaptive Differential Privacy (DP) mechanism with decaying noise into the communication process of Federated NCF.
-This protects the exchange of model gradients against inference attacks.
-For effective knowledge extrapolation, it is intricate to strike a balance between preserving learned representations locally and leveraging the common knowledge shared globally.
+[JiangFedNCF2022] introduced an adaptive Differential Privacy (DP) mechanism with decaying noise into the communication process of Federated NCF.
+This protects the exchange of gradients against inference attacks.
+For effective sample extrapolation, it is intricate to strike a balance between preserving learned representations locally and leveraging the common knowledge shared globally.
 To address this, Li et al.
-[49] designed a gating dual-encoder structure based on Variational Auto-Encoder (VAE) for personalized CF in a federated setting.
+[li2024personalized] designed a gating dual-encoder structure based on Variational Auto-Encoder (VAE) for federated CF.
 They introduced a gating network followed by two encoders, one for pertaining local knowledge and the other for learning user representations in the global latent space.
@@
 This facilitates global knowledge fusion but also increases the burden on end devices.
-One may ask that is it possible to make local knowledge “invisible”?
+So is it possible to make local knowledge “invisible”?
 Singhal et al.
-[50] offered their answer by presenting a novel reconstruction-based framework FedRecon.
-Inspired by Model-Agnostic Meta Learning (MAML), FedRecon allows stateless participants to rebuild their local model every time they join the federation with the knowledge provided by the global model.
-This facilitates more flexible collaboration at scale and, more importantly, indicates that personalized knowledge can be extrapolated by leveraging globally shared knowledge.
+[singhal2021federated] offered their answer by allowing stateless participants to rebuild their local model every time they join the federation with the knowledge provided by the global model.
+This enables flexible collaboration at scale and, more importantly, indicates that personalized knowledge can be extrapolated by leveraging globally shared knowledge.
 IV-B Federated Data Augmentation
-Data augmentation, as a powerful technique for diversifying the data, can also enable effective expansion of knowledge when deployed in a federated system.
+Data augmentation is a powerful technique for diversifying the data when deployed in a federated system.
 The value of federated data augmentation is two-fold.
 On the one hand, it directly introduces new knowledge—typically represented by synthetic samples—to the set of local data.
-On the other hand, multiple data owners can collaboratively perform data augmentation under the guidance of a global coordinator, which brings higher consistency of data distribution across the parties in the federation.
-The implementations of augmentation generally involve applying transforms on existing samples or synthesizing new samples based on a priori.
-Note that the synthesis of data and the training of downstream models can be decoupled, hence a simple implementation is to perform data augmentation in the local pre-processing stage without interleaving with the following FL process.
+On the other hand, multiple participants can collaboratively perform data augmentation, which brings higher consistency of data distribution across the parties in the federation.
+Data augmentation generally involves applying transforms on existing samples or synthesizing new samples based on a priori.
+Note that the synthesis of data and the training of downstream models can be decoupled, hence a simple implementation is to perform data augmentation in the local pre-processing stage without interleaving with the following learning process.
 Kundu et al.
-[51] followed this approach and developed a federated monkeypox disease detection model on the basis of offline image augmentation.
+[Kundu2024monkeypox] followed this approach and developed a federated monkeypox disease detection model on the basis of offline image augmentation.
 Specifically, a Cycle GAN is pre-trained and deployed locally to generate additional samples to enhance local training.
 Similarly, Ullah et al.
-[52] developed an FL-based smart healthcare system, where they employed oversampling in order to address the class imbalance of chest X-ray imaging data.
-More specifically, SMOTE is adopted to mitigate the scarcity of minority-class samples such as those with “COVID-19” and “Viral” labels [53].
-One potential limit is that unreasonable samples could be generated without using domain-specific knowledge as constraints.
-Despite being easy-to-implement, offline data augmentation is agnostic of any global knowledge and thus may not serve well for the global model’s convergence.
+[Ullah2024xray] developed an FL-based smart healthcare system, where they employed offline oversampling in order to address the class imbalance of chest X-ray imaging data.
+More specifically, SMOTE is adopted to mitigate the scarcity of minority-class samples such as those with “COVID-19” and “Viral” labels [Chowdhury2020covid19].
+Despite being easy-to-implement, offline data augmentation is agnostic of any global knowledge and thus may not serve well for the downstream tasks.
 The potential limit of offline data augmentation can be explained by considering the boundary of knowledge for each domain, wherein offline data augmentation is essentially analogous to linear and non-linear interpolation within the domain.
-By contrast, through introducing exogenous information for local data augmentation (i.e., collaborative data augmentation), more generalizable knowledge can be obtained to expand the domain.
-This could lead to higher gains for the downstream tasks.
-TABLE II: Summarizing the studies related to federated knowledge expansion in terms of methodological highlights, drawbacks, datasets for experiments, and the form of knowledge for exchange.
+By contrast, through introducing exogenous information for local data augmentation (i.e., collaborative data augmentation), more generalizable knowledge can be obtained to expand the domain, leading to higher gains for the downstream tasks.
+Although the collaboration of different parties naturally provides a rich set of knowledge sources for data augmentation, how to share the knowledge without compromising user privacy remains the key challenge.
+Essentially, a participant needs to perceive the distribution of data outside its domain without raw data exchange.
+On this point, sharing a generator (i.e., a generative model that produces new samples) has proved to be an useful mechanism.
+Wen et al.
+[Wen2022CVAE] proposed a generative model-based federated data augmentation strategy termed FedDA.
+They deployed and collaboratively trained a Conditional AutoEncoder (CVAE) on each participant.
+Instead of sharing the whole generator, they extracted the decoder’s hidden-layer features as the shared knowledge and introduced a knowledge distillation loss to guide the training.
+This effectively reduces communication overheads.
+For privacy preservation, noise samples are mixed into each local dataset.
+Intuitively, one can adapt traditional augmentation, such as the popular MixUp [Zhang2018mixup] strategy, to the TAF setting.
+Given a loss criterion $\ell$ and a model $h_{\theta}$, naïve Mixup globally yields the following loss over any generated samples $(\boldsymbol{x}^{\prime},y^{\prime})$:
+$\ell_{rawmix}(h_{\theta}(\boldsymbol{x}^{\prime}),y^{\prime})=\ell\Big(h_{\theta}\big((1-\lambda)\boldsymbol{x}_{i}+\lambda\boldsymbol{x}_{j}\big),(1-\lambda)y_{i}+\lambda y_{j}\Big),$ (2)
+where the raw data from two participants $i$ and $j$ are involved, posing privacy issues.
+To achieve privacy-preserving augmentation, Yoon et al.
+[yoon2021fedmix] presented Mean Augmented Federated Learning (MAFL), a simple framework that enables participants to exchange the average of local samples (called “mashed” data) along with model updates.
+Further, through approximating the global MixUp loss (Eq.
+2) with Taylor expansion, they derived that the approximation can be done using local samples and the averaged samples from other participants (Eq.
+(3)):
+$\ell_{FedMix}=\mu\ell\big(h_{\theta}(\mu\boldsymbol{x}_{i}),y_{i}\big)+\lambda\ell\big(h_{\theta}(\mu\boldsymbol{x}_{i}),\bar{y}_{j}\big)+\lambda\frac{\partial\ell}{\partial\boldsymbol{x}}\bar{\boldsymbol{x}}_{j},$ (3)
+where $\mu=1-\lambda$, $\bar{\boldsymbol{x}}_{j}=\frac{1}{|\xi|}\sum_{j\in\xi}\boldsymbol{x}_{j}$ and $\bar{y}_{j}=\frac{1}{|\xi|}\sum_{j\in\xi}y_{j}$ represent the mean of features and labels over a private data batch $\xi$ from other clients, respectively.
+The derivative $\frac{\partial\ell}{\partial\boldsymbol{x}}$ is evaluated at $\boldsymbol{x}=\mu\boldsymbol{x}_{i}$ and $y=y_{i}$.
+This motivated them to devise the FedMix [yoon2021fedmix] algorithm that performs collaborative MixUp augmentation and training based on a decomposable loss function.
+Zhang et al.
+[Zhang2023fedm] further extended this framework for regression tasks using bilateral neighborhood expansion and evaluated it on various public datasets.
+Empirical results show that their algorithms yielded stronger performance than FedMix [yoon2021fedmix].
+A special case of federated data augmentation is the reconstruction of samples using knowledge from multiple domains, which has proved very useful for accelerated medical imaging [yan2024cross, ahmed2025fedgraphmri].
+To adapt to the disparity of imaging operators between different sites, Elmas et al.
+[elmas2023mri] deployed an unconditional GAN at each site to generate synthetic high-quality samples, with a private discriminator to fit site-specific knowledge and a globally shared generator for prior knowledge embedding.
+TABLE II: Summarizing the studies related to collaborative data expansion in terms of methodological highlights, drawbacks, datasets for experiments, and the form of knowledge for exchange.
 Refs. Category Knowledge carrier Highlights Drawbacks Datasets
-[42] Fed. knowledge extrapolation predictors & filters $\bullet$ remote data reconstruction via predictors
+[Sanyal2019fff] Fed. sample extrapolation predictors & filters $\bullet$ remote data reconstruction via predictors
 $\bullet$ dynamic communication frequency based on error monitoring
 $\bullet$ uniform filter parameters across all devices MHEALTH
-[43, 45] Fed.
-knowledge extrapolation factor matrices $\bullet$ decoupling private and shareable knowledge via matrix factorization $\bullet$ linear growth of communication cost with the raw matrix’s dimensions MovieLens, FilmTrust, Epinions, and in-house data
-[44] Fed. knowledge extrapolation binarized matrices $\bullet$ finer categorization of private knowledge
+[FCF2019, yang2022dpmf] Fed.
+sample extrapolation factor matrices $\bullet$ decoupling private and shareable knowledge via matrix factorization $\bullet$ linear growth of communication cost with the raw matrix’s dimensions MovieLens, FilmTrust, Epinions, and in-house data
+[yang2024discrete] Fed. sample extrapolation binarized matrices $\bullet$ finer categorization of private knowledge
 $\bullet$ binarization and caching
 $\bullet$ limited to CF applications JD, Tmall, User Behavior, and MovieLens
-[47, 48, 49] Fed. knowledge extrapolation latent vectors and prediction models $\bullet$ better protection of private knowledge
+[perifanis2022federated, JiangFedNCF2022, li2024personalized] Fed. sample extrapolation latent vectors and prediction models $\bullet$ better protection of private knowledge
 $\bullet$ dilution of knowledge in model aggregation due to sparsity
@@
 MovieLens, Lastfm, Foursquare NYC, Ta-Feng, Amazon-Instant-Video, and QB-article
-[50] Fed. knowledge extrapolation the global model $\bullet$ model-agnostic and stateless collaboration
+[singhal2021federated] Fed. sample extrapolation the global model $\bullet$ model-agnostic and stateless collaboration
 $\bullet$ generalizable to diverse tasks
 $\bullet$ lack of consideration of data heterogeneity MovieLens and Stack Overflow
-[51, 52] Federated data augmentation task-specific models $\bullet$ easy to implement
+[Kundu2024monkeypox, Ullah2024xray] Federated data augmentation task-specific models $\bullet$ easy to implement
 $\bullet$ low risk of privacy leakage
@@
 COVID-19 radiography, pox virus, and chest X-ray images
-[54] Federated data augmentation attention scores from generative models $\bullet$ CVAE for local augmentation
+[Wen2022CVAE] Federated data augmentation attention scores from generative models $\bullet$ CVAE for local augmentation
 $\bullet$ light-weight knowledge exchange
 $\bullet$ high local training cost Fashion-MNIST and CIFAR-10
-[55, 56] Federated data augmentation the average of local samples $\bullet$ controllable anonymity
+[yoon2021fedmix, Zhang2023fedm] Federated data augmentation the average of local samples $\bullet$ controllable anonymity
 $\bullet$ low communication costs
 $\bullet$ hard to generalize for being loss-dependent FEMNIST, FMNIST, CIFAR-10, CIFAR-100, CINIC-10, Airfoil, NO2, and UTKFace
-[57, 58] Federated data augmentation prediction model $\bullet$ Augmented images with cross-institute knowledge $\bullet$ tightly-coupled with the task fastMRI and IXI
-[59] Federated data augmentation shared generator $\bullet$ decoupled knowledge with decoupled models $\bullet$ imbalanced amount of knowledge for the generator and discriminator fastMRI, IXI, BRATS, and in-house data
-Although the collaboration of different clients naturally provides a rich set of knowledge sources for data augmentation, how to share the knowledge without compromising user privacy remains the key challenge.
-Essentially, a client needs to perceive the distribution of data outside its domain without raw data exchange.
-On this point, sharing a generator (i.e., a generative model that produces new samples) has proved to be an useful mechanism.
-Wen et al.
-[54] proposed a generative models-based federated data augmentation strategy (FedDA), where a Conditional AutoEncoder (CVAE) is deployed and collaboratively trained on each client to tackle non-IID data distribution.
-Instead of sharing the whole generator, they extracted the decoder’s hidden-layer features as the shared knowledge and introduced a knowledge distillation loss to guide the training.
-This effectively reduces communication overheads.
-For privacy preservation, noise samples are mixed into each local dataset.
-Intuitively, one can adapt traditional data augmentation, such as the popular MixUp [38] strategy, to the federated setting.
-However, the difficulty here is that raw samples need to be shared if we want to mix up local and global knowledge.
-Over the generated samples $(\boldsymbol{x}^{\prime},y^{\prime})$ this “global MixUp” strategy theoretically corresponds to minimizing the following loss:
-$\ell_{rawmix}(h_{\theta}(\boldsymbol{x}^{\prime}),y^{\prime})=\ell\Big{(}h_{\theta}\big{(}(1-\lambda)\boldsymbol{x}_{i}+\lambda\boldsymbol{x}_{j}\big{)},(1-\lambda)y_{i}+\lambda y_{j}\Big{)}$ (4)
-where $\ell$ and $h_{\theta}$ is the task-specific loss criterion and the model to be trained.
-To achieve privacy preservation, Yoon et al.
-[55] first presented Mean Augmented Federated Learning (MAFL), a simple framework that enables clients to exchange the average of local samples (called “mashed” data) along with model updates.
-Further, through approximating the global MixUp loss (Eq.
-4) with Taylor expansion, they derived that the approximation can be done using local samples and the averaged samples from other clients (Eq.
-(5)):
-$\ell_{FedMix}=\mu\ell\big{(}h_{\theta}(\mu\boldsymbol{x}_{i}),y_{i}\big{)}+\lambda\ell\big{(}h_{\theta}(\mu\boldsymbol{x}_{i}),\bar{y}_{j}\big{)}+\lambda\frac{\partial\ell}{\partial\boldsymbol{x}}\bar{\boldsymbol{x}}_{j},$ (5)
-where $\mu=1-\lambda$, $\bar{\boldsymbol{x}}_{j}=\frac{1}{|\xi|}\sum_{j\in\xi}\boldsymbol{x}_{j}$ and $\bar{y}_{j}=\frac{1}{|\xi|}\sum_{j\in\xi}y_{j}$ represent the mean of features and labels over a private data batch $\xi$ from other clients, respectively.
-The derivative $\frac{\partial\ell}{\partial\boldsymbol{x}}$ is evaluated at $\boldsymbol{x}=\mu\boldsymbol{x}_{i}$ and $y=y_{i}$.
-This motivated them to devise the FedMix [55] algorithm that performs collaborative MixUp augmentation and training based on a decomposable loss function.
-Zhang et al.
-[56] further extended this framework for regression tasks using bilateral neighborhood expansion and evaluated it on various public datasets.
-Empirical results show that their algorithms yielded stronger performance than FedMix [55].
-A special case of federated data augmentation is the reconstruction of samples using knowledge from multiple domains, which has proved very useful for accelerated medical imaging such as Magnetic Resonance Imaging (MRI) [57, 58].
-To adapt to the disparity of imaging operators between different sites, Elmas et al.
-[59] deployed an unconditional GAN at each site to generate synthetic high-quality samples, with a private discriminator to fit site-specific knowledge and a globally shared generator for prior knowledge embedding.
-A tailored reconstruction loss was used to guide the training of local generators.
+[yan2024cross, ahmed2025fedgraphmri] Federated data augmentation prediction model $\bullet$ Augmented images with cross-institute knowledge $\bullet$ tightly-coupled with the task fastMRI and IXI
+[elmas2023mri] Federated data augmentation shared generator $\bullet$ decoupled knowledge with decoupled models $\bullet$ imbalanced amount of knowledge for the generator and discriminator fastMRI, IXI, BRATS, and in-house data
 IV-C Summary
-In this section we revisit two representative lines of research that share the common idea of expanding user/client-specific domains of knowledge in a federation.
-Federated knowledge extrapolation mainly deals with data sparsity and missing information in local domains, whilst federated data augmentation focuses on mitigating non-IID local data distribution.
-The two approaches methodologically differ in the two aspects.
-  • The way new knowledge is introduced: knowledge extrapolation typically relies on representation learning to map multiple sources of information into one or several vector spaces, where sparse features are transformed into dense embeddings.
-These embeddings and the encoders act as the carriers in the process of knowledge aggregation and distribution.
-On this basis, predictions can be made to extend the local domain of knowledge, for example, by extrapolating unobserved interactions in the recommender system.
-By contrast, data augmentation can be performed independently to enhance the representations of knowledge, which is essentially analogous to linear interpolation [^7] In fact, it is exactly the idea behind the MixUp strategy..
-Through collaboration, knowledge from outside the domain can be introduced by information fusion at sample level.
+Federated sample extrapolation mainly deals with data sparsity and missing information in local domains, whilst federated data augmentation focuses on mitigating non-IID local data distribution.
+We summarize related works in Table II and compare them in the following aspects:
+  • The way new knowledge is introduced: sample extrapolation typically relies on representation learning to map multiple sources of information into one or several vector spaces, using embeddings and encoder as the knowledge carriers.
+By contrast, data augmentation is essentially analogous to interpolation which fuses information at sample level.
 This directly results in an expanded size of data and typically does not depend on any task-specific models.
-  • Motivation and objectives: the extrapolation of knowledge is usually coupled with the optimization of task-specific objectives such as the accuracy of item recommendations.
-For example, in federated collaborative filtering, knowledge from different users are implicitly fused in the process of the global item vector update and local user vector update, guided by a cost function that quantifies the reconstruction loss.
-Federated data augmentation, on the other hand, aims to regulate the local data distribution towards the IID setting.
-In most cases, augmentation is performed without an explicit objective of optimization.
-Each of the two approaches has certain limitations, as summarized in Table II.
-For collaborative knowledge extrapolation, domain-specific knowledge is embedded as factor matrices or vectors that are optimized under the supervision of known observations.
-This means that inaccurate information could be inferred and disseminated once the training samples are very noisy or manipulated by malicious clients.
-For collaborative data augmentation, neither interpolation-based or generative model-based schemes can theoretically guarantee the fidelity of synthetic samples.
-Certain mechanisms are necessary to constrain the process of augmentation against invalid patterns [60, 61].
-So far, this problem still needs further investigation in the federated settings.
-V Methods for Knowledge Filtering
-A federated system is usually characterized by a rich set of data sources which, on the one hand, facilitate a collaborative way of learning but, on the other hand, bring about knowledge redundancy as well as the risk of data poisoning [62, 63].
-The issue has become increasingly concerning with the advance of generative models that are able to fabricate adversarial samples.
-For example, a hyperdimensional computing (HDC) model is employed in [64] to craft poisoned samples that can bypass byzantine-robust defenses.
-This calls for more effective mechanisms for knowledge filtering.
-We in this paper define knowledge filtering as the techniques for excluding misinformation in form of improper data (and its derivatives [^8] Including but not limited to embeddings, models and gradients.) and distinguishing different regions of the domain in terms of value for learning.
-Knowledge filtering can involve engineering and algorithm design at both levels across and within the domains.
-This section reviews the emerging studies that seek to address the discrepancy of knowledge regarding its validity and utility for the downstream task.
-V-A Client-level Filtering
-Poisoned datasets can severely threaten the validity of knowledge, which in turn causes performance degradation of the federated model learned.
-Being targeted or untargeted, data poisoning attacks are generally assumed to be carried out by one or multiple participants.
-Therefore, a straightforward approach to defending against data poisoning is to identify and exclude these malicious clients.
-Although a variety of schemes have been developed for the generation of adversarial samples that are difficult to detect individually, they still inevitably change the distribution of data, for instance, with flipped labels or fake samples blended in.
+  • Motivation and objectives: the extrapolation of samples is sometimes coupled with the optimization of task-specific objectives such as the accuracy of item recommendations.
+Federated data augmentation, in many cases, can be performed in a task-agnostic fashion.
+  • Limitations: For federated extrapolation, inaccurate information could be disseminated once the training samples are very noisy or manipulated by malicious clients.
+For augmentation, it is hard to theoretically guarantee the fidelity of synthetic samples.
+V Collaborative Data Refinement
+Federation can bring information redundancy as well as the risk of data poisoning to the participants [cina2024machine, Wang2023temporal].
+We in this paper define data refinement as the techniques for excluding misinformation that exhibits knowledge incoherence.
+Figure 5: (Left) results of our experiments over 100 clients under four disparate conditions of data that involve IID, non-IID and low-quality on-device data.
+(Right) the FedProf framework for dataset-level filtering.
+Experiment details are described in [wu2023fedprof], with more observations included.
+V-A Dataset-level Filtering
+A coarse-grained approach to filtering out misleading knowledge is to identify and exclude the malicious participants (and their datasets).
+Although tailored adversarial samples are difficult to detect individually, they still inevitably change the distribution of data.
 Through empirical studies on label-flipping attacks, Khuu et al.
-[65] found that, with a single round of attack, a small subset of adversarial clients can result in dramatic drop in the model’s recall on the targeted class as well as notable decrease of overall accuracy.
-Inspired by the Shapley Additive Explanation (SHAP), they developed a client filtering algorithm using Support Vector Machines (SVMs) to classify clients based on four features derived from SHAP values.
-However, SHAP values are computation-intensive for each client as we need to calculate a score matrix of all image-label pairs.
-By contrast, detecting data poisoning from the perspective of model is low-cost and more easy-to-implement.
-Model-based filtering approaches are based on the fact that local training on poisoned data will result in model update that are likely to deviate from that of benign clients.
-For example, the FedZZ [66] framework detects untargeted data poisoning attack based on the cosine similarity between a client’s update vector and a reference zone-aggregated vector.
-The proposed zone-based detection scheme groups the clients into multiple zones and calibrates the zone set progressively by accuracy.
+[Khuu2024shap] found that, with a single round of attack, a small subset of adversarial participants can result in dramatic drop in the model’s performance.
+Inspired by the Shapley Additive Explanation (SHAP), they developed a dataset filtering algorithm using Support Vector Machines (SVMs) to classify participants based on four features derived from SHAP values.
+However, SHAP values are computation-intensive for each participant as we need to calculate a score matrix of all image-label pairs.
+By contrast, Model-based filtering approaches are low-cost and more easy-to-implement, based on the fact that poisoned model updates are likely to deviate from benign updates.
+For example, the FedZZ [kumar2024precision] framework detects poisoned domains based on the cosine similarity between a client’s update vector and a reference zone-aggregated vector.
+The proposed zone-based detection scheme groups the participants into multiple zones and calibrates the zone set progressively by accuracy.
 Similarly, Wei et al.
-[67] investigated the impact of targeted data poisoning attacks on local training by examining the eigenvalues of the covariances of gradient updates.
-Through theoretical analysis and empirical study, they proved that gradients from adversarial clients and honest clients form clearly separable clusters, which provides forensic evidence for accurate client filtering.
-Nonetheless, challenges remain as it was also observed that the timing of attacks greatly influences the effectiveness of statistic-based defense strategies [67].
-Apart from adversarial clients with malicious intention, low-quality information such as noise and irrelevant samples may exist in local data.
-In our previous work [68], we conducted empirical studies inspecting the impact of low-quality information on the process of collaborative learning (Fig.
+[Wei2024eigen] proposed to examine the eigenvalues of the covariance matrix of gradient updates.
+Through theoretical analysis and empirical study, they proved that gradients from adversarial participants and honest participants form clearly separable clusters, which provides forensic evidence for accurate dataset-level filtering.
+Li et al.
+[li2024mitigating] proposed a confidence-aware defense (CAD) strategy that is insensitive to the types of poisoning.
+By introducing the definition of confidence score and SuperLoss [castells2020superloss], the rationale behind CAD is that tampered data will cause a significant drop of confidence score in local training, based on which malicious participants can be detected.
+Apart from adversarial participants with malicious intention, low-quality information such as noise and irrelevant samples may exist in local data.
+In our previous work [wu2023fedprof], we empirically investigated the impact of low-quality information on the process of collaborative learning (left panel of Fig.
 5).
 The result demonstrates that poor data conditions significantly slow down the convergence whilst undermining the global model’s performance.
-Figure 5: Results of our experiments over 100 clients under four disparate conditions of data that involve IID, non-IID and low-quality on-device data.
-Experiment details are described in [68], with more observations included.
-Instead of dirty labels, in [68] we examined the performance drop caused by low-quality clients that trained on local data affected by various forms of noise.
-We discovered that the engagement of low-quality knowledge can be as detrimental as pathological data distribution, and more importantly, the existence of both led to the worst results of convergence and final accuracy.
-To address the problem, a novel client filtering strategy FedProf [68] is presented, where the latent representations from the global model’s hidden layer are extracted and compressed into client-specific profiles.
-Based on the Kullback–Leibler (KL) divergence between the profiles, low-quality clients are filtered out before each FL round probabilistically.
-Noise data filtering at client level can be practically imperative especially for Internet of Things (IoT) applications.
+We discovered that the engagement of low-quality knowledge can be as detrimental as pathological data distribution.
+To address the problem, a novel participant filtering strategy FedProf [wu2023fedprof] is presented (right panel of Fig.
+5).
+The key aspect of this strategy is termed “representation profiling” where the latent output from the global model’s hidden layer are extracted and compressed into participant-specific profiles.
+Based on the Kullback–Leibler (KL) divergence between the profiles, low-quality datasets are filtered out probabilistically before each FL round.
+Noise data filtering at dataset level can be practically imperative especially for Internet of Things (IoT) applications.
 This is because real-time sensing is inherently susceptible to environmental disturbance and hardware flaws.
@@
 Targeting at this problem, Lei et al.
-[69] presented a filtering framework OES-Fed that works on the vehicle level.
+[lei2022oes] presented a filtering framework OES-Fed that works on the vehicle level.
 The framework first identifies a group of reference vehicles (i.e., vehicles with high-quality data) based on model accuracy, and sets their model parameters as the initial centers for K-means clustering.
@@
 In addition, Kalman filters and exponential smoothing are employed to incorporate the historical changes of model parameters.
-This enables dynamic assessment of local knowledge in response to Continual Learning (CL) scenarios [70] where new data keep rolling in.
-V-B Sample-level Filtering
-Studies on sample-level filtering endeavor to improve the quality of local knowledge by differentiating each sample within the local datasets.
-In contrast to the coarse-grained approaches that either keep or filter out entire local data domains, sample-level filtering corresponds to a fine-grained roadmap for enhancing the gain of local training and in turn facilitating global model convergence.
-However, we face unique challenges as it is typically not allowed in FL to access the raw attributes of each individual instance of data.
-This problem is especially intractable in the healthcare scenarios where the Electronic Health Records (EHRs) of patients are highly sensitive and probably heterogeneous across different organizations in the federation.
-By heterogeneous it means that samples from different clients scatter in inconsistent feature spaces that correspond to different permutations of features.
-To address this, Thakur et al.
-[71] proposed a knowledge abstraction and filtering framework, where the core component is the client-specific filtering modules that can be cascaded.
-Each filtering module starts with a native encoder that map local data to a unified latent space and ends with a gating operator that uses a globally shared knowledge abstraction vector to filter local samples’ representations.
-The novelty of this framework boils down to the knowledge abstraction vector which is trained collaboratively to embed consensus knowledge and guides the filtering of low knowledge.
-Nonetheless, the mechanism cannot directly identify and filter out problematic samples.
-Motivated by the knowledge distillation paradigm [72, 73], Lu et al.
-[74] proposed to employ conditional generators on each client to generate proxy samples for ensemble distillation.
-To address the loss of fidelity in synthetic data, the authors connected a filter to the head of each local model whose outputs are then modulated based on the reference labels.
-This helps in refining local knowledge without the need for a public data pool.
-As a matter of fact, knowledge filtering can be performed simultaneously at client level and sample level.
-Adopting such a hybrid approach, Xu et al.
-[75] cast light on cloud-edge collaboration where “dirty” data on untrustworthy devices may significantly compromise the learning process.
-A novel framework called Safe is presented, which at the client level resorts to Alternating Direction Method of Multipliers (ADMM) for detecting suspicious clients that behave like data poisoning attackers.
-At sample level, data samples on these (potentially) malicious are then filtered based on K-means clustering and Euclidean distance.
-A shortcoming here is that the Euclidean distance in the raw feature space may not serve as a good criterion in the sample-level filtering in the presence of high dimensionality and sparsity.
-Moreover, Safe was only evaluated on the learning task of Support Vector machines (SVMs), which means that whether the framework can generalize to deep learning tasks remains questionable.
-TABLE III: Summarizing the studies related to federated knowledge filtering in terms of methodological highlights, drawbacks, datasets for experiments, and the form of knowledge for exchange.
+This enables dynamic assessment of local knowledge in response to Continual Learning (CL) scenarios [liang2025diffusion] where new data keep rolling in.
+TABLE III: Summarizing the studies related to collaborative data refinement in terms of methodological highlights, drawbacks, datasets for experiments, and the form of knowledge for exchange.
 Refs. Category Knowledge carrier Highlights Drawbacks Datasets
-[65] Client-level filtering derived features from SHAP $\bullet$ model-independency and statistical interpretability $\bullet$ high computational cost of SHAP MNIST and Fashion-MNIST
-[66, 67] Client-level filtering gradients or update vectors $\bullet$ scalable and easy-to-implement
+[Khuu2024shap] Dataset-level filtering derived features from SHAP $\bullet$ model-independency and statistical interpretability $\bullet$ high computational cost of SHAP MNIST and Fashion-MNIST
+[kumar2024precision, Wei2024eigen] Dataset-level filtering gradients or update vectors $\bullet$ scalable and easy-to-implement
 $\bullet$ effective against diverse types of data pollution
@@
 Fashion-MNIST, EMNIST, CIFAR10, and LFW
-[68] Client-level filtering statistically compressed representations $\bullet$ very low cost of communication
+[wu2023fedprof] Dataset-level filtering statistically compressed representations $\bullet$ very low cost of communication
 $\bullet$ stronger privacy protection
 $\bullet$ requirement for dense layers in the model structure GasTurbine, EMNIST, and CIFAR-10
-[69] Client-level filtering task-specific models $\bullet$ scalable and effective against diverse types of data pollution
+[lei2022oes] Dataset-level filtering task-specific models $\bullet$ scalable and effective against diverse types of data pollution
 $\bullet$ historical behavior included for smoothing
 $\bullet$ relatively high cost of communication MNIST, CIFAR-10, and Vehicle Classification dataset
-[71] Sample-level filtering knowledge abstraction vector and the prediction model $\bullet$ abstraction of global knowledge for local filtering
+[thakur2024knowledge] Sample-level filtering knowledge abstraction vector and the prediction model $\bullet$ abstraction of global knowledge for local filtering
 $\bullet$ localized filtering module adaptable to heterogeneous domains
@@
 Fashion-MNIST, SVHN, CIFAR-10, and Digits-5
-[75] Client+sample filtering model parameters $\bullet$ knowledge filtering at both levels using a multi-stage design
+[xu2022safe] Dataset+sample filtering model parameters $\bullet$ knowledge filtering at both levels using a multi-stage design
 $\bullet$ edge-cloud hierarchical system for stronger privacy
@@
 Iris and WineQuality
-[76] Fed. data re-weighting task-specific models $\bullet$ IWDS for achieving fast convergence without loss of accuracy $\bullet$ knowledge outside local domains unexploited CIFAR-10 and Fashion-MNIST
-[77] Fed. data re-sampling task-specific models $\bullet$ easy to implement with any classic re-sampling methods $\bullet$ knowledge outside local domains unexploited MIMIC-III
-[78] Fed. data re-weighting Gini coefficients and models $\bullet$ learnable strategy for client-specific re-weighting control
+[tang2021data] Federated data re-weighting task-specific models $\bullet$ IWDS for achieving fast convergence without loss of accuracy $\bullet$ knowledge outside local domains unexploited CIFAR-10 and Fashion-MNIST
+[chen2024implementing] Federated data re-sampling task-specific models $\bullet$ easy to implement with any classic re-sampling methods $\bullet$ knowledge outside local domains unexploited MIMIC-III
+[ouyang2024ddpg] Federated data re-weighting Gini coefficients and models $\bullet$ learnable strategy for client-specific re-weighting control
 $\bullet$ local Gini coefficients reflecting local distributional knowledge
@@
 MNIST, FMNIST, and CIFAR-10
+V-B Sample-level Filtering
+In contrast to the coarse-grained approaches that either keep or filter out entire local data domains, sample-level filtering identifies the deviants in each domain.
+However, it is typically not allowed in a federated setting to access the raw attributes of each individual data instances.
+This problem is especially intractable in the healthcare scenarios where the Electronic Health Records (EHRs) of patients are highly sensitive and probably heterogeneous across different organizations in the federation.
+By “heterogeneous” it means that samples from different clients scatter in inconsistent feature spaces that correspond to different permutations of features.
+To address this, Thakur et al.
+[thakur2024knowledge] proposed a knowledge abstraction and filtering framework, where the core component is the participant-specific filtering modules that can be cascaded.
+Each filtering module starts with a native encoder that map local data to a unified latent space and ends with a gating operator that uses a globally shared knowledge abstraction vector to filter local samples’ representations.
+Nonetheless, the mechanism cannot directly identify and filter out problematic samples.
+Motivated by the knowledge distillation paradigm [le2025fedmekt, chen2023metafed], Lu et al.
+[lu2024datafree] proposed to employ local conditional generators to generate proxy samples for ensemble distillation.
+A filter is connected to the head of each local model whose outputs are then modulated based on the reference labels.
+This helps in refining local knowledge without the need for a public data pool.
+As a matter of fact, data refinement can be performed simultaneously at dataset level and sample level.
+Adopting such a hybrid approach, Xu et al.
+[xu2022safe] cast light on cloud-edge collaboration where “dirty” data on untrustworthy devices may significantly compromise the learning process.
+A novel framework called Safe is presented, which at the dataset level resorts to Alternating Direction Method of Multipliers (ADMM) for detecting suspicious participants that behave like data poisoning attackers.
+At sample level, data samples on these (potentially) malicious are then filtered based on K-means clustering and Euclidean distance.
+A shortcoming here is that the Euclidean distance in the raw feature space may not serve as a good criterion due to high dimensionality and sparsity.
+Moreover, Safe was only evaluated on the learning task of Support Vector machines (SVMs), which means that whether the framework can generalize to deep learning tasks remains questionable.
 V-C Federated Data Re-balancing
-The notorious problem of non-IID data across the local domains in the federation not only hampers the convergence of the global model, but also leads to the drop of accuracy when the final model is evaluated on a presumably balanced test set.
-From the angle of knowledge, these phenomena can be explained from two aspects.
-First, theoretical evidence [79, 80] has revealed the convergence speed of the global model, on a certain ML task, is very much determined by the divergence of parameter updates (or accumulated gradients) from the clients.
-This essentially comes down to discrepancy in local knowledge as reflected by the conditional distribution of features and labels, which results in diverging local gradients towards different local optima.
-Second, non-IID data across the system usually implies that clients could have conflicting knowledge.
-Once learned and aggregated into a global model, client-specific knowledge gets diluted or neutralized, which consequently undermines the model’s ability to generalize.
-Data re-balancing has proved its effectiveness in class-imbalanced learning where re-sampling is the most common technique adopted in the literature [81, 82].
+Data imbalance, usually associated with lopsided marginal probabilities $P(X)$ and $P(Y)$, creates difficulties for knowledge discovering.
+Data re-balancing has proved its effectiveness in class-imbalanced learning where re-sampling is the most common technique adopted in the literature [shi2023advances, zhou2025heterogeneous].
 By scaling up (down) the sampling probabilities of tail (dominant) classes, the local model is learned on a reshaped distribution of data.
 This can also be interpreted as a means of knowledge filtering where class-specific knowledge is kept by a different ratio.
-Considering that local data can be extremely imbalanced in non-IID settings, it is natural to explore collaborative knowledge filtering by means of data re-balancing.
 Tang et al.
-[76] investigated the impact of the celebrated class-balanced loss (CBLoss) [83] when applied to the local training of FL.
+[tang2021data] investigated the impact of the celebrated class-balanced loss (CBLoss) [cui2019cbloss] when applied to the federated setting.
 Two interesting observations were made.
-First, data re-sampling narrowed the gap between the probability distributions of label sampling between different clients, which consequently accelerates the global model’s convergence.
-Second, though reaching the plateau at a faster rate, re-sampling resulted in an inferior model that yielded lower accuracy on local data.
+First, data re-sampling narrowed the gap of the probability distributions of label sampling between different clients.
+Second, though reaching the plateau at a faster rate, re-sampling can result in an inferior model that yielded lower accuracy on local data.
 These two findings motivated them to devise a Imbalanced Weight Decay Sampling (IWDS) strategy that gradually smooths an imbalanced sampling towards uniform sampling.
@@
 with $\beta_{t}<1$ being the sampling parameter that decays from $\beta_{m}$ to $\beta_{0}$:
-$\beta_{t}=\beta_{m}+(\beta_{0}-\beta_{m})\cdot\gamma^{t}$ (7)
+$\beta_{t}=\beta_{m}+(\beta_{0}-\beta_{m})\cdot\gamma^{t},$ (5)
 where $\gamma$ is the decay rate parameter and $t$ is the index of the federated round.
@@
 $r^{(t)}_{k,i,j}=\frac{w^{(t)}_{k,i}}{w^{(t)}_{k,j}}=\frac{1-\beta_{t}^{n_{k,j}}}{1-\beta_{t}^{n_{k,i}}}.$ (6)
-This strategy leverages imbalanced re-sampling to speed up the convergence at the early stage of FL and gradually approaches uniform sampling for higher final accuracy [76], i.e., $r^{(t)}_{k,i,j}$ gets closer to 1 with $\beta_{t}$ decreasing.
+This strategy leverages imbalanced re-sampling to speed up the convergence at the early stage of collaboration and gradually approaches uniform sampling for higher final accuracy [tang2021data], i.e., $r^{(t)}_{k,i,j}$ gets closer to 1 with $\beta_{t}$ decreasing.
 Although distributional information is utilized locally to enhance local training, knowledge outside the local domain is left unexploited.
-Ouyang et al. [78] formulated a different weighting rule for each sample as a power function of class frequency. After normalization the sampling probability $P_{k,c}$ for a class-$c$ instance is:
+Ouyang et al.
+[ouyang2024ddpg] formulated a different weighting rule for each sample as a power function of class frequency.
+After normalization the sampling probability $P_{k,c}$ for a class-$c$ instance is:
 $P_{k,c}=\frac{n_{k,c}^{\beta_{k}}}{\sum_{i=1}^{C}n_{k,i}^{\beta_{k}}},$ (7)
-where $C$ is total number of classes and the client-specific parameter $\beta_{k}$ controls the sampling probability distribution on client $u_{k}$.
+where $C$ is total number of classes and the participant-specific parameter $\beta_{k}$ controls the sampling probability distribution on participant $u_{k}$.
 To dynamically determine $\{\beta_{k}|k=1,2,\ldots\}$ for the training cohort, a global coordinating agent based on Deep Reinforcement Learning (DRL) was developed.
 More specifically, the agent collects local and global model parameters as input state, makes decisions on $\{\beta_{k}|k=1,2,\ldots\}$, and receives rewards defined based on class-specific accuracy and Gini coefficients of local distribution.
-Most of the classic re-sampling techniques for imbalanced learning can be easily applied to the local training step in FL.
+Most of the classic re-sampling techniques for imbalanced learning can be easily applied to the local operation phase in TAF.
 This corresponds to a naïve implementation of federated data re-balancing where clients act independently based on local distribution.
-Chen [77] conducted empirical studies to investigate the impacts of local data re-balancing on the efficacy of FL.
-Over-sampling, under-sampling and the combination of these two were evaluated on a healthcare-related logistic regression task over ten clients.
-They also filtered out the outliers in the pre-processing stage by z-score.
-Two interesting finds were reported: i) Under-sampling caused a slight performance decrease in the final model’s accuracy and precision, and ii) the hybrid method (combining over-sampling and under-sampling) showed the best performance across the board in all metrics.
+Chen [chen2024implementing] conducted empirical studies to investigate its impacts on some downstream learning tasks.
+Over-sampling, under-sampling and the combination of these two were evaluated.
+Two interesting finds were reported: i) Under-sampling caused a slight performance loss in the final model’s accuracy and precision, and ii) the hybrid method (combining over-sampling and under-sampling) showed the best performance across the board in all metrics.
 These observations imply a strong potential of data re-balancing in the federation even without any knowledge exchange between the domains.
 However, whether the phenomenon still holds for more complex tasks remains under-explored.
+We summarize these solutions in Table III.
 V-D Summary
-In this section we demystify the representative approaches to knowledge filtering within a federation.
-With rising concerns of data-centric attacks, accurately identifying adversarial clients that train on poisoned local data can be a top priority with regard to system security and robustness.
-We argue that fine-grained filtering at sample level can be employed jointly with client selection to provide more reliable evidence for adversary detection.
-As of data re-balancing, existing approaches stand on an implicit assumption that uniform data distribution is ideal for local training.
+With rising concerns of data-centric attacks, dataset-level filtering can minimize the risk of involving poisonous data.
+We argue that fine-grained filtering at sample level can be employed jointly with participant screening to provide more reliable evidence for adversary detection.
+As of data re-balancing, existing approaches stand on an implicit assumption that uniform data distribution is ideal for the downstream task.
 However, this may not always hold.
-On the one hand, the local learner could under-fit on the “special knowledge” in local data when more attention is paid to the tail classes [76].
+On the one hand, the local learner could under-fit on the “special knowledge” [tang2021data].
 On the other hand, there is no sufficient theoretical evidence that supports imbalanced sampling, provided that the test data distribution is unknown.
-Although extensive effort has been made in addressing class-imbalanced learning, collaborative knowledge filtering for regression tasks, e.g.
+Besides, collaborative data refinement for regression tasks, e.g.
 drug response modeling, is hardly explored.
-Moreover, existing studies pay little attention to the subtle line between data imbalance and non-IID data, the understanding of which can help us better reshape the knowledge locally and globally.
-In Table III we summarize the strength and limits of existing studies in relevance to knowledge filtering in the federation.
-VI Label and Feature Space Correction
-In this section we review the advance of research on label and feature space correction, by which we refer to the branch of studies that employ collaborative mechanisms to make correction or transformation on data labels or features.
-Provided that the quality of local data can be dubious especially in a cross-device federated system, it is important to rectify the samples in each local domain.
-In contrast to knowledge filtering that aims to exclude misinformation, amendments of samples usually involves data engineering techniques and feature engineering methods.
-Hence, in the following content we survey the state-of-the-art under two different categories: federated data correction and federated feature engineering.
+Moreover, existing studies pay little attention to the subtle line between data imbalance and non-IIDness of data, the understanding of which can help us better redefine how data can refined locally and globally.
+VI Collective Harmonization of Data
+This section revisits branches of studies that employ collaborative mechanisms to make correction or transformation on data labels or features so that they can be utilized later in a harmonized fashion.
 VI-A Federated Data Correction
-As discussed in Section V, data noise commonly exists in a federation.
-Among various types and forms of noise, mislabeled samples are believed to have salient impact on supervised learning.
-Many researchers endeavored to address the problem of noisy labels in the context of FL, that is, federated noisy label learning (FNLL) [84, 85, 86, 87].
-Considering the tie between a client and its local domain, it is intuitive to first identify the “noisy clients” [88, 89].
-Xu et al.
-[88] proposed FedCorr, a multi-stage label denoising framework that streamlines a pre-processing stage and a fine-tuning stage prior to the traditional FL process, all executed in a federated fashion.
-In the federated pre-processing stage, two distinct mechanisms are employed for detecting noisy clients and correcting noisy labels.
-On the one hand, each client computes an estimate of the Local Intrinsic Dimensionality (LID) based on its local model’s predictions, which are collected by the central server for noisy client detection.
-As an useful measure that connects dimensionality and discriminability [90, 91], the LID for a vector $\boldsymbol{x}\sim\mathcal{D}$ is defined as:
+Mislabeled samples can have salient impact.
+This motivates researches on federated noisy label learning (FNLL) [jiang2024tackling, wu2023fednoro].
+Considering the tie between a participant and its local domain, Xu et al.
+[xu2022fedcorr] proposed FedCorr, a multi-stage label denoising framework that streamlines a pre-processing stage and a fine-tuning stage prior to the traditional learning process, all executed in a federated fashion.
+In the pre-processing stage, each participant computes an estimate of the Local Intrinsic Dimensionality (LID) based on its local model’s predictions, which are collected by the central server for noisy dataset detection.
+As an useful measure that connects dimensionality and discriminability [houle2013dimensionality, karger2002finding], the LID for a vector $\boldsymbol{x}\sim\mathcal{D}$ is defined as:
 ${LID}_{x}=\lim_{r\rightarrow 0}\lim_{\epsilon\rightarrow 0}\frac{\log F_{d_{x}}((1+\epsilon)r)-\log F_{d_{x}}(r)}{\log(1+\epsilon)},$ (8)
 where $d_{x}$ is the distance between $\boldsymbol{x}$ and a randomly drawn vector $\boldsymbol{x}^{\prime}$ from $\mathcal{D}$ and $F_{d_{x}}(\cdot)$ represents the cumulative distribution function (CDF) of $d_{x}$.
-As models trained on noisy data tend to report larger LID scores, the server then uses this rule to spot these clients for label correction.
-FedCorr [88] adopts a confidence-based rule to find and correct mislabeled data on noisy clients.
+As models trained on noisy data tend to report larger LID scores, the server then uses this rule to spot these participants for label correction.
+FedCorr [xu2022fedcorr] adopts a confidence-based rule to find and correct mislabeled data on noisy datasets.
 Specifically, the subset of samples (controlled by a proportion $\pi$) that yield the highest losses is first selected, following which the predictions of the global model $f_{G}$ are used for relabeling.
 Formally, the final subset of samples $D^{\prime}_{k}$ (for a noisy client $u_{k}$) to be relabeled is defined in (9):
-$D^{\prime}_{k}=\big{\{}(\boldsymbol{x},y)|(\boldsymbol{x},y)\in\underset{\hat{D}\subseteq D_{k}\atop|\hat{D}|=\pi|D_{k}|}{\arg\max}\sum_{(\boldsymbol{x},y)\in\hat{D}}\ell(f_{G}(\boldsymbol{x}),y)\\ \text{ and }\max f_{G}(\boldsymbol{x})\geq\theta\big{\}},$ (11)
-where $\ell$ is the loss function and $D_{k}$ is the local data from client $u_{k}$. The labels of $D^{\prime}_{k}$ are then corrected to the predictions by the global model $f_{G}$.
-There are two major defects in FedCorr [88]: i) incompatibility to general loss functions other than cross entropy, and ii) susceptibility to dishonest clients who tamper with their LID scores.
+$D^{\prime}_{k}=\big\{(\boldsymbol{x},y)|(\boldsymbol{x},y)\in\\ \underset{\hat{D}\subseteq D_{k}\atop|\hat{D}|=\pi|D_{k}|}{\arg\max}\sum_{(\boldsymbol{x},y)\in\hat{D}}\ell(f_{G}(\boldsymbol{x}),y)\text{ and }\max f_{G}(\boldsymbol{x})\geq\theta\big\},$ (9)
+where $\ell$ is the loss function and $D_{k}$ is the local data from participant $u_{k}$. The labels of $D^{\prime}_{k}$ are then corrected to the predictions by the global model $f_{G}$.
+There are two major defects in FedCorr [xu2022fedcorr]: i) incompatibility to general loss functions other than cross entropy, and ii) susceptibility to dishonest participants who tamper with their LID scores.
 Targeting at the second issue, Wang et al.
-[89] built upon FedCorr and redesigned its federated pre-processing stage to achieve robust label correction.
-The key idea behind is to craft an efficient Zero-Knowledge Proof (ZKP) [92] protocol by which the server (as the verifier) can verify the computation integrity of the LID scores reported by the clients (as the provers) in a batched manner.
+[wang2024icde] built upon FedCorr and redesigned its federated pre-processing stage to achieve robust label correction.
+The key idea behind is to craft an efficient Zero-Knowledge Proof (ZKP) [campanelli2022zkp] protocol by which the server (as the verifier) can verify the computation integrity of the LID scores reported by the participants (as the provers) in a batched manner.
 The proofs are generated both before and after label correction.
@@
 Similarly, the FedELC framework by Jiang et al.
-[84] detects clients that carry noisy data and corrects them in two separate stages.
-The first stage is almost the same as in [88, 85], whilst the second stage creates learnable soft labels for data correction optimized towards the minimization of a linear combination of three different entropy losses.
-FedELC was evaluated on a rich set of noise data settings including the CIFAR-N [93] dataset from Amazon Mechanical Turk which features a high noise rate of human-annotated labels.
+[jiang2024tackling] detects participants that carry noisy data and corrects them in two separate stages.
+The first stage is almost the same as in [xu2022fedcorr, wu2023fednoro], whilst the second stage creates learnable soft labels for data correction optimized towards the minimization of a linear combination of three different entropy losses.
+FedELC was evaluated on a rich set of noise data settings including the CIFAR-N [wei2021cifar-n] dataset from Amazon Mechanical Turk which features a high noise rate of human-annotated labels.
 Noisy label detection and correction can also be executed directly at sample level.
-Based on label uncertainty estimation [94], Zeng et al.
-[95] started from utilizing a well-trained global model’s predictions to estimate the probability that samples of a class $i\in\mathcal{C}$ are correctly labeled in a local dataset $D_{k}$:
+Based on label uncertainty estimation [northcutt2021confident], Zeng et al.
+[zeng2022clc] started from utilizing a well-trained global model’s predictions to estimate the probability $P_{k,i}$ that samples of a class $i\in\mathcal{C}$ are correctly labeled in a local dataset $D_{k}$, given in (10):
 $P_{k,i}=$ $\sum_{j\in\mathcal{C}}p(y=i|y^{*}=j)p(y^{*}=j|y=i)$
 $\approx$ $\sum_{j\in\mathcal{C}}p(y=i|y^{*}=j)\mathbb{E}_{\boldsymbol{x}\in D_{k,y=i}}p(y^{*}=j;\boldsymbol{x},\theta_{G})$
-$=$ $\mathbb{E}_{\boldsymbol{x}\in D_{k,y=i}}\sum_{j\in\mathcal{C}}p(y=i|y^{*}=j)p(y^{*}=j;\boldsymbol{x},\theta_{G})$
 $=$ $\mathbb{E}_{\boldsymbol{x}\in D_{k,y=i}}\sum_{j\in\mathcal{C}}p(y=i|y^{*}=j;\boldsymbol{x},\theta_{G})p(y^{*}=j;\boldsymbol{x},\theta_{G})$
@@
 $=$ $\frac{1}{|D_{k,y=i}|}\sum_{v=1}^{|D_{k,y=i}|}p(y_{v}=i;\boldsymbol{x}_{v},\theta_{G}),$ (10)
-where $D_{k,y=i}$ is the data subset on client $u_{k}$ labeled as class $i$, $\mathcal{C}\subset\mathbb{W}$ is the label space, and $\theta_{G}$ denotes the global model’s parameter.
-Next, the global estimate of the probability of correctly-labeled class $i$ can be obtained using weighted average consensus [95] over the population $\mathcal{K}$:
+where $D_{k,y=i}$ is the data subset on participant $u_{k}$ labeled as class $i$, $\mathcal{C}\subset\mathbb{W}$ is the label space, and $\theta_{G}$ denotes the global model’s parameter.
+Next, the global estimate of the probability of correctly-labeled class $i$ can be obtained using weighted average consensus [zeng2022clc] over the population $\mathcal{K}$:
 $P_{i}=\sum_{k\in\mathcal{K}}\frac{|D_{k,y=i}|}{\sum_{k\in\mathcal{K}}|D_{k,y=i}|}P_{k,i},$ (11)
-which resembles the model aggregation rule in the canonical FedAvg [96].
+which resembles the model aggregation rule in the canonical FedAvg [mcmahanfl].
 A practical concern of federated data correction is that the participants can be self-interested, which means that they may not stick to the protocol if their gain of participation cannot cover the cost of data correction.
 In light of this issue, Yan et al.
-[97] first modeled the payoff of clients as a function of the model’s accuracy, payment from the task publisher, and the data correction cost.
-It is worth noting that the authors employed an iterative denoising strategy that corrects a fraction of $\alpha$ noise samples per round.
+[yan2024incent] first modeled the payoff of participants as a function of the model’s accuracy, payment from the task publisher, and the data correction cost.
+The authors employed an iterative denoising strategy that corrects a fraction of $\alpha$ noise samples per round.
 To reduce the fraction of noisy data from $\epsilon$ to $\epsilon^{\prime}$ given a denoising cost per sample $c$ and data size $|D_{k}|$, the total cost of data correction for client $u_{k}$ is:
@@
 Yan et al.
-[97] considered two different situations of federated denoising.
-In the first case, clients are cooperative and work towards to the optimal social welfare, i.e., the maximization of the sum of client payoffs.
-In the second case, every client is assumed self-interested and only motivated to maximize its own payoff, which corresponds to a non-cooperative game.
+[yan2024incent] considered two different situations of federated denoising.
+In the first case, participants are cooperative and work towards to the optimal social welfare, i.e., the maximization of the sum of client payoffs.
+In the second case, every participant is assumed self-interested and only motivated to maximize its own payoff, which corresponds to a non-cooperative game.
 VI-B Federated Feature Engineering
-Prior to the boom of deep learning, feature engineering had proved very useful in alleviating the “curse of dimensionality” and allowing the model to learn more generalizable knowledge from the data.
-In the context of KAF, by federated feature engineering we refer to a new thread of research with particular focus on collaborative knowledge augmentation by means of aligning, selecting and transforming the feature space (Fig.
+In TAF, by federated feature engineering we refer to a new thread of research with particular focus on aligning, selecting and transforming the feature space (Fig.
 6).
-This process can either be independent of the downstream task or somehow coupled with a specific federated learning protocol.
-In both cases, feature engineering across multiple domains in the federation poses new challenges in several aspects.
-  • Disordered or misaligned attributes: in cross-institutional cooperation it is common to deal with different views of data where the local datasets are subject to different permutations of attributes [71].
+This process can either be independent of the downstream task or somehow coupled with a specific protocol, motivated to address three challenges:
+  • Disordered or misaligned attributes: in cross-institutional cooperation it is common to deal with different views of data where the local datasets are subject to different permutations of attributes [thakur2024knowledge].
   • Heterogeneous feature space: from a single domain to a multitude of domains with partial overlap in the feature space, the heterogeneity brings great difficulty to the design of autonomic feature engineering.
   • Varied significance of features: non-IID settings usually mean bias and special knowledge within each domain, which will greatly influence the importance of features when measured statistically.
-These challenges have catalyzed rising interest in federated heterogeneous learning and also motivated cutting-edge studies on new feature engineering paradigms.
+These challenges have catalyzed cutting-edge studies on new feature engineering paradigms.
 In this section we discuss the advance in federated feature engineering, with particular focuses on collaborative feature selection and feature transformation.
-Figure 6: A schematic illustration of collaborative feature engineering under the KAF paradigm.
-Fed-FiS [98] represents an early exploration of collaborative feature selection in the federation based on information theory.
+Figure 6: A schematic illustration of federated feature engineering.
+Fed-FiS [banerjee2021fedfis] represents an early exploration of collaborative feature selection in the federation based on information theory.
 The study follows a training-free, task-agnostic approach to assessing the importance of features.
-Locally, a client $u_{k}$ independently evaluates the feature-feature mutual information (FFMI) and the feature-class mutual information (FCMI) over the local feature and label set $\{f^{(k)}_{1},f^{(k)}_{2},\ldots\}\cup C^{(k)}$ by Eqs.
+Locally, a participant $u_{k}$ independently evaluates the feature-feature mutual information (FFMI) and the feature-class mutual information (FCMI) over the local feature and label set $\{f^{(k)}_{1},f^{(k)}_{2},\ldots\}\cup C^{(k)}$ by Eqs.
 (13) and (14):
@@
 ${FFCI}^{k}_{i}=\sum_{x\in f^{(k)}_{i}}\sum_{c\in C^{(k)}}p(x,c)\log\frac{p(x,c)}{p(x)p(c)}.$ (14)
-Based on these mutual information-based measures, each client performs clustering to select a subset of features which are then uploaded to the central server.
-In the second stage, the server uses FFMI and FCMI again to score each feature, which in turn helps the clients identify the best local feature subset.
+Based on these mutual information-based measures, each participant performs clustering to select a subset of features which are then uploaded to the central server.
+In the second stage, the server uses FFMI and FCMI again to score each feature, which in turn helps the participants identify the best local feature subset.
 Banerjee et al.
-[99] further extended this framework using an improved ranking scheme based on Pareto optimization where each feature is treated as a solution characterized by FFMI and FCMI.
-The features are then ranked based on the relation of domination and reported back to the clients.
+[banerjee2024cost] further extended this framework using an improved ranking scheme based on Pareto optimization where each feature is treated as a solution characterized by FFMI and FCMI.
+These solutions are then ranked based on how many others dominate them and that they dominate.
+Rankings are reported back to the clients.
 Fu et al.
-[100] studied the feature selection problem in the Vertical Federated Learning (VFL) setting where the set of entities are associated with consistent identifies but represented by different features.
-The proposed framework, FEAST, involves three types of clients for collaborative feature selection: a selected party, an active party, and candidate parties.
+[fu2023feast] studied the feature selection problem in the Vertical Federated Learning (VFL) setting where the set of entities are associated with consistent identifiers but represented by different features.
+The proposed framework, FEAST, involves three types of participants for collaborative feature selection: a selected party, an active party, and candidate parties.
 In the pre-processing stage, discretization is first applied to each feature, combined with stratified sampling to reduce computation and communication costs.
@@
 Specifically, a party is selected to report its statistical variables which are generated by grouping local features and re-encoding the instances.
-The candidate parties receive the statistical variables and score each feature based on Conditional Mutual Information (CMI).
-The feature scores are then collected by an active party to perform global ranking to produce the best feature subset $S_{k}$ (s.t., $|S_{k}|=m$ given $m$ as the number of local features to select) for each client $u_{k}$, which collectively represents the solution to the following optimization:
-$\min_{S_{k}\subseteq\mathcal{X}_{k}}H(Y|\{S_{1},S_{2},\ldots\}).$ (17)
-where $H(Y|\{S_{1},S_{2},\ldots\})$ is conditional entropy of the label $Y$ given the union of locally selected feature subsets $\{S_{1},S_{2},\ldots\}$.
-This framework offers useful insights for communication-efficient exchange of knowledge for collaborative feature selection, but the protocol does not consider dishonest parties (clients) in the system.
+The candidate parties receive the statistical variables and score each feature based on Conditional Mutual Information (CMI) [yang1999feature].
+The feature scores are then collected by an active party to perform global ranking to produce the best feature subset $S_{k}$ (s.t., $|S_{k}|=m$ given $m$ as the number of local features to select) for each participant $u_{k}$, which collectively represents the solution to the following optimization:
+$\min_{S_{k}\subseteq\mathcal{X}_{k}}H(Y|\{S_{1},S_{2},\ldots\}),$ (15)
+where $H(Y|\{S_{1},S_{2},\ldots\})$ is the conditional entropy of the label $Y$ given the union of locally selected feature subsets $\{S_{1},S_{2},\ldots\}$.
+This framework offers useful insights for communication-efficient exchange of knowledge for collaborative feature selection, but the protocol does not consider dishonest participants in the system.
 As a matter of fact, feature engineering techniques can be utilized to defend against data poisoning attacks strengthened by feature manipulation.
-This is also termed feature poisoning (FP) in the literature [101].
-Experimental evidence has demonstrated that, despite being highly disguised, FP attacks can be effectively countered by the Random Deep Feature Selection (RDFS) [102] strategy.
+This is also termed feature poisoning (FP) in the literature [Nowroozi2024rdfs].
+Experimental evidence has demonstrated that, despite being highly disguised, FP attacks can be effectively countered by the Random Deep Feature Selection (RDFS) [Nowroozi2025fp] strategy.
 Beside feature selection, feature transformation represents another typical approach to data engineering.
@@
 However, it is not trivial for data holders in a federated system to perform feature transformation collaboratively.
-A practical problem that stands out is the computation cost incurred by performance evaluation provided numerous combinations of transformed features.
+A practical problem that stands out is the computation cost incurred by performance evaluation given numerous combinations of transformed features.
 Moreover, privacy issues arise since transformation can involve multiple features coming from different domains, which necessitates Secure Multi-Party Computation (MPC).
 To address these issues, Fang et al.
-[103] designed FLFE, a federated feature engineering framework for efficient, privacy-preserving multi-party collaboration.
+[fang2020flfe] designed FLFE, a federated feature engineering framework for efficient, privacy-preserving multi-party collaboration.
 The core techniques employed include a sketching scheme for quantizing local features, a DNN-based classifier for judging the value of features, and a mask-based feature exchange protocol for transformation.
 This brings advantages in communication efficiency and also preserves data privacy without homomorphic encryption that limits the diversity of transformations.
-TABLE IV: Summarizing the studies related to federated label and feature space correction in terms of methodological highlights, drawbacks, datasets for experiments, and the form of knowledge for exchange.
+TABLE IV: Summarizing the studies related to collective harmonization of data in terms of methodological highlights, drawbacks, datasets for experiments, and the form of knowledge for exchange.
 Refs. Category Knowledge carrier Highlights Drawbacks Datasets
-[88, 89] Fed. label correction local intrinsic dimensionality (LID) and models $\bullet$ global pre-processing plus “clean” clients-only fine-tuning
+[xu2022fedcorr, wang2024icde] Federated label correction local intrinsic dimensionality (LID) and models $\bullet$ global pre-processing plus “clean” clients-only fine-tuning
 $\bullet$ local knowledge encapsulated in LID
 $\bullet$ incompatibility to general loss functions CIFAR-10, CIFAR-100, Clothing1M, and simulated data
-[84] Fed. label correction warm-up models and class-wise loss vectors $\bullet$ distance-aware aggregation to mitigate noisy knowledge
+[jiang2024tackling] Federated label correction warm-up models and class-wise loss vectors $\bullet$ distance-aware aggregation to mitigate noisy knowledge
 $\bullet$ learnable soft labels for local correction
 $\bullet$ sensitive to the setting of multiple hyper-parameters CIFAR-10/100, CIFAR-10/100-N, Clothing1M
-[95] Fed.
-label correction class-wise confidence thresholds and models $\bullet$ consensus-based method to ensemble decentralized knowledge $\bullet$ reliance on a pre-trained model for label uncertainty estimation MNIST, USC-HAD, CIFAR-10, and Clothing1M
-[97] Fed. label correction denoising rate decisions and models $\bullet$ payoff modeling for game-based interaction
+[zeng2022clc] Federated label correction class-wise confidence thresholds and models $\bullet$ consensus-based method to ensemble decentralized knowledge $\bullet$ reliance on a pre-trained model for label uncertainty estimation MNIST, USC-HAD, CIFAR-10, and Clothing1M
+[yan2024incent] Federated label correction denoising rate decisions and models $\bullet$ payoff modeling for game-based interaction
 $\bullet$ iterative denoising with a tunable fraction
 $\bullet$ reliance on accurate estimate of noise rate of private data MNIST and CIFAR-10
-[98, 99] Fed. feature selection MI-based scores and feature subsets $\bullet$ low communication costs
+[banerjee2021fedfis, banerjee2024cost] Federated feature selection MI-based scores and feature subsets $\bullet$ low communication costs
 $\bullet$ model-agnostic and task-independent
@@
 SL-KDD99 and anonymized credit card transactions (ACC)
-[100] Fed. feature selection statistical variables and feature scores $\bullet$ MI-based and model-independent
+[fu2023feast] Federated feature selection statistical variables and feature scores $\bullet$ MI-based and model-independent
 $\bullet$ efficient exchange of statistical knowledge
@@
 MIMIC III, PhysioNet, Census-Income, and Nomao
-[104] Fed. feature selection feature subsets $\bullet$ low communication cost
+[zhang2023fshfl] Federated feature selection feature subsets $\bullet$ low communication cost
 $\bullet$ unsupervised and model-agnostic feature selection
@@
 Bot-IoT, ACC, KDD99, and DEFT
-[105] Fed. feature transform. expression of derived features $\bullet$ low communication cost and reduced search space $\bullet$ very high cost of feature subset evaluation OpenML586 and Airfoil
-[106] Fed. feature selection probability vectors associated with features $\bullet$ probabilistic feature subset selection and Bayesian aggregation
+[cassara2022tvt] Federated feature selection probability vectors associated with features $\bullet$ probabilistic feature subset selection and Bayesian aggregation
 $\bullet$ significantly reduced network traffic
 $\bullet$ convergence guarantee for IID data only MAV and WEarable Stress and Affect Detection (WESAD)
-[103] Fed. feature transform. sketches of features $\bullet$ Quantile Data Sketch and mask-based feature exchange
+[overman2024federated] Federated feature transform expression of derived features $\bullet$ low communication cost and reduced search space $\bullet$ very high cost of feature subset evaluation OpenML586 and Airfoil
+[fang2020flfe] Federated feature transform sketches of features $\bullet$ Quantile Data Sketch and mask-based feature exchange
 $\bullet$ model-free evaluation
 $\bullet$ pre-training required for the DNN-based feature judge 166 datasets from OpenML
-A major obstacle for adapting Automated Feature Engineering (AutoFE) methods (e.g., DIFER [107], OpenFE [108] and Cognito [109]) to the federated system is the prohibitive cost of communication.
-To this end, Overman and Klabjan [105] have made early effort and devised a federated framework for collaborative AutoFE.
-In the horizontal FL setting, clients independently carry out AutoFE and report the expression of derived features to the server.
+A major obstacle for adapting Automated Feature Engineering (AutoFE) methods (e.g., OpenFE [zhang2023openfe]) to the federated system is the prohibitive cost of communication.
+To this end, Overman and Klabjan [overman2024federated] have made early effort and devised a federated framework for collaborative AutoFE.
+In their framework, participants independently carry out AutoFE and report the expression of derived features to the server.
 The server then randomly applies masks to the union of these features to generate feature subsets.
-An iterative halving strategy is used to find the best mask (e.g., optimal feature subset) after running multiple rounds of standard FL for performance evaluation.
-For vertical FL setting, feature values across the entire local dataset need to be shared after homomorphic encryption so as to fuse features from different domains.
-For hybrid settings, importance of common features across multiple domains can be assessed progressively with the training process.
-Federated feature engineering can facilitate the refinery of knowledge in a broad range of scenarios.
+An iterative halving strategy is used to find the best mask (e.g., optimal feature subset) after running multiple rounds of learning for performance evaluation.
+For participants with heterogeneous feature space, feature values across the entire local datasets need to be shared after homomorphic encryption so as to fuse features from different domains.
+For hybrid settings, importance of common features across multiple domains can be assessed progressively with the process.
+Federated feature engineering can facilitate the refinement of knowledge in a broad range of scenarios.
 For example, Zhang et al.
-[104] proposed a federated feature selection framework for Internet of Things (IoT) applications where each device (as a client) has very limited computational resources.
-In this framework, local features $\{f^{(k)}_{1},f^{(k)}_{2},\ldots\}$ on each client $u_{k}$ are first filtered using a threshold defined based on Symmetric Uncertainty (SU) and Feature Average Relevance (FAR):
+[zhang2023fshfl] proposed a federated feature selection framework for Internet of Things (IoT) applications where each device (as a participant) has very limited computational resources.
+In this framework, local features $\{f^{(k)}_{1},f^{(k)}_{2}$$\ldots\}$ on each client $u_{k}$ are first filtered using a threshold defined based on Symmetric Uncertainty (SU) and Feature Average Relevance (FAR):
 ${SU}(f^{(k)}_{i},f^{(k)}_{j})=2\Big[\frac{{MI}(f^{(k)}_{i},f^{(k)}_{j})}{H(f^{(k)}_{i})+H(f^{(k)}_{j})}\Big$], (16)
 ${FAR}(f^{(k)}_{i})=\frac{1}{d_{k}}\sum_{j=1}^{d_{k}}{SU}(f^{(k)}_{i},f^{(k)}_{j})$, (17)
-where $d_{k}$ is the feature dimensionality for client $u_{k}$, $H(\cdot)$ is the entropy measure and $MI(\cdot,\cdot)$ denotes the mutual information between two features.
-Next, One-Class Support Vector Machine (OCSVM) and hierarchical clustering are employed on each client to identify the subset of most significant features.
-The server then calculates the intersection of these subsets and synchronizes with the clients.
+where $d_{k}$ is the feature dimensionality for participant $u_{k}$, $H(\cdot)$ is the entropy measure and $MI(\cdot,\cdot)$ denotes the mutual information between two features.
+Next, One-Class Support Vector Machine (OCSVM) and hierarchical clustering are employed on each participant to identify the subset of most significant features.
+The server then calculates the intersection of these subsets and synchronizes with the participants.
 The solution was evaluated on four classification tasks in relevance to IoT applications.
-Experimental results demonstrated that competitive and even higher performance could be achieved with the federated feature subset, with improved training efficiency.
+Experimental results demonstrated that competitive and even higher performance could be achieved with the federated feature subset, with improved efficiency.
 Cassará et al.
-[106] argued that a small subset of features are adequate for collaborative learning in interconnected Autonomous Driving Systems (ADSs).
+[cassara2022tvt] argued that a small subset of features are adequate for an interconnected Autonomous Driving Systems (ADSs).
 To this end, they developed a federated feature selection algorithm that combines on-vehicle feature filtering and edge-based feature subset aggregation.
-Locally, each participating vehicle (as a client) optimizes a feature selection probability vector such that the conditional entropy $H(y|S)$ (given feature subset $S$ and labels $y$) is minimized, which corresponds to an Associated Stochastic Problem (ASP).
+Locally, each participating vehicle optimizes a feature selection probability vector such that the conditional entropy $H(y|S)$ (given feature subset $S$ and labels $y$) is minimized, which corresponds to an Associated Stochastic Problem (ASP).
 The edge server then aggregates these probability vectors into a consensus one and feeds it back to the vehicles.
@@
 The Kolmogorov-Smirnov test with preset parameters are used for checking convergence.
-By evaluating this algorithm, the authors claimed that a 99% compression rate for network traffic was achieved since as few as 24 (out of 2166) features were selected for collaborative training on the MAV dataset, with negligible impact on model performance.
-This finding indicates that federated feature engineering can be as effective as model-based compression methods (e.g., quantization, pruning, etc.) [110] in terms of saving learning costs.
+By evaluating this algorithm, the authors claimed that a 99% compression rate for network traffic was achieved since as few as 24 (out of 2166) features were selected for collaborative training on the MAV dataset, with negligible impact on downstream tasks.
+This finding indicates that federated feature engineering can be as effective as model-based compression methods (e.g., quantization, pruning, etc.) [gupta2022compression] in terms of cost saving.
 VI-C Summary
-In this section we highlight emerging studies that work on the dimensions in which data are represented through collaborative effort from multiple parties.
-We argue that they symbolize an important approach to knowledge augmentation in federation.
-On the one hand, federated data correction attempts to amend the misinformation carried by the data to avoid erroneous knowledge, with particular focus on the labels.
-Reference knowledge, in most of the label-correction methods, is encapsulated in the global model.
-On the other hand, federated feature engineering focuses on identifying the most important attributes or the most expressive combination of them.
+In this section we highlight emerging studies that tailor the dimensions in which data are represented.
+Federated data correction attempts to amend the misinformation carried by the data to avoid erroneous knowledge, with particular focus on the labels.
+Reference knowledge, in most of the label-correction methods, is encapsulated in a global model.
+This limits their application to task-agnostic scenarios.
+Federated feature engineering focuses on identifying the most important attributes or the most expressive combination of them.
 In this context, the statistical properties of features are often utilized as the measure of knowledge they can provide.
-Although different in methodologies (summarized in Table IV, the two lines of research have some common points in terms of implementation.
+Their pros and cons are summarized in Table IV.
+Although being different in methodologies, the two lines of research have much in common.
 First, both data correction and feature engineering are typically implemented as part of a multi-stage procedure in the federation.
-For example, data correction is usually cooperated by the clients in the pre-processing stage before the actual training process such as FL.
+For example, data correction is usually cooperated by the participants in the pre-processing stage before the downstream tasks.
 Similarly, federated feature engineering usually involves repetitive local feature assessment and client-server information exchange.
 In addition, we discover that both approaches can be sensitive to the discrepancy of data distribution across different domains.
-A unified posterior (that describes the conditional probability of labels) or an optimal feature subspace may not work well for every individual client in the federated system.
+A unified posterior (that describes the conditional probability of labels) or an optimal feature subspace may not work well for every individual participant in the federated system.
 VII Open Challenges and Questions
-Through revisiting diverse approaches towards more effective, efficient learning from various perspectives of knowledge, we argue that there is a large space ahead to explore in the domain.
-The advance in knowledge discovery combined with the promising potential of federated learning is expected to open up new opportunities for knowledge-centric, learning-driven collaboration in the future.
-In the paper we highlight urgent challenges and open questions that involve but are not limited to the ways knowledge is learned, represented and exchanged as well as how to enhance Large Language Models (LLMs) with KAF.
-We expect to offer unique insights for research beyond knowledge for learning.
-VII-A Decentralized Peer-to-Peer Knowledge Augmentation
-Extensive effort has been made in decentralized machine learning since it has natural advantages in scalability, flexibility and byzantine-robustness [111, 112].
-However, existing approaches are mostly learning-centric, which means that their objectives are shaped towards the optimization of a specific group of models or a family of relevant tasks.
-This makes the collaboration costly and hard to reproduce, because the model as the knowledge carrier is very sensitive to the uncertainty in federated training.
-This uncertainty boils down to the population involved as well as the sequence of knowledge propagation in the network, which also raises the cost of reproducing the model.
-Recent studies on decentralized matrix factorization (DMF) present a domain-specific path to expand local knowledge domain via peer-to-peer (P2P) communications [113, 114, 115].
-DMF can be regarded as a special case of decentralized knowledge augmentation.
+Through revisiting diverse approaches which bring generic knowledge into local data domains in a federated system, we argue that there is a large space ahead to explore especially in the aspects of efficiency and efficacy.
+In this section we highlight opportunities for future research in relevance to TAF.
+VII-A Decentralized Peer-to-Peer Knowledge Seeking
+Decentralization has natural advantages in scalability, flexibility and byzantine-robustness [vincent2023systematic].
+Recent studies on decentralized matrix factorization (DMF) present a domain-specific path to expand local data via peer-to-peer (P2P) communications [hou2023bfrecsys, huang2024towards].
+DMF can be regarded as a special case of decentralized data augmentation, yet it is hard to deal with general forms of data other than matrices.
 With a novel vision, Abdelmoniem et al.
-[116] described a decentralized knowledge routing framework where a client can describe the knowledge requested and find another well-matched client for collaborative learning.
+[abdelmoniem2025] described a decentralized knowledge routing framework where a participant can describe the knowledge requested and find another well-matched peer for collaborative learning.
 This is done by sharing artificial samples with the proxy devices (called knowledge discoverers) responsible for request routing and handling.
-Even though the objective employed is still task-specific and learning-oriented (deficiencies discussed in Sec.
-I still apply), it inspires us to explore new directions for decentralized knowledge exchange and knowledge matching between domains.
-Clues for further study: knowledge abstraction and matching, blockchain, zero-knowledge proof, and knowledge routing.
+Though the objective employed is still task-specific and learning-oriented, it inspires us to explore new directions for decentralized knowledge exchange and knowledge matching between domains.
+Clues for further study: knowledge abstraction and matching [wu2023fedprof, ahmed2024fi], blockchain, zero-knowledge proof, and knowledge routing [abdelmoniem2025].
 VII-B Light-weight Knowledge Courier and Efficient Exchange
-A key factor in knowledge augmentation in federation is the exchange of knowledge between distributed local domains.
-For example, the item factor matrix is shared in the federation as the carrier of domain-specific knowledge in federated collaborative filtering, whilst the global model takes this role in federated label denoising.
+A key factor in task-agnostic federation is the exchange of knowledge between distributed local domains.
+For example, the item factor matrix is shared as the carrier of domain-specific knowledge in federated collaborative filtering, whilst the global model takes this role in federated label denoising.
 Whatever form it takes for knowledge exchange, the reduction of communication cost persists as a top priority in practice.
 On the application layer, the roadmap for efficient knowledge exchange mainly encompasses two approaches: i) decreasing the communication payload incurred by the courier of knowledge, and ii) reducing the number of communication rounds if iterative interactions are needed.
-Although FL is more communication-efficient than traditional distributed learning, it is worth noting that models, model updates and gradients are “heavy”.
-Provided that it is not necessary to exchange the “full picture” of knowledge in many situations such as federated knowledge filtering and label space correction, how to craft a flexible protocol to avoid information redundancy emerges as new problem to be resolved.
-Clues for further study: model-free knowledge representation, information bottleneck, and Bayesian compression.
-VII-C Federated LLM Augmentation with Exogenous Knowledge
-It is impossible for any model, even as complex as the latest LLMs [5, 6, 117], to learn and store comprehensive knowledge that covers all domains of training data.
-Moreover, language models usually suffer in reproducing accurate knowledge, a.k.a.
-the problem of hallucinations.
-Recent studies have paid attention to the federated Retrieval Augmentation Generation (RAG) [118], which is a promising technique for introducing exogenous knowledge to assist in model reasoning.
-These attempts target at addressing the “information scattering” effect [119] by searching multiple domains, yet the trade-off between performance gain and retrieval cost is mostly overlooked.
-The effectiveness and efficiency of federated RAG hinge on two factors, that is, whether a selected exogenous knowledge base well matches the given query, and whether the content retrieved can provide qualified knowledge.
+Existing studies often exchange models, but it is worth noting that models, model updates and gradients are “heavy”.
+Provided that it is not necessary to exchange the “full picture” of knowledge in many situations such as federated data filtering and harmonization, how to craft a flexible protocol to avoid information redundancy emerges as new problem to be resolved.
+Clues for further study: model-free knowledge representation, information bottleneck [yang2024ib], and Bayesian compression [yang2025continual].
+VII-C Federated Retrieval of Exogenous Knowledge with Confidentiality
+Latest LLM-empowered dialogue systems [guo2025deepseek, openaio1systemcard] commonly adopt Retrieval-Augmented Generation (RAG) to bring in up-to-date, verified information so as to produce accurate knowledge, a.k.a.
+reducing hallucinations.
+With rising awareness of intellectual properties, recent studies have paid attention to the federated Retrieval-Augmented Generation (FedRAG) [shojaee2025federated] with requirements of context confidentiality [addison2024cfedrag].
+This makes a special case of TAF where knowledge exchange (including queries and context) happens between multiple knowledge bases (KBs) and a reasoning agent.
+In this scenario, few studies stand on the KB side and they hardly account for the cost of communication and the price of retrieval service.
 On the one hand, it is very expensive in time to reliably determine whether a huge database contains the piece of knowledge desired.
 On the other hand, the scarcity of user feedback brings difficulties for assessing the usefulness of contents retrieved.
-To harness the potential of federated retrieval for enhanced LLM services, deeper investigation into the evaluation of exogenous knowledge is needed.
-Clues for further study: federated RAG, federated search systems, query-content evaluation, and knowledge base summary.
-VII-D KAF for Unlabeled data
-We observed that the majority of researches in relevance to knowledge augmentation and federated learning were designed for or evaluated on supervised learning tasks.
-This is mainly attributed to the prior knowledge implied by the labels that usually serve as the most important criterion for many operations.
+In addition, it is worthwhile to explore how multiple knowledge bases can cooperate with each other, e.g., in a TAF setting, to provide refined and logically consolidated retrieval capability.
+Clues for further study: federated RAG with confidentiality [addison2024cfedrag], federated search systems [guerraoui2025efficient], query-content evaluation [yu2024knowledge], and scattered knowledge bases [wu2025multirag].
+VII-D Task-agnostic Collaboration over Incomplete Data
+We observed that the majority of researches in the federated setting assume the completeness of data, overlooking the fact that real-world data have missing attributes and labels.
+Especially, many existing studies rely on data labels for knowledge extraction and empirically associate non-IIDness to the marginal distribution of labels.
 For example, feature engineering typically relies on feature-label mutual information, and data re-balancing strategies re-weight data points by class label.
-It is also worth mentioning that most existing empirical studies on federated learning associate non-IIDness to the marginal distribution of labels.
-However, this largely limits their application to a broader categories of unlabeled data such as text corpora, surveillance footage and browsing records.
-We argue that the main challenge for KAF over unlabeled data lies in the perplexity of knowledge within each domain.
-This means that, without the a priori from annotations, the increased uncertainty of data (e.g., measured in entropy) may result in lower confidence in the knowledge learned.
-In this case, more attention should be paid to the intrinsic structures of data and the impact of non-IID data beyond classification.
-Clues for further study: active learning, contrastive learning, soft labels, and knowledge distillation.
-VII-E Can KAF Make Independent Training Superior over FL?
-We argue that refined representation of knowledge is what collaborative learning can bring back to data.
-This is the key point that distinguishes KAF from what have been explored in the relevant field such as Federated Learning.
-As discussed, knowledge augmentation in federation can enhance the local knowledge in each domain by means of expansion, filtering and correction.
-This creates possibility for each participant to learn independently after the augmentation.
-But how competitive will it be as compared to Federated Learning?
-The answer is uncertain and hinges on multiple factors including local data size, the test data distribution, and the utility of knowledge gained via KAF.
-In terms of performance, KAF can increase the size and diversity of samples necessary for stochastic gradient-based local training, with augmented knowledge from foreign domains for generalizable model.
-In terms of cost, FL struggles in communication efficiency with large models and huge client population, whilst KAF plus independent training has tempting potential when equipped with efficient knowledge exchange protocols.
-From both perspectives, the superiority of KAF is still an open-ended question.
-Clues for further study: benchmarking, sample-efficient learning, and marginal gains of collaboration.
-VII-F Can Knowledge be Borrowed and Unlearned in a Federation?
-“The right to be forgotten”, as stated in privacy-protecting regulations, has given rise to the interest in the research of Machine Unlearning [120, 121].
-With potential value in achieving security, privacy, usability, and fidelity in various application scenarios [121], the concept of machine unlearning has also been adopted in FL systems to enable a specific client’s knowledge to be removed from what has been learned by the model [122].
-However, existing approaches are basically limited to the unlearning of class-oriented knowledge [123].
-More importantly, the process of unlearning can cost as much as re-training in terms of time and resources.
-This calls for further study on approximate federated unlearning to enable efficient and generalizable knowledge erasure.
-In the context of KAF, collaboratively augmented data can be reckoned as “borrowed” knowledge.
-New mechanisms are required in order to accurately identify the synthetic data points that characterize the designated part of knowledge to be forgotten.
-Clues for further study: federated machine unlearning, model inversion, gated knowledge transfer, noise learning, and influence-aware filtration.
+This means that they cannot apply to collaboration over incomplete data.
+This largely limits their application to a broader categories of unlabeled data such as text corpora, surveillance footage and browsing records.
+We argue that it is important while challenging for the data owners in TAF to represent incomplete data in a traditional form.
+For example, it is impossible to swap class-wise information over unlabeled datasets.
+In this case, we suggest further exploration into flexible representations of data based on their intrinsic structures and robust protocols for knowledge exchange between parties of varied data conditions.
+Clues for further study: fuzzy representations, auxiliary supervision [yang2024depth], sample fusion, and knowledge-data alignment [zhao2024weakly] .
+VII-E Can TAF Make Autonomy Superior over FL?
+We argue that the gain of collaborative learning becomes marginal when the participants’ data get larger and diversified.
+This is the key factor that limits the sustainability of learning-centric collaboration such as FL.
+On the contrary, task-agnostic federation liberates the participants from having to work on a specific model.
+This creates possibility for each participant to learn independently after the collaboration.
+But how competitive will it be regarding the downstream task when compared to federated learning?
+The answer is uncertain and hinges on multiple factors including local data size, the test data distribution, and the utility of data gained via TAF.
+In terms of performance, TAF can increase the size and diversity of samples necessary for independent training, with augmented knowledge from foreign domains for generalizable model.
+In terms of cost, FL struggles in communication efficiency with large models and huge participant population, whilst TAF plus independent training has tempting potential when equipped with efficient knowledge exchange protocols.
+From both perspectives, it is still an open-ended question.
+Clues for further study: benchmarking, query-based knowledge transfer [alballa2025qkt], and autonomous organization [zhang2025srfl].
+VII-F Can TAF Support Knowledge Erasure?
+“The right to be forgotten”, as stated in privacy-protecting regulations, has motivated recent studies in unlearning knowledge gained in a group [chen2024fmu].
+However, many existing unlearning approaches sit on unrealistic assumptions such as storing the entire trace of gradients or forcing all participants to engage in the unlearning stage [tarun2024fast].
+More importantly, they are yet to be evaluated on realistic unlearning scenarios where the shared knowledge needs to be “withdrawn” any time on demand.
+In contrast to these traditional settings, TAF opens up new opportunities for knowledge erasure.
+On the one hand, collaboratively augmented data can be reckoned as “borrowed” knowledge which is easy to remove.
+On the other hand, we believe it is also possible to accelerate any downstream unlearning task by exchanging artificial samples dedicated for “overwriting” specific knowledge.
+In both ways TAF can support efficient knowledge erasure.
+It also calls for innovative data engineering and knowledge exchange protocols in response to this special demand.
+Clues for further study: Augmentation-based unlearning [falcao2025data, Choi2024CVPR], Adversarial Mixup [peng2025adversarialmixup], inhibited synthetic data [zhang2025toward], influence-aware filtration [dang2025efficient], and noise learning [chien2024langevin].
 VIII Conclusion
-As the value of data assets and privacy issues arise, it is of utmost importance for organizations to make full use of proprietary data whilst seeking cooperation with their counterparts.
-Although the collaborative paradigm of federated learning has paved the way for the sharing of knowledge across domains with privacy preservation, research in the field hardly explored the possibilities beyond model-centric learning.
-We argue that knowledge sharing is not just limited to the purposes of model optimization, but can also serve as the key to local domain augmentation, which substantially diversifies the way how participants can benefit from the federation.
-With this vision, we in this paper present knowledge augmentation in federation, a conceptual paradigm for the expansion, filtration and correction of knowledge across decentralized domains.
-We revisit emerging research in the field via the lens of knowledge and provide a category of studies according to their methodologies.
-Through comparison and discussion we point out that the carrier of knowledge and the protocol for knowledge exchange are the most influential factors for KAF to be effective and efficient.
-Based on our investigation, we further highlight a number of open challenges that are not adequately addressed by existing effort, and also pose two open-ended questions to inspire more theoretical and empirical exploration in the relevant threads of research.
-Acknowledgment
-This work is supported by the National Natural Science Foundation of China, National Social Science Foundation of China Post-funded project, Basic and Applied Basic Research Project of Guangzhou, and the Fundamental Research Funds for the Central Universities.
+As the value of data assets and privacy issues arise, it is of utmost importance for data owners to make full use of proprietary information whilst seeking cooperation with their counterparts.
+Although federated learning-based paradigms have paved the way for privacy-preserving collaboration, its learning-centric nature greatly limits its feasibility over cost-sensitive, self-interested data owners.
+We argue that it is worth thinking of a new scenario by shifting the focus to data-centric and task-irrelevant cooperation, which substantially diversifies the way how participants can benefit from the federation.
+With this vision, we first conceptualize a new scenario of collaboration termed Task-Agnostic Federation, and then investigate latest research that enables exchange of generic knowledge to achieve collaborative data expansion, data refinement and data harmonization over decentralized data.
+Through a comprehensive literature review, we point out that the carrier of knowledge and the protocol for knowledge exchange are the most influential factors for TAF to be effective, efficient and extendable.
+Based on our investigation, we further highlight a number of open challenges that are not yet adequately addressed by existing effort, and also pose two open-ended questions to inspire more theoretical and empirical exploration in the domain.
+Acknowledgement
+This work is supported by the National Natural Science Foundation of China (62402198, U23B2027), National Social Science Foundation of China Post-funded Project (20FGLB034), Basic and Applied Basic Research Project of Guangzhou (2025A04J2212), the Fundamental Research Funds for the Central Universities (21624348), and New Generation Artificial Intelligence-National Science and Technology Major Project (2025ZD0123605).
+An earlier preprint version of this article entitled “Knowledge Augmentation in Federation: Rethinking What Collaborative Learning Can Bring Back to Decentralized Data” was posted on arXiv (DOI arXiv:2503.03140).